### v1

# PADetBench: Towards Benchmarking Physical Attacks against Object Detection

Abstract
Physical attacks against object detection have gained increasing attention due to their significant practical implications. However, conducting physical experiments is extremely time-consuming and labor-intensive. Moreover, physical dynamics and cross-domain transformation are challenging to strictly regulate in the real world, leading to unaligned evaluation and comparison, severely hindering the development of physically robust models. To accommodate these challenges, we explore utilizing realistic simulation to thoroughly and rigorously benchmark physical attacks with fairness under controlled physical dynamics and cross-domain transformation. This resolves the problem of capturing identical adversarial images that cannot be achieved in the real world. Our benchmark includes 20 physical attack methods, 48 object detectors, comprehensive physical dynamics, and evaluation metrics. We also provide end-to-end pipelines for dataset generation, detection, evaluation, and further analysis. In addition, we perform 8064 groups of evaluation based on our benchmark, which includes both overall evaluation and further detailed ablation studies for controlled physical dynamics. Through these experiments, we provide in-depth analyses of physical attack performance and physical adversarial robustness, draw valuable observations, and discuss potential directions for future research. Codebase: GitHub Repository.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable success in various research fields and industrial applications, such as computer vision [65], natural language processing [64], and speech recognition [61]. However, previous studies [78, 27, 2, 41, 3] have shown that DNNs are susceptible and vulnerable to adversarial attacks, which can be divided into digital attacks and physical attacks according to different attack domains in computer vision tasks. Technically, digital attacks are usually performed by adding imperceptible perturbations to the input image of victim models after imaging, while physical attacks can cheat DNNs by directly modifying the physical properties of the targets before imaging, such as changing the texture of the object [77, 114], and adding stickers [96, 44]. Intuitively, physical attacks are more dangerous and practical than digital attacks because they can be easily implemented in the real world with fewer prerequisites, which arouses significant concerns, especially in safety-critical applications like autonomous driving [92, 6], security surveillance [63, 95], remote sensing [94, 48], and so on.
Object detection is a fundamental and pragmatic task in computer vision, which has been widely deployed in various intelligent systems [126, 113]. Consequently, most of the physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has attracted increasing attention in recent years. However, the absence of regulated and easy-to-follow benchmarks hinders the development of physical attack and physically robust detection methods. The main reasons for discouraging the development of physical attack benchmarks are as follows. 1) Time-consuming and expensive: To evaluate the physical attack performance and physical adversarial robustness of object detection models, researchers need to conduct a large number of experiments in the real world, which is time-consuming and expensive. 2) Physical dynamics align: To ensure comparison fairness, physical dynamics should be strictly controlled and kept the same, while this can’t be achieved in real-world scenarios since it is impossible to take two identical pictures. 3) Cross-domain loss: Physical attacks are commonly conducted by elaborating conspicuous adversarial patterns that should be able to survive after transforming from the physical to the digital domain, while this cross-domain loss is uncontrollable. 4) Hard to compare: With the progressive evolution of physical attacks, physical adversarial perturbations have evolved from 2D space to 3D space, which builds a barrier to fairly compare different kinds of physical attack methods. As the above challenges exist, it is tough to effectively verify the physical attack efficacy and physical adversarial robustness of object detection models without thorough evaluation and impartial comparisons. As a result, the researchers can not identify the actual progress of physical adversarial attacks and physical adversarial robustness development, which may gear down the progress of the research field.
In this paper, we propose a fair and rigorous benchmarking of physical attacks on object detection models. We utilize CARLA [17], a highly realistic autonomous driving simulator, to benchmark physical attacks under consistent physical dynamics such as weather, viewing angle, and distance. These conditions are challenging to align for impartial comparisons in the real world due to the impossibility of capturing two identical photographs. Our benchmark includes 20 physical attack methods, clean and random noised objects for comparison, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, and a comprehensive pipeline for evaluating physical attacks and subsequent analysis. We conduct exhaustive evaluation of each physical attack against a wide range of object detectors under various physical dynamic conditions, using multiple evaluation metrics. Following our experimental procedures, we delve into detailed analyses concerning the performance of physical attacks and the robustness against physical adversarial actions. We highlight intriguing findings and discuss potential avenues for future research.
In summary, our contributions are as follows:
We propose a robust and equitable benchmark for physical attacks against object detection models. This benchmark deeply explores the potential of real-world simulators to consistently evaluate physical attacks under a variety of continuous physical dynamics.
We offer 8064 groups of evaluation on 20 physical attacks against 48 object detectors, which are conducted under continuous physical dynamics with diverse evaluation metrics, providing a holistic perspective on the performance of attack and detection.
We delve into detailed analyses of the fooling efficacy of physical attacks and the robustness of object detection models against these attacks. Through these analyses, we highlight valuable findings and discuss potential avenues for future research in this area.
2 Related work
2.1 Object detection
Object detection serves as a cornerstone in computer vision, with the primary objective of identifying and localizing objects within images or video sequences. With the advent of deep learning, DNNs have been widely adopted for object detection, significantly improving performance. R-CNN [26] and its successors [25, 70, 56, 66, 99, 106, 75], have made substantial strides in improving detection speed and accuracy by introducing region proposal networks and shared convolution computations. Single Shot MultiBox Detector (SSD) [55] and You Only Look Once (YOLO) [67] series [67, 68, 69, 1, 38, 43, 86, 37, 87, 85] further accelerate the detection process by eliminating the need for region proposals and performing detection in a single pass, making them suitable for real-time applications. Recently, more advanced architectures based on transformers have been proposed, such as DETR [7], DAB-DETR [52], ViTDet [47], DINO [104], and Co-DETR [125], which further push the boundaries of object detection performance by leveraging attention mechanisms. Despite these advancements, object detection in adversarial environments remains a challenging task, necessitating continuous research and development in this area.
2.2 Physical attack
Kurakin et al. [41] first illustrate that machine learning systems remain susceptible to adversarial examples, even within these physical world contexts. They substantiate this by introducing adversarial images, captured via a cell phone camera, which still can dramatically degrade the performance of intelligent vision systems. Brown et al. [2] first design adversarial perturbations in patch form, i.e., adversarial patch, which localizes adversarial perturbations to a specific region of the image without imperceptibility constraints. Adversarial patches are more practical and effective in the real world, as they can be easily printed and attached to objects to fool detectors [73, 81, 100, 124, 122, 91, 121, 34, 112, 72, 35, 29]. Nonetheless, the adversarial patches mentioned above are overly noticeable, potentially arousing the suspicion of the victims. To mitigate this concern, scholars have suggested the generation of adversarial patches in natural styles to circumvent drawing attention [36, 32, 28]. Beyond adversarial patches, the academic community has also investigated physical perturbations in a multitude of forms, including light [31, 98], viewpoint [15], and 3D objects [51], among others. To reconcile the disparity between 2D perturbations and 3D physical space, the concept of adversarial perturbations has been expanded into the 3D realm [111, 88, 76, 77, 117]. This extension has proven to be more efficacious and applicable in real-world scenarios. The considerable variation in physical perturbations and experimental settings presents a hurdle to the equitable comparison of disparate physical attack methodologies.
2.3 Robustness benchmark
The benchmarking of adversarial attacks is crucial for evaluating the robustness of DNNs-based models and further developing more robust ones. To address the issue of overestimated robustness, [12] first establish a standardized benchmark of adversarial robustness, which accurately reflects the robustness of the considered models within a reasonable computational budget. Wu et al. [97] build a comprehensive benchmark for backdoor attacks in image classification models. [60] provides an easy-to-use benchmark to assess the performance of object detection models under conditions of deteriorating image quality, such as image distortions or diverse weather conditions instead of adversarial perturbations. In contrast, [115] benchmark the adversarial robustness of image classifiers in black-box settings. Except for RGB images, Dong et al. [14] benchmark the robustness of 3D object detection to common corruptions from weather, sensor, motion, etc., in both LiDAR and camera data. Li et al. [45] towards benchmarking and assessing the visual naturalness of physical adversarial perturbations. Hingun et al. [30] construct a large-scale benchmark for evaluating adversarial patches with traffic sign dataset. CARLA [17], as a realistic autonomous driving simulator, has been progressively involved in the research of physical adversarial robustness. [62] presents CARLA-GEAR: a dataset generator for evaluating the adversarial robustness of vision models. The work in [109] put forth a pipeline for instance-level data generation, leveraging the CARLA simulator. Utilizing this pipeline, they created the DCI dataset and carried out experiments involving three detectors and three physical attacks. In summary, there is still a lack of a comprehensive and rigorous benchmark for physical attacks against object detection models. We aim to fill the gap in this work and significantly lower the threshold of this path with easy-to-follow instructions and codebase.
3 PADetBench
The benchmark encompasses four integral facets: datasets, physical attack implementations, object detection models, and comprehensive evaluation & analysis procedures. From a technical standpoint, we have engineered each constituent of the benchmark as modular, end-to-end pipelines within our publicly accessible codebase, ensuring straightforward adoption and replication.
3.1 Dataset Generation
It is common to use COCO [50], PASCAL VOC [19], KITTI [23], etc., as benchmark datasets for object detection. However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility required to create manipulated, real-world adversarial scenarios. Physical attacks typically entail altering the physical attributes of objects before capturing their images. To fairly and accurately evaluate and compare such attacks, experiments necessitate applying perturbations in real-world conditions with controlled physical dynamics, which are excessively time-consuming, labor-intensive, and theoretically infeasible. Simulated environments, like CARLA [17], present a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
This work contributes an end-to-end pipeline for dataset generation within our codebase, significantly streamlining the dataset generation process and enhancing research productivity. Our pipeline prioritizes user-friendliness, enabling researchers to swiftly generate datasets embodying diverse physical conditions through a concise series of steps. These conditions encompass variations in weather, viewing angles, and distances, along with the capacity to impose physical perturbations on objects. Comprehensively, our pipeline supports over 10 distinct environments ranging from downtowns to small towns and rural landscapes (detailed listing available in the Appendix), coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures. It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning both location and orientation. To ensure accessibility, we accompany the pipeline with step-by-step guidelines for personalizing object perturbations and seamlessly integrating these modifications within CARLA’s [17] simulation framework.
Our benchmark comprises three categories of datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies. To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline.
Moreover, our pipeline facilitates the automatic generation of supplementary annotations, including 2D and 3D bounding boxes, depth maps, and instance segmentation maps. Consequently, our benchmark extends its utility beyond 2D object detection, also catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
3.2 Physical attacks
We select 20 representative physical attack methods, in which FCA [88], DTA [76], ACTIVE [77], 3D2Fool [114], POOPatch [11], RPAU [54], APPA [48], and CAMOU [111], are for vehicle detection, while DAP [29], AdvPattern [95], UPC [36], NatPatch [32], MTD [13], AdvCaT [33], AdvTexture [34], AdvTshirt[101], AdvPatch [81], LAP [79], InvisCloak [100], and AdvCam [18], are for person detection. The corresponding physical perturbations of these methods are imported into Unreal Engine 4 for CARLA (CarlaUE4) [17], as shown in the physical attacks part of Fig. 1, to generate the physical adversarial datasets. We adhere to two principles similar to [97] when selecting physical attacks. First, the methods are representative or advanced in the research field, which can serve as baseline and state-of-the-art (SOTA) methods for comparison, respectively. Second, physical attacks are easily conducted and with reproducible performance, which can be conveniently followed and reproduced by other researchers. Since our benchmark evaluates physical attacks based on their crafted perturbations, novel physical attack methods can be easily integrated into the benchmark by following the provided pipeline. We will continue to update the physical attacks in the benchmark to keep pace with the latest research progress.
Backbone | Category | Detectors |
CNN | One-stage | ATSS[107], AutoAssign[119], CenterNet[118], CentripetalNet[16] |
CornerNet[42], DDOD[10], DyHead[99], EfficientNet[80], FCOS[82], | ||
FoveaBox[40], FreeAnchor[110], FSAF[120], GFL[46],LD[116], NAS-FPN[24], | ||
PAA[39], RetinaNet[49], RTMDet[58], TOOD[20], VarifocalNet[105], | ||
YOLOX[22], YOLOv5[38], YOLOv6[43], YOLOv7[86], YOLOv8[37] | ||
Two-stage | Faster R-CNN[70], Cascade R-CNN[5], Cascade RPN[84], | |
Double Heads[99], FPG[8],Grid R-CNN[56], Guided Anchoring[89], | ||
HRNet[74], Libra R-CNN[66], PAFPN[53],RepPoints[102], | ||
Res2Net[21], ResNeSt[103], SABL[90], Sparse R-CNN[75] | ||
Transformer | - | DETR[7], Conditional DETR[59], DDQ[108], DAB-DETR[52], |
Deformable DETR[123], DINO[104],PVT[93], PVTv2[93] |
3.3 Object detectors
We choose 48 object detectors in the same principles as choosing physical attack methods, covering mainstream object detectors, such as YOLO series [38, 43, 86, 37, 22] (One-stage) and R-CNN series [26, 25, 70, 4, 75], which are based on CNN. Except for canonical detectors, we also include transformer-based detectors, such as DETR [7], Conditional DETR [59], Deformable DETR [123], DAB-DETR [52], and DINO [104]. All the selected detectors are listed in Table 1 according to their characteristics. Our benchmark provides the end-to-end pipeline for object detection evaluation based on MMDetection [9]. Consequently, it is convenient to integrate new detectors into the benchmark, and the benchmark can also be easily extended to evaluate other vision tasks, such as 3D object detection, instance segmentation, and depth estimation.
3.4 Evaluation and analysis
To rigorously assess the efficacy of physical attacks on object detection systems, we furnish baseline datasets: clean datasets (without perturbations) and those infused with randomized noise (incorporating arbitrary disturbances in -bounded space). This dual-baseline approach sets the stage for a thorough examination.
Quantifying performance entails employing evaluation metrics that consider both detection accuracy and the performance of the attacks. These metrics comprise by several widely adopted indicators, including average recall (AR) and Attack Success Rate (ASR). AR is calculated as the average of the recall rates:
where TP and FN denote the true positive and false negative counts of the detector, respectively. AR signifies the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given target. On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as:
where and denote the average recall rates of the attack and clean datasets, respectively. ASR provides a direct measure of the extent to which the attacks undermine the detector’s confidence in its predictions.
Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating deeper insights into model behavior. These include Grad-CAM [71] for visualizing the regions of input data that contribute most to the model’s prediction, Shapley value [57] to quantify the individual feature contributions, and t-SNE [83] for reducing dimensionality and visualizing high-dimensional data in a more interpretable manner. These additions empower users to conduct comprehensive analyses beyond mere performance evaluation.
4 Experiments
4.1 Experimental setup
Datasets. 1) Overall experiments. We generate overall datasets with 3 objects, 10 weather conditions, 2 altitude angles, 8 azimuth angles, 5 radius values, 3 spawn points, and 20 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison. 2) Ablation Studies. We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, altitude angle within a hemispherical space. Accomplishing this involves generating focused sub-benchmarks, each consisting of 100 samples.
Physical attacks. we generate 24 datasets, including 20 physically noised datasets that correspond to 20 physical attacks, an extra 2 clean datasets and 2 randomly noised datasets for comparison of vehicle detection and person detection, respectively.
Object detectors. We evaluate 48 object detectors, as shown in Table 1, by integrating MMDetection [9] into our evaluation pipeline, in which the selected detectors are trained on the COCO [50] dataset.
Therefore, we conduct a total of 8064 () groups of the experiment, which are conducted with NVIDIA Geforce 4090.
4.2 Overall experiments and analysis
We present the comprehensive results of vehicle detection against physical attacks in Fig. 2. Additionally, Fig. 3 and Fig. 4 show visualized analyses of the experimental results from detection and attack perspectives, respectively. Experimental results of person detection and corresponding detailed numerical results are listed in the Appendix. From these analyses, several key observations emerge:
Detection perspective: 1) Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing up to 50%, as shown in Fig. 4. However, pedestrian detection performance is less affected regarding various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Fig. 3. The potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as shown in the ablation experiments, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks. 2) For different detectors, the performance varies significantly, with some detectors exhibiting superior robustness against physical attacks like EfficientNet, YOLO series and RTMDet of one-stage detectors. Besides, DDQ also shows significant adversarial robustness among transformer-based detectors. While the rest of the detectors show varying lower levels of robustness, the SOTA detection performance is not necessarily indicative of adversarial robustness. Consequently, the benchmark also serves as a robustness indicator.
Attack perspective: 1) For vehicle detection, different physical attacks exhibit varying levels of effectiveness, with some attacks achieving ASR values exceeding 70% like ACTIVE, and others failing to surpass 20%. Most of the physical attacks hard to fool the latest SOTA detectors, such as EfficientNet, YOLO series, and RTMDet. This phenomenon is caused by the victim models of the attack method lagging behind the development of the detection method, which also motivates us to fill this gap. 2) For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with the majority of attacks achieving ASR values below 20%, even lower than 0% (see more explanation in the Appendix). The relatively strongest attack method is AdvTexture, which elaborates on a 2D patch but with tricks for 3D space. This also demonstrates the gap between 2D perturbations and 3D physical space.
4.3 Ablation experiments and analysis
Except for the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors. We show the results of 3 physical dynamics, including weather, distance, and camera viewing angle, in Fig. 5 and Fig. 6, respectively. More experiments on other dynamics are provided in the Appendix. From these analyses, several key observations emerge: 1) Physical attack performance can be easily swayed by physical dynamics. This phenomenon is consistent with our intuition and emphasizes the importance of strictly aligning physical dynamics when evaluating physical attacks, which are often underestimated by previous works. 2) We also observe a gap between the attack performance of our benchmark and the reported performance in the original papers (supplemented in Appendix). Two reasons may contribute to this gap: the first is the adopted SOTA detectors in our benchmark, which are more robust than the victim models in the original papers, and the second is that our benchmark provides more comprehensive and strict evaluation datasets and physical dynamics, which are more challenging for the attack methods. Please refer to the Appendix for more experiments, detailed analysis and discussion.
5 Discussion
Limitations. For now, PADetBench mainly focuses on the physical attacks on object detection models. In the future, we will extend the benchmark to evaluate other vision tasks, such as instance segmentation, 3D object detection, and depth estimation.
Potential impacts. 1) Positive impacts: In-depth understanding of the robustness of object detection models against physical attacks, which can guide the development of more robust models. 2) Negative impacts: The benchmark may be misused to conduct physical attacks in real life, which may pose a threat to the security-critical applications involving intelligent visual perception.
6 Conclusion
In conclusion, this work aims to meet the critical need for a standardized and rigorous benchmark for physical attacks on object detection models. Our benchmark leverages the CARLA simulator for consistent physical dynamics and includes 20 attack methods, 48 detectors, comprehensive dynamics, and metrics from different perspectives. It also provides end-to-end pipelines for dataset generation, attack and detection evaluation, and experimental results analysis. Our extensive evaluation and analyses provide valuable insights for future research directions. The supporting codebase is publicly available and we will continuously supplement it, aiding the development of physically robust object detection methods.
7 Acknowledgment
The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.
Checklist
1.
For all authors…
Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
Did you describe the limitations of your work? [Yes]
Did you discuss any potential negative societal impacts of your work? [Yes]
Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
2.
If you are including theoretical results…
Did you state the full set of assumptions of all theoretical results? [N/A]
Did you include complete proofs of all theoretical results? [N/A]
3.
If you ran experiments (e.g. for benchmarks)…
Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] Please refer to the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack for the code and instructions.
Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] This benchmark does not involve a training process.
Did you report error bars (e.g., to the random seed after running experiments multiple times)? [N/A]
Did you include the total amount of computing and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
4.
If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
- (a)
Did you mention the license of the assets? [Yes] Please refer to the Appendix for the license.
Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the GitHub repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack) for details.
Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
5.
If you used crowdsourcing or conducted research with human subjects…
Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
Appendix A Additional content of the Benchmark
A.1 Mini-test
We kindly invite the reviewers and readers to participate in a mini-test to discriminate the real-world images and the simulated images as shown in Fig. 7, the answer is revealed in its caption.
Town | Description |
Town1 | A small, simple town with a river and several bridges. |
Town2 | A small simple town with a mixture of residential and commercial buildings. |
Town3 | A larger, urban map with a roundabout and large junctions. |
Town4 | A small town embedded in the mountains with a special "figure of 8" infinite highway. |
Town5 | Squared-grid town with cross junctions and a bridge. |
It has multiple lanes per direction. Useful to perform lane changes. | |
Town6 | Long many lane highways with many highway entrances and exits. It also has a Michigan left. |
Town7 | A rural environment with narrow roads, corn, barns and hardly any traffic lights. |
Town8 | Secret "unseen" town used for the Leaderboard challenge. |
Town9 | Secret "unseen" town used for the Leaderboard challenge. |
Town10 | A downtown urban environment with skyscrapers, residential buildings and an ocean promenade. |
Town11 | A Large Map that is undecorated. |
Serves as a proof of concept for the Large Maps feature. | |
Town12 | A Large Map with numerous different regions, |
including high-rise, residential and rural environments. |
Full list of the optional maps, where Town8 and Town9 are unseen for competition. Please refer to CARLA [17] documentary for more details.
A.2 Corresponding config files of the selected detectors
The corresponding config files of the selected detectors are listed in Table 3.
Number | Config Files | Detectors |
1 | atss_r50_fpn_1x_coco | ATSS[107] |
2 | autoassign_r50-caffe_fpn_1x_coco | AutoAssign[119] |
3 | centernet-update_r50-caffe_fpn_ms-1x_coco | CenterNet[118] |
4 | centripetalnet_hourglass104_16xb6-crop511-210e-mstest_coco | CentripetalNet[16] |
5 | cornernet_hourglass104_10xb5-crop511-210e-mstest_coco | CornerNet[42] |
6 | ddod_r50_fpn_1x_coco | DDOD[10] |
7 | atss_r50_fpn_dyhead_1x_coco | DyHead[99] |
8 | retinanet_effb3_fpn_8xb4-crop896-1x_coco | EfficientNet[80] |
9 | fcos_x101-64x4d_fpn_gn-head_ms-640-800-2x_coco | FCOS[82] |
10 | fovea_r50_fpn_4xb4-1x_coco | FoveaBox[40] |
11 | freeanchor_r50_fpn_1x_coco | FreeAnchor[110] |
12 | fsaf_r50_fpn_1x_coco | FSAF[120] |
13 | gfl_r50_fpn_1x_coco | GFL[46] |
14 | ld_r50-gflv1-r101_fpn_1x_coco | LD[116] |
15 | retinanet_r50_nasfpn_crop640-50e_coco | NAS-FPN[24] |
16 | paa_r50_fpn_1x_coco | PAA[39] |
17 | retinanet_r50_fpn_1x_coco | RetinaNet[49] |
18 | rtmdet_s_8xb32-300e_coco | RTMDet[58] |
19 | tood_r50_fpn_1x_coco | TOOD[20] |
20 | vfnet_r50_fpn_1x_coco | VarifocalNet[105] |
21 | yolov5_l-p6-v62_syncbn_fast_8xb16-300e_coco | YOLOv5[38] |
22 | yolov6_l_syncbn_fast_8xb32-300e_coco | YOLOv6[43] |
23 | yolov7_l_syncbn_fast_8x16b-300e_coco | YOLOv7[86] |
24 | yolov8_l_syncbn_fast_8xb16-500e_coco | YOLOv8[37] |
25 | yolox_l_fast_8xb8-300e_coco | YOLOX[22] |
26 | faster-rcnn_r50_fpn_1x_coco | Faster R-CNN[70] |
27 | cascade-rcnn_r50_fpn_1x_coco | Cascade R-CNN[5] |
28 | cascade-rpn_faster-rcnn_r50-caffe_fpn_1x_coco | Cascade RPN[84] |
29 | dh-faster-rcnn_r50_fpn_1x_coco | Double Heads[99] |
30 | faster-rcnn_r50_fpg_crop640-50e_coco | FPG[8] |
31 | grid-rcnn_r50_fpn_gn-head_2x_coco | Grid R-CNN[56] |
32 | ga-faster-rcnn_x101-32x4d_fpn_1x_coco | Guided Anchoring[89] |
33 | faster-rcnn_hrnetv2p-w18-1x_coco | HRNet[74] |
34 | libra-retinanet_r50_fpn_1x_coco | Libra R-CNN[66] |
35 | faster-rcnn_r50_pafpn_1x_coco | PAFPN[53] |
36 | reppoints-moment_r50_fpn_1x_coco | RepPoints[102] |
37 | faster-rcnn_res2net-101_fpn_2x_coco | Res2Net[21] |
38 | faster-rcnn_s50_fpn_syncbn-backbone+head_ms-range-1x_coco | ResNeSt[103] |
39 | sabl-faster-rcnn_r50_fpn_1x_coco | SABL[90] |
40 | sparse-rcnn_r50_fpn_1x_coco | Sparse R-CNN[75] |
41 | detr_r50_8xb2-150e_coco | DETR[7] |
42 | conditional-detr_r50_8xb2-50e_coco | Conditional DETR[59] |
43 | ddq-detr-4scale_r50_8xb2-12e_coco | DDQ[108] |
44 | dab-detr_r50_8xb2-50e_coco | DAB-DETR[52] |
45 | deformable-detr_r50_16xb2-50e_coco | Deformable DETR[123] |
46 | dino-4scale_r50_8xb2-12e_coco | DINO[104] |
47 | retinanet_pvt-t_fpn_1x_coco | PVT[93] |
48 | retinanet_pvtv2-b0_fpn_1x_coco | PVTv2[93] |
Appendix B Additional content of the Experiments
B.1 Genetated data for the ablation studies
We provide the generated data samples for the ablation studies in Fig. 8.
B.2 Detailed illustration of the gap between the attack performance of our benchmark and the reported performance in the original papers
Our benchmark addresses two important aspects that were not considered in the original papers:
1.
Strict physical dynamics: Our benchmark leverages the CARLA simulator, which provides a realistic and consistent physical environment for evaluating physical attacks on object detection models. This ensures that the attacks are evaluated under conditions that closely resemble the physical world, allowing for more accurate assessment of the robustness of the models.
2.
Evaluation in 3D space perspective: Unlike the original papers, which mainly focus on 2D evaluation, our benchmark evaluates the attack performance in a 3D space perspective. This is crucial because physical attacks can exploit vulnerabilities in the 3D structure of the objects, such as comprehensive and continuous view, scale, distance, etc. By considering the 3D space perspective, our benchmark provides a more comprehensive evaluation of the models’ robustness against physical attacks.
By addressing these two aspects, our benchmark provides a more realistic and comprehensive evaluation of the attack performance of object detection models, bridging the gap between the reported performance in the original papers and the real-world scenarios.
B.3 Supplemented experiments analysis and discussion
B.3.1 Detection Perspective
Vehicle Detection Perspective:
Physical attacks on vehicle detection systems pose a substantial challenge due to the specialized nature of the perturbations crafted to deceive these models. These attacks can lead to a drastic decline in average recall rates, reaching as low as 50%. This high level of vulnerability is largely attributed to the complex dynamics in the 3D environment where vehicles operate. Physical attacks on vehicle detectors exploit this three-dimensional context, introducing perturbations that consider real-world factors such as lighting, perspective, occlusion, and motion, making them more effective in disrupting the model’s performance.
On the other hand, pedestrians, operating in a somewhat simpler 2D plane, seem to be less affected by similar adversarial attacks, with a decrease in average recall rates of less than 20%. Adversarial examples targeting pedestrian detection typically involve 2D patches, which might be more straightforward to apply in specific scenarios but may not account for the full range of real-world complexities. As a result, there is an urgent demand to establish a comprehensive and stringent benchmark to systematically evaluate the resilience of these models against physical attacks, facilitating research and development towards more secure systems.
Pedestrian Detection Perspective:
In contrast to vehicle detection, pedestrian detectors exhibit a certain level of inherent robustness, potentially due to the simpler constraints imposed on the recognition process. Nevertheless, as seen across various detectors, the extent of this robustness varies widely. Models like EfficientNet, YOLO series, RTMDet (one-stage detectors), and DDQ (transformer-based detectors) demonstrate commendable resistance to physical attacks. The superior performance of DDQ could be linked to the attention mechanisms inherent to transformer architectures, which are capable of capturing global spatial dependencies, thus mitigating the impact of adversarial perturbations.
However, it is evident from the benchmark results that not all state-of-the-art (SOTA) detectors offer comparable adversarial robustness. Many detectors exhibit varying degrees of vulnerability, indicating that peak accuracy in standard detection tasks does not automatically guarantee resilience against adversarial threats. Consequently, this benchmarking framework not only identifies areas of weakness for refinement but also contributes to a better understanding of the interplay between detection performance and adversarial robustness in real-world deployments.
In conclusion, understanding and mitigating the effects of physical attacks in both vehicle and pedestrian detection domains can greatly benefit deep learning and computer vision research. By developing more robust models resistant to such attacks, we can enhance the safety and reliability of autonomous systems that rely on accurate object detection, ultimately fostering advancements in the fields of automotive technology, smart city infrastructure, and robotics. Furthermore, this benchmark would encourage researchers to explore defensive techniques and novel architectures that better withstand both digital and physical adversarial threats, pushing the boundaries of deep learning and computer vision capabilities.
B.3.2 Attack Perspective
From the attacker’s viewpoint, the effectiveness of physical attacks on deep learning-based vehicle detection systems is highly variant. Certain methodologies, such as ACTIVE, achieve astonishingly high success rates in defeating the detectors, with ASR values surpassing 70%. However, the majority of current attacks struggle to maintain comparable performance, often failing to reach even 20% ASR. This discrepancy can be partly attributed to the rapid advancements in detection algorithms, with the latest state-of-the-art models like EfficientNet, YOLO series, and RTMDet demonstrating increased resilience against known attacks. This disparity in the evolutionary pace between attackers and defenders underscores the importance of continuous research and innovation in adversarial attacks to keep pace with the evolving landscape of detection techniques.
Moreover, this evolving dynamic underscores a critical need for a more dynamic and collaborative ecosystem in deep learning and computer vision research. By closing the gap between attack methods and detector capabilities, the field will likely see increased robustness and security measures, ultimately benefiting automotive safety and other real-world applications relying on these systems.
On the other hand, when evaluating person detection, the outcome of physical attacks exhibits a different pattern, with ASR values typically remaining below 20%, and often even below 0%, which indicates that the attack method is less effective than random guessing and the eye-catching perturbation may arouse more attention than the object itself. Additionally, the variable transferability of these attack methodologies across different detectors leads to a wide disparity in ASR values. In certain instances, this manifests as negative ASR figures, indicative of a backfiring effect where the detectors become more adept at identifying targets in the presence of attempted attacks.
The significantly lower effectiveness of these attacks on pedestrian detection models highlights the comparative advantages of their 2D nature against primarily 2D adversarial perturbations. Nevertheless, the AdvTexture method, despite being a 2D approach, manages to incorporate 3D considerations, achieving higher ASRs compared to other attacks. This underscores the pivotal role of incorporating 3D awareness into attack strategies to exploit the vulnerabilities of pedestrian detectors more effectively.
These contrasting observations highlight the need for more sophisticated attack methods in the domain of pedestrian detection. By advancing the understanding of how 2D techniques can be adapted or combined with 3D concepts, attackers can create more potent adversarial samples, driving defender-side innovation to fortify models further. Such advancements will ultimately contribute to the progression of the field by promoting the design of more secure and reliable computer vision systems, particularly relevant in surveillance, autonomous navigation, and smart city infrastructures.
In summary, the diverse outcomes of physical attacks on both vehicle and person detection emphasize the importance of ongoing research and competition between attack and defense approaches. As the attacks become more intricate and align with the complex nature of real-world scenarios, deep learning and computer vision models will adapt, increasing their resilience and overall functionality. This continuous push-and-pull between adversaries and protectors fosters the evolution of robust, secure, and accurate object-detection technologies essential for numerous applications, including automotive safety, surveillance, and urban automation.
B.4 Additional overall experimental results
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 4, 5, 6, 7, 8, 9, 10, and 11. In addition, the visualized evaluation results are shown in Fig. 9, 10, 11, 12, 13, 14, and 15.
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.83 | 0.477 | 0.231 | 0.545 | 0.606 | 0.502 | 0.434 | 0.678 | 0.235 | 0.532 |
AutoAssign | 0.786 | 0.574 | 0.37 | 0.559 | 0.589 | 0.609 | 0.415 | 0.722 | 0.487 | 0.648 |
CenterNet | 0.839 | 0.558 | 0.297 | 0.552 | 0.57 | 0.58 | 0.426 | 0.742 | 0.412 | 0.521 |
CentripetalNet | 0.78 | 0.685 | 0.558 | 0.725 | 0.687 | 0.725 | 0.527 | 0.801 | 0.501 | 0.648 |
CornerNet | 0.748 | 0.586 | 0.438 | 0.652 | 0.593 | 0.653 | 0.458 | 0.779 | 0.429 | 0.582 |
DDOD | 0.838 | 0.708 | 0.433 | 0.695 | 0.686 | 0.745 | 0.548 | 0.785 | 0.694 | 0.646 |
DyHead | 0.876 | 0.611 | 0.385 | 0.566 | 0.725 | 0.614 | 0.574 | 0.671 | 0.402 | 0.73 |
EfficientNet | 0.881 | 0.711 | 0.506 | 0.687 | 0.721 | 0.764 | 0.638 | 0.763 | 0.71 | 0.665 |
FCOS | 0.933 | 0.804 | 0.676 | 0.838 | 0.795 | 0.855 | 0.658 | 0.894 | 0.76 | 0.824 |
FoveaBox | 0.814 | 0.597 | 0.294 | 0.514 | 0.645 | 0.548 | 0.467 | 0.649 | 0.469 | 0.618 |
FreeAnchor | 0.81 | 0.51 | 0.381 | 0.611 | 0.563 | 0.643 | 0.431 | 0.638 | 0.336 | 0.582 |
FSAF | 0.788 | 0.51 | 0.233 | 0.566 | 0.559 | 0.537 | 0.432 | 0.661 | 0.364 | 0.529 |
GFL | 0.852 | 0.509 | 0.202 | 0.456 | 0.626 | 0.485 | 0.485 | 0.63 | 0.201 | 0.602 |
LD | 0.825 | 0.554 | 0.305 | 0.563 | 0.658 | 0.548 | 0.463 | 0.664 | 0.29 | 0.591 |
NAS-FPN | 0.87 | 0.623 | 0.473 | 0.662 | 0.673 | 0.695 | 0.5 | 0.764 | 0.382 | 0.655 |
PAA | 0.808 | 0.582 | 0.474 | 0.621 | 0.605 | 0.619 | 0.501 | 0.685 | 0.567 | 0.64 |
RetinaNet | 0.85 | 0.511 | 0.349 | 0.565 | 0.653 | 0.568 | 0.479 | 0.684 | 0.43 | 0.584 |
RTMDet | 0.875 | 0.717 | 0.625 | 0.771 | 0.733 | 0.753 | 0.736 | 0.821 | 0.676 | 0.72 |
TOOD | 0.781 | 0.495 | 0.353 | 0.522 | 0.584 | 0.557 | 0.462 | 0.615 | 0.37 | 0.572 |
VarifocalNet | 0.874 | 0.472 | 0.205 | 0.424 | 0.628 | 0.529 | 0.419 | 0.573 | 0.208 | 0.538 |
YOLOv5 | 0.886 | 0.76 | 0.744 | 0.762 | 0.807 | 0.821 | 0.788 | 0.857 | 0.812 | 0.75 |
YOLOv6 | 0.907 | 0.824 | 0.747 | 0.834 | 0.833 | 0.887 | 0.776 | 0.896 | 0.851 | 0.784 |
YOLOv7 | 0.906 | 0.774 | 0.762 | 0.829 | 0.822 | 0.866 | 0.786 | 0.903 | 0.774 | 0.803 |
YOLOv8 | 0.929 | 0.791 | 0.761 | 0.812 | 0.838 | 0.873 | 0.793 | 0.917 | 0.74 | 0.803 |
YOLOX | 0.908 | 0.766 | 0.683 | 0.783 | 0.817 | 0.851 | 0.794 | 0.849 | 0.702 | 0.782 |
Faster R-CNN | 0.772 | 0.375 | 0.141 | 0.338 | 0.509 | 0.46 | 0.369 | 0.612 | 0.268 | 0.438 |
Cascade R-CNN | 0.802 | 0.483 | 0.297 | 0.488 | 0.574 | 0.607 | 0.407 | 0.673 | 0.334 | 0.532 |
Cascade RPN | 0.805 | 0.53 | 0.291 | 0.452 | 0.588 | 0.55 | 0.441 | 0.662 | 0.452 | 0.548 |
Double Heads | 0.797 | 0.521 | 0.295 | 0.539 | 0.537 | 0.621 | 0.404 | 0.713 | 0.445 | 0.516 |
FPG | 0.846 | 0.678 | 0.486 | 0.714 | 0.671 | 0.749 | 0.503 | 0.806 | 0.47 | 0.65 |
Grid R-CNN | 0.795 | 0.472 | 0.244 | 0.513 | 0.529 | 0.65 | 0.399 | 0.699 | 0.392 | 0.494 |
Guided Anchoring | 0.904 | 0.723 | 0.555 | 0.747 | 0.748 | 0.781 | 0.524 | 0.828 | 0.738 | 0.717 |
HRNet | 0.76 | 0.547 | 0.29 | 0.52 | 0.512 | 0.571 | 0.336 | 0.738 | 0.462 | 0.511 |
Libra R-CNN | 0.78 | 0.49 | 0.294 | 0.527 | 0.563 | 0.492 | 0.486 | 0.664 | 0.334 | 0.54 |
PAFPN | 0.8 | 0.497 | 0.206 | 0.463 | 0.562 | 0.54 | 0.413 | 0.682 | 0.383 | 0.457 |
RepPoints | 0.847 | 0.576 | 0.222 | 0.525 | 0.547 | 0.557 | 0.427 | 0.712 | 0.565 | 0.523 |
Res2Net | 0.874 | 0.64 | 0.494 | 0.698 | 0.72 | 0.74 | 0.482 | 0.797 | 0.601 | 0.716 |
ResNeSt | 0.837 | 0.502 | 0.352 | 0.535 | 0.587 | 0.499 | 0.538 | 0.493 | 0.407 | 0.555 |
SABL | 0.796 | 0.46 | 0.262 | 0.501 | 0.563 | 0.535 | 0.423 | 0.648 | 0.359 | 0.484 |
Sparse R-CNN | 0.774 | 0.469 | 0.257 | 0.398 | 0.604 | 0.418 | 0.44 | 0.518 | 0.316 | 0.532 |
DETR | 0.636 | 0.114 | 0.048 | 0.033 | 0.351 | 0.17 | 0.333 | 0.198 | 0.025 | 0.339 |
Conditional DETR | 0.793 | 0.554 | 0.408 | 0.575 | 0.644 | 0.617 | 0.525 | 0.7 | 0.457 | 0.671 |
DDQ | 0.809 | 0.55 | 0.457 | 0.531 | 0.649 | 0.676 | 0.631 | 0.626 | 0.453 | 0.629 |
DAB-DETR | 0.838 | 0.391 | 0.194 | 0.308 | 0.616 | 0.473 | 0.488 | 0.576 | 0.163 | 0.526 |
Deformable DETR | 0.827 | 0.642 | 0.371 | 0.528 | 0.641 | 0.67 | 0.46 | 0.662 | 0.525 | 0.626 |
DINO | 0.78 | 0.351 | 0.236 | 0.322 | 0.543 | 0.56 | 0.423 | 0.526 | 0.217 | 0.49 |
PVT | 0.828 | 0.719 | 0.355 | 0.648 | 0.711 | 0.802 | 0.547 | 0.853 | 0.592 | 0.52 |
PVTv2 | 0.845 | 0.666 | 0.494 | 0.763 | 0.621 | 0.844 | 0.425 | 0.803 | 0.704 | 0.476 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.238 | 0.156 | 0.077 | 0.182 | 0.183 | 0.164 | 0.133 | 0.212 | 0.087 | 0.166 |
AutoAssign | 0.238 | 0.183 | 0.126 | 0.189 | 0.182 | 0.199 | 0.139 | 0.236 | 0.164 | 0.212 |
CenterNet | 0.238 | 0.167 | 0.093 | 0.178 | 0.165 | 0.182 | 0.14 | 0.23 | 0.126 | 0.156 |
CentripetalNet | 0.228 | 0.215 | 0.164 | 0.225 | 0.206 | 0.22 | 0.161 | 0.245 | 0.155 | 0.194 |
CornerNet | 0.218 | 0.184 | 0.132 | 0.198 | 0.175 | 0.199 | 0.142 | 0.237 | 0.129 | 0.173 |
DDOD | 0.242 | 0.21 | 0.127 | 0.221 | 0.202 | 0.234 | 0.174 | 0.242 | 0.211 | 0.193 |
DyHead | 0.26 | 0.196 | 0.121 | 0.18 | 0.221 | 0.191 | 0.18 | 0.21 | 0.126 | 0.227 |
EfficientNet | 0.252 | 0.225 | 0.168 | 0.217 | 0.219 | 0.239 | 0.206 | 0.243 | 0.232 | 0.21 |
FCOS | 0.277 | 0.251 | 0.211 | 0.266 | 0.246 | 0.272 | 0.214 | 0.285 | 0.236 | 0.256 |
FoveaBox | 0.235 | 0.184 | 0.089 | 0.165 | 0.191 | 0.17 | 0.145 | 0.196 | 0.149 | 0.193 |
FreeAnchor | 0.241 | 0.159 | 0.111 | 0.19 | 0.179 | 0.205 | 0.138 | 0.205 | 0.098 | 0.184 |
FSAF | 0.231 | 0.171 | 0.077 | 0.187 | 0.181 | 0.181 | 0.141 | 0.213 | 0.129 | 0.179 |
GFL | 0.244 | 0.169 | 0.064 | 0.152 | 0.192 | 0.163 | 0.157 | 0.201 | 0.069 | 0.192 |
LD | 0.235 | 0.176 | 0.095 | 0.181 | 0.205 | 0.176 | 0.146 | 0.208 | 0.094 | 0.187 |
NAS-FPN | 0.256 | 0.199 | 0.146 | 0.209 | 0.213 | 0.215 | 0.16 | 0.251 | 0.124 | 0.208 |
PAA | 0.237 | 0.178 | 0.139 | 0.189 | 0.178 | 0.195 | 0.157 | 0.208 | 0.173 | 0.194 |
RetinaNet | 0.249 | 0.169 | 0.108 | 0.194 | 0.209 | 0.189 | 0.166 | 0.227 | 0.155 | 0.191 |
RTMDet | 0.254 | 0.227 | 0.185 | 0.235 | 0.224 | 0.232 | 0.239 | 0.241 | 0.209 | 0.221 |
TOOD | 0.231 | 0.155 | 0.109 | 0.159 | 0.173 | 0.176 | 0.14 | 0.183 | 0.118 | 0.169 |
VarifocalNet | 0.248 | 0.144 | 0.063 | 0.127 | 0.188 | 0.164 | 0.132 | 0.175 | 0.062 | 0.162 |
YOLOv5 | 0.259 | 0.227 | 0.223 | 0.23 | 0.237 | 0.249 | 0.244 | 0.253 | 0.246 | 0.221 |
YOLOv6 | 0.256 | 0.245 | 0.218 | 0.251 | 0.242 | 0.262 | 0.238 | 0.263 | 0.25 | 0.229 |
YOLOv7 | 0.265 | 0.241 | 0.225 | 0.252 | 0.246 | 0.262 | 0.241 | 0.272 | 0.227 | 0.243 |
YOLOv8 | 0.276 | 0.246 | 0.239 | 0.254 | 0.252 | 0.269 | 0.256 | 0.283 | 0.236 | 0.246 |
YOLOX | 0.263 | 0.233 | 0.212 | 0.236 | 0.237 | 0.253 | 0.248 | 0.251 | 0.217 | 0.233 |
Faster R-CNN | 0.212 | 0.117 | 0.042 | 0.11 | 0.159 | 0.145 | 0.111 | 0.193 | 0.087 | 0.137 |
Cascade R-CNN | 0.232 | 0.152 | 0.088 | 0.15 | 0.182 | 0.19 | 0.135 | 0.214 | 0.115 | 0.169 |
Cascade RPN | 0.229 | 0.157 | 0.083 | 0.132 | 0.175 | 0.157 | 0.137 | 0.201 | 0.122 | 0.166 |
Double Heads | 0.238 | 0.168 | 0.09 | 0.179 | 0.171 | 0.205 | 0.143 | 0.231 | 0.15 | 0.164 |
FPG | 0.247 | 0.222 | 0.149 | 0.224 | 0.21 | 0.233 | 0.161 | 0.265 | 0.152 | 0.216 |
Grid R-CNN | 0.231 | 0.154 | 0.078 | 0.173 | 0.17 | 0.208 | 0.135 | 0.225 | 0.127 | 0.164 |
Guided Anchoring | 0.269 | 0.239 | 0.176 | 0.244 | 0.235 | 0.243 | 0.174 | 0.265 | 0.244 | 0.229 |
HRNet | 0.219 | 0.171 | 0.091 | 0.164 | 0.159 | 0.173 | 0.114 | 0.236 | 0.148 | 0.165 |
Libra R-CNN | 0.234 | 0.161 | 0.094 | 0.18 | 0.179 | 0.169 | 0.159 | 0.222 | 0.115 | 0.175 |
PAFPN | 0.219 | 0.147 | 0.057 | 0.145 | 0.17 | 0.168 | 0.122 | 0.205 | 0.129 | 0.139 |
RepPoints | 0.251 | 0.18 | 0.068 | 0.167 | 0.172 | 0.185 | 0.141 | 0.234 | 0.189 | 0.166 |
Res2Net | 0.25 | 0.195 | 0.154 | 0.217 | 0.212 | 0.22 | 0.162 | 0.244 | 0.192 | 0.22 |
ResNeSt | 0.23 | 0.156 | 0.101 | 0.158 | 0.174 | 0.154 | 0.169 | 0.142 | 0.126 | 0.162 |
SABL | 0.233 | 0.146 | 0.08 | 0.159 | 0.173 | 0.177 | 0.139 | 0.199 | 0.123 | 0.155 |
Sparse R-CNN | 0.238 | 0.159 | 0.095 | 0.137 | 0.195 | 0.145 | 0.154 | 0.172 | 0.128 | 0.168 |
DETR | 0.186 | 0.047 | 0.017 | 0.01 | 0.113 | 0.059 | 0.111 | 0.065 | 0.009 | 0.105 |
Conditional DETR | 0.236 | 0.183 | 0.129 | 0.183 | 0.199 | 0.205 | 0.172 | 0.211 | 0.154 | 0.212 |
DDQ | 0.232 | 0.164 | 0.133 | 0.152 | 0.185 | 0.2 | 0.189 | 0.182 | 0.136 | 0.187 |
DAB-DETR | 0.239 | 0.115 | 0.057 | 0.09 | 0.178 | 0.145 | 0.143 | 0.157 | 0.05 | 0.143 |
Deformable DETR | 0.229 | 0.187 | 0.106 | 0.147 | 0.177 | 0.197 | 0.136 | 0.188 | 0.158 | 0.171 |
DINO | 0.232 | 0.11 | 0.075 | 0.104 | 0.172 | 0.18 | 0.139 | 0.161 | 0.078 | 0.156 |
PVT | 0.229 | 0.229 | 0.106 | 0.213 | 0.224 | 0.254 | 0.177 | 0.26 | 0.199 | 0.163 |
PVTv2 | 0.24 | 0.214 | 0.154 | 0.252 | 0.195 | 0.275 | 0.147 | 0.244 | 0.237 | 0.149 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.973 | 0.808 | 0.518 | 0.84 | 0.883 | 0.811 | 0.732 | 0.924 | 0.549 | 0.826 |
AutoAssign | 0.946 | 0.846 | 0.703 | 0.876 | 0.849 | 0.888 | 0.763 | 0.938 | 0.856 | 0.901 |
CenterNet | 0.976 | 0.926 | 0.674 | 0.893 | 0.901 | 0.897 | 0.806 | 0.97 | 0.873 | 0.895 |
CentripetalNet | 0.943 | 0.867 | 0.753 | 0.908 | 0.901 | 0.918 | 0.749 | 0.958 | 0.81 | 0.853 |
CornerNet | 0.948 | 0.8 | 0.692 | 0.905 | 0.847 | 0.902 | 0.697 | 0.959 | 0.722 | 0.829 |
DDOD | 0.95 | 0.936 | 0.756 | 0.923 | 0.9 | 0.934 | 0.863 | 0.948 | 0.949 | 0.909 |
DyHead | 0.977 | 0.836 | 0.606 | 0.792 | 0.903 | 0.845 | 0.838 | 0.843 | 0.685 | 0.927 |
EfficientNet | 0.977 | 0.974 | 0.912 | 0.975 | 0.974 | 0.97 | 0.951 | 0.973 | 0.966 | 0.948 |
FCOS | 0.981 | 0.974 | 0.93 | 0.968 | 0.951 | 0.975 | 0.931 | 0.972 | 0.939 | 0.959 |
FoveaBox | 0.955 | 0.873 | 0.623 | 0.806 | 0.896 | 0.832 | 0.748 | 0.881 | 0.768 | 0.878 |
FreeAnchor | 0.973 | 0.841 | 0.855 | 0.928 | 0.864 | 0.939 | 0.781 | 0.89 | 0.68 | 0.887 |
FSAF | 0.95 | 0.836 | 0.567 | 0.886 | 0.823 | 0.84 | 0.632 | 0.918 | 0.659 | 0.8 |
GFL | 0.978 | 0.837 | 0.575 | 0.81 | 0.908 | 0.839 | 0.836 | 0.913 | 0.549 | 0.903 |
LD | 0.98 | 0.893 | 0.579 | 0.862 | 0.927 | 0.871 | 0.743 | 0.917 | 0.676 | 0.91 |
NAS-FPN | 0.975 | 0.916 | 0.768 | 0.932 | 0.944 | 0.946 | 0.817 | 0.937 | 0.766 | 0.925 |
PAA | 0.966 | 0.923 | 0.861 | 0.952 | 0.905 | 0.937 | 0.895 | 0.938 | 0.951 | 0.92 |
RetinaNet | 0.98 | 0.859 | 0.764 | 0.915 | 0.934 | 0.89 | 0.775 | 0.921 | 0.729 | 0.9 |
RTMDet | 0.982 | 0.954 | 0.943 | 0.971 | 0.958 | 0.971 | 0.975 | 0.98 | 0.99 | 0.937 |
TOOD | 0.908 | 0.763 | 0.666 | 0.83 | 0.834 | 0.826 | 0.717 | 0.847 | 0.622 | 0.853 |
VarifocalNet | 0.977 | 0.802 | 0.486 | 0.77 | 0.9 | 0.832 | 0.671 | 0.866 | 0.466 | 0.829 |
YOLOv5 | 0.975 | 0.979 | 0.967 | 0.974 | 0.977 | 0.974 | 0.974 | 0.972 | 0.985 | 0.966 |
YOLOv6 | 0.985 | 0.982 | 0.968 | 0.974 | 0.983 | 0.979 | 0.974 | 0.984 | 0.978 | 0.973 |
YOLOv7 | 0.962 | 0.924 | 0.931 | 0.948 | 0.939 | 0.958 | 0.932 | 0.96 | 0.932 | 0.931 |
YOLOv8 | 0.975 | 0.919 | 0.903 | 0.93 | 0.942 | 0.96 | 0.916 | 0.965 | 0.885 | 0.917 |
YOLOX | 0.955 | 0.877 | 0.853 | 0.902 | 0.91 | 0.945 | 0.926 | 0.918 | 0.868 | 0.891 |
Faster R-CNN | 0.846 | 0.479 | 0.268 | 0.493 | 0.595 | 0.593 | 0.417 | 0.752 | 0.337 | 0.532 |
Cascade R-CNN | 0.854 | 0.539 | 0.382 | 0.591 | 0.65 | 0.689 | 0.448 | 0.78 | 0.368 | 0.602 |
Cascade RPN | 0.973 | 0.884 | 0.741 | 0.933 | 0.898 | 0.957 | 0.885 | 0.967 | 0.868 | 0.891 |
Double Heads | 0.849 | 0.597 | 0.416 | 0.654 | 0.621 | 0.725 | 0.459 | 0.819 | 0.488 | 0.594 |
FPG | 0.912 | 0.763 | 0.624 | 0.848 | 0.761 | 0.866 | 0.587 | 0.907 | 0.61 | 0.748 |
Grid R-CNN | 0.873 | 0.585 | 0.39 | 0.672 | 0.653 | 0.794 | 0.488 | 0.836 | 0.495 | 0.614 |
Guided Anchoring | 0.975 | 0.962 | 0.914 | 0.966 | 0.962 | 0.966 | 0.839 | 0.961 | 0.929 | 0.956 |
HRNet | 0.844 | 0.655 | 0.429 | 0.687 | 0.627 | 0.732 | 0.423 | 0.875 | 0.532 | 0.609 |
Libra R-CNN | 0.959 | 0.828 | 0.599 | 0.824 | 0.892 | 0.775 | 0.805 | 0.92 | 0.563 | 0.858 |
PAFPN | 0.856 | 0.61 | 0.337 | 0.591 | 0.661 | 0.659 | 0.49 | 0.805 | 0.434 | 0.569 |
RepPoints | 0.978 | 0.903 | 0.595 | 0.874 | 0.885 | 0.884 | 0.833 | 0.945 | 0.883 | 0.848 |
Res2Net | 0.911 | 0.692 | 0.541 | 0.757 | 0.763 | 0.793 | 0.511 | 0.852 | 0.646 | 0.761 |
ResNeSt | 0.929 | 0.646 | 0.497 | 0.727 | 0.701 | 0.691 | 0.685 | 0.802 | 0.515 | 0.716 |
SABL | 0.856 | 0.545 | 0.387 | 0.646 | 0.656 | 0.636 | 0.488 | 0.774 | 0.407 | 0.571 |
Sparse R-CNN | 0.959 | 0.877 | 0.513 | 0.759 | 0.913 | 0.746 | 0.733 | 0.882 | 0.605 | 0.871 |
DETR | 0.746 | 0.328 | 0.232 | 0.256 | 0.468 | 0.383 | 0.457 | 0.369 | 0.234 | 0.468 |
Conditional DETR | 0.962 | 0.831 | 0.73 | 0.931 | 0.865 | 0.934 | 0.881 | 0.964 | 0.839 | 0.917 |
DDQ | 0.983 | 0.976 | 0.972 | 0.979 | 0.975 | 0.975 | 0.977 | 0.974 | 0.983 | 0.969 |
DAB-DETR | 0.98 | 0.909 | 0.928 | 0.97 | 0.948 | 0.968 | 0.946 | 0.924 | 0.91 | 0.934 |
Deformable DETR | 0.954 | 0.907 | 0.704 | 0.902 | 0.879 | 0.924 | 0.766 | 0.905 | 0.91 | 0.88 |
DINO | 0.975 | 0.895 | 0.883 | 0.953 | 0.923 | 0.958 | 0.922 | 0.953 | 0.912 | 0.924 |
PVT | 0.948 | 0.953 | 0.827 | 0.936 | 0.957 | 0.956 | 0.901 | 0.963 | 0.954 | 0.886 |
PVTv2 | 0.973 | 0.942 | 0.884 | 0.952 | 0.934 | 0.973 | 0.835 | 0.967 | 0.941 | 0.867 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.374 | 0.318 | 0.206 | 0.341 | 0.342 | 0.321 | 0.296 | 0.371 | 0.219 | 0.325 |
AutoAssign | 0.385 | 0.341 | 0.293 | 0.366 | 0.335 | 0.367 | 0.32 | 0.39 | 0.345 | 0.374 |
CenterNet | 0.396 | 0.365 | 0.275 | 0.373 | 0.353 | 0.373 | 0.342 | 0.408 | 0.344 | 0.361 |
CentripetalNet | 0.378 | 0.362 | 0.313 | 0.386 | 0.365 | 0.384 | 0.31 | 0.404 | 0.321 | 0.35 |
CornerNet | 0.387 | 0.337 | 0.286 | 0.382 | 0.343 | 0.38 | 0.292 | 0.401 | 0.282 | 0.336 |
DDOD | 0.366 | 0.363 | 0.302 | 0.371 | 0.347 | 0.379 | 0.357 | 0.383 | 0.366 | 0.358 |
DyHead | 0.378 | 0.338 | 0.254 | 0.32 | 0.35 | 0.338 | 0.342 | 0.336 | 0.255 | 0.364 |
EfficientNet | 0.387 | 0.397 | 0.377 | 0.395 | 0.393 | 0.394 | 0.399 | 0.402 | 0.406 | 0.38 |
FCOS | 0.401 | 0.4 | 0.382 | 0.399 | 0.389 | 0.405 | 0.389 | 0.418 | 0.37 | 0.397 |
FoveaBox | 0.371 | 0.337 | 0.246 | 0.327 | 0.344 | 0.333 | 0.304 | 0.353 | 0.303 | 0.344 |
FreeAnchor | 0.384 | 0.332 | 0.333 | 0.367 | 0.345 | 0.38 | 0.324 | 0.365 | 0.244 | 0.357 |
FSAF | 0.37 | 0.337 | 0.229 | 0.362 | 0.33 | 0.342 | 0.271 | 0.377 | 0.257 | 0.325 |
GFL | 0.375 | 0.337 | 0.238 | 0.336 | 0.354 | 0.341 | 0.348 | 0.373 | 0.226 | 0.357 |
LD | 0.372 | 0.352 | 0.236 | 0.351 | 0.367 | 0.348 | 0.314 | 0.367 | 0.268 | 0.36 |
NAS-FPN | 0.38 | 0.367 | 0.307 | 0.373 | 0.378 | 0.375 | 0.34 | 0.384 | 0.296 | 0.373 |
PAA | 0.399 | 0.371 | 0.337 | 0.385 | 0.357 | 0.39 | 0.376 | 0.388 | 0.383 | 0.369 |
RetinaNet | 0.392 | 0.343 | 0.298 | 0.371 | 0.378 | 0.367 | 0.34 | 0.39 | 0.301 | 0.364 |
RTMDet | 0.357 | 0.322 | 0.293 | 0.345 | 0.319 | 0.351 | 0.34 | 0.362 | 0.304 | 0.311 |
TOOD | 0.345 | 0.293 | 0.261 | 0.332 | 0.312 | 0.329 | 0.283 | 0.327 | 0.236 | 0.323 |
VarifocalNet | 0.376 | 0.31 | 0.188 | 0.304 | 0.352 | 0.329 | 0.277 | 0.338 | 0.173 | 0.321 |
YOLOv5 | 0.364 | 0.364 | 0.366 | 0.371 | 0.358 | 0.378 | 0.377 | 0.373 | 0.354 | 0.355 |
YOLOv6 | 0.357 | 0.361 | 0.343 | 0.363 | 0.352 | 0.37 | 0.359 | 0.37 | 0.351 | 0.343 |
YOLOv7 | 0.366 | 0.356 | 0.339 | 0.36 | 0.355 | 0.371 | 0.358 | 0.376 | 0.338 | 0.352 |
YOLOv8 | 0.376 | 0.358 | 0.354 | 0.369 | 0.357 | 0.376 | 0.368 | 0.385 | 0.336 | 0.354 |
YOLOX | 0.362 | 0.33 | 0.325 | 0.345 | 0.338 | 0.368 | 0.367 | 0.358 | 0.309 | 0.332 |
Faster R-CNN | 0.303 | 0.187 | 0.099 | 0.194 | 0.228 | 0.236 | 0.163 | 0.302 | 0.134 | 0.203 |
Cascade R-CNN | 0.316 | 0.213 | 0.147 | 0.232 | 0.255 | 0.272 | 0.185 | 0.312 | 0.145 | 0.239 |
Cascade RPN | 0.375 | 0.333 | 0.276 | 0.351 | 0.341 | 0.358 | 0.353 | 0.376 | 0.307 | 0.341 |
Double Heads | 0.32 | 0.239 | 0.165 | 0.27 | 0.249 | 0.299 | 0.2 | 0.333 | 0.193 | 0.238 |
FPG | 0.347 | 0.31 | 0.249 | 0.35 | 0.302 | 0.352 | 0.239 | 0.379 | 0.249 | 0.306 |
Grid R-CNN | 0.326 | 0.23 | 0.155 | 0.273 | 0.259 | 0.312 | 0.205 | 0.33 | 0.188 | 0.244 |
Guided Anchoring | 0.38 | 0.396 | 0.373 | 0.39 | 0.385 | 0.384 | 0.35 | 0.399 | 0.363 | 0.379 |
HRNet | 0.303 | 0.239 | 0.159 | 0.252 | 0.228 | 0.264 | 0.167 | 0.333 | 0.188 | 0.226 |
Libra R-CNN | 0.383 | 0.326 | 0.233 | 0.334 | 0.357 | 0.315 | 0.333 | 0.385 | 0.223 | 0.337 |
PAFPN | 0.304 | 0.226 | 0.127 | 0.228 | 0.249 | 0.254 | 0.185 | 0.304 | 0.167 | 0.214 |
RepPoints | 0.387 | 0.356 | 0.235 | 0.355 | 0.349 | 0.372 | 0.358 | 0.393 | 0.364 | 0.336 |
Res2Net | 0.341 | 0.261 | 0.208 | 0.302 | 0.284 | 0.305 | 0.21 | 0.338 | 0.237 | 0.292 |
ResNeSt | 0.341 | 0.257 | 0.189 | 0.274 | 0.261 | 0.27 | 0.268 | 0.302 | 0.201 | 0.271 |
SABL | 0.318 | 0.213 | 0.151 | 0.26 | 0.253 | 0.257 | 0.204 | 0.303 | 0.16 | 0.226 |
Sparse R-CNN | 0.395 | 0.367 | 0.219 | 0.324 | 0.38 | 0.32 | 0.312 | 0.381 | 0.249 | 0.361 |
DETR | 0.334 | 0.144 | 0.105 | 0.106 | 0.203 | 0.171 | 0.204 | 0.158 | 0.085 | 0.203 |
Conditional DETR | 0.387 | 0.341 | 0.32 | 0.404 | 0.342 | 0.406 | 0.367 | 0.391 | 0.328 | 0.374 |
DDQ | 0.393 | 0.389 | 0.388 | 0.394 | 0.39 | 0.389 | 0.403 | 0.391 | 0.374 | 0.389 |
DAB-DETR | 0.404 | 0.366 | 0.384 | 0.406 | 0.385 | 0.413 | 0.404 | 0.376 | 0.36 | 0.383 |
Deformable DETR | 0.378 | 0.352 | 0.277 | 0.351 | 0.333 | 0.372 | 0.3 | 0.351 | 0.339 | 0.34 |
DINO | 0.388 | 0.353 | 0.353 | 0.39 | 0.364 | 0.388 | 0.372 | 0.384 | 0.346 | 0.37 |
PVT | 0.349 | 0.371 | 0.319 | 0.365 | 0.37 | 0.379 | 0.365 | 0.382 | 0.387 | 0.34 |
PVTv2 | 0.372 | 0.372 | 0.359 | 0.386 | 0.365 | 0.399 | 0.354 | 0.39 | 0.384 | 0.343 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.54 | 0.517 | 0.498 | 0.428 | 0.419 | 0.473 | 0.522 | 0.468 | 0.495 | 0.458 | 0.385 | 0.454 | 0.492 | 0.514 |
AutoAssign | 0.491 | 0.466 | 0.454 | 0.314 | 0.36 | 0.423 | 0.456 | 0.43 | 0.403 | 0.41 | 0.346 | 0.427 | 0.453 | 0.484 |
CenterNet | 0.524 | 0.476 | 0.477 | 0.408 | 0.39 | 0.469 | 0.483 | 0.436 | 0.437 | 0.45 | 0.372 | 0.43 | 0.474 | 0.524 |
CentripetalNet | 0.526 | 0.53 | 0.524 | 0.48 | 0.349 | 0.524 | 0.508 | 0.51 | 0.471 | 0.473 | 0.405 | 0.472 | 0.515 | 0.526 |
CornerNet | 0.517 | 0.51 | 0.505 | 0.403 | 0.295 | 0.488 | 0.494 | 0.449 | 0.444 | 0.42 | 0.345 | 0.414 | 0.48 | 0.506 |
DDOD | 0.481 | 0.48 | 0.448 | 0.359 | 0.416 | 0.421 | 0.47 | 0.445 | 0.453 | 0.433 | 0.329 | 0.409 | 0.442 | 0.45 |
DyHead | 0.474 | 0.483 | 0.485 | 0.406 | 0.402 | 0.464 | 0.501 | 0.454 | 0.473 | 0.433 | 0.4 | 0.433 | 0.467 | 0.474 |
EfficientNet | 0.457 | 0.431 | 0.442 | 0.398 | 0.418 | 0.406 | 0.431 | 0.399 | 0.407 | 0.403 | 0.394 | 0.396 | 0.422 | 0.433 |
FCOS | 0.45 | 0.438 | 0.429 | 0.364 | 0.383 | 0.41 | 0.433 | 0.407 | 0.404 | 0.407 | 0.356 | 0.409 | 0.425 | 0.448 |
FoveaBox | 0.543 | 0.53 | 0.536 | 0.473 | 0.475 | 0.482 | 0.54 | 0.481 | 0.523 | 0.483 | 0.374 | 0.458 | 0.498 | 0.533 |
FreeAnchor | 0.537 | 0.522 | 0.493 | 0.414 | 0.396 | 0.431 | 0.492 | 0.443 | 0.446 | 0.411 | 0.331 | 0.447 | 0.48 | 0.513 |
FSAF | 0.554 | 0.551 | 0.529 | 0.444 | 0.439 | 0.485 | 0.538 | 0.496 | 0.515 | 0.457 | 0.379 | 0.479 | 0.512 | 0.527 |
GFL | 0.57 | 0.541 | 0.532 | 0.431 | 0.453 | 0.495 | 0.509 | 0.478 | 0.495 | 0.48 | 0.398 | 0.459 | 0.52 | 0.547 |
LD | 0.57 | 0.54 | 0.524 | 0.401 | 0.397 | 0.486 | 0.517 | 0.484 | 0.489 | 0.477 | 0.385 | 0.484 | 0.519 | 0.535 |
NAS-FPN | 0.442 | 0.436 | 0.433 | 0.365 | 0.391 | 0.4 | 0.451 | 0.399 | 0.394 | 0.398 | 0.32 | 0.399 | 0.413 | 0.435 |
PAA | 0.464 | 0.464 | 0.457 | 0.402 | 0.389 | 0.43 | 0.451 | 0.432 | 0.447 | 0.402 | 0.322 | 0.409 | 0.441 | 0.463 |
RetinaNet | 0.522 | 0.543 | 0.497 | 0.438 | 0.425 | 0.459 | 0.505 | 0.483 | 0.478 | 0.454 | 0.384 | 0.475 | 0.489 | 0.528 |
RTMDet | 0.533 | 0.482 | 0.515 | 0.52 | 0.466 | 0.46 | 0.495 | 0.437 | 0.472 | 0.47 | 0.459 | 0.449 | 0.466 | 0.5 |
TOOD | 0.474 | 0.5 | 0.475 | 0.384 | 0.376 | 0.453 | 0.503 | 0.453 | 0.486 | 0.441 | 0.371 | 0.435 | 0.457 | 0.475 |
VarifocalNet | 0.492 | 0.505 | 0.481 | 0.387 | 0.395 | 0.443 | 0.504 | 0.444 | 0.469 | 0.445 | 0.368 | 0.436 | 0.47 | 0.506 |
YOLOv5 | 0.481 | 0.46 | 0.472 | 0.448 | 0.403 | 0.453 | 0.484 | 0.454 | 0.485 | 0.418 | 0.35 | 0.42 | 0.452 | 0.459 |
YOLOv6 | 0.467 | 0.445 | 0.461 | 0.456 | 0.435 | 0.438 | 0.444 | 0.446 | 0.459 | 0.437 | 0.423 | 0.432 | 0.444 | 0.449 |
YOLOv7 | 0.463 | 0.438 | 0.48 | 0.446 | 0.343 | 0.401 | 0.438 | 0.403 | 0.457 | 0.402 | 0.372 | 0.366 | 0.429 | 0.437 |
YOLOv8 | 0.434 | 0.421 | 0.431 | 0.432 | 0.402 | 0.415 | 0.421 | 0.416 | 0.429 | 0.416 | 0.405 | 0.409 | 0.415 | 0.417 |
YOLOX | 0.448 | 0.436 | 0.457 | 0.457 | 0.382 | 0.432 | 0.46 | 0.425 | 0.46 | 0.433 | 0.393 | 0.412 | 0.426 | 0.437 |
Faster R-CNN | 0.541 | 0.547 | 0.497 | 0.416 | 0.425 | 0.456 | 0.532 | 0.468 | 0.456 | 0.448 | 0.341 | 0.432 | 0.512 | 0.534 |
Cascade R-CNN | 0.559 | 0.551 | 0.539 | 0.431 | 0.445 | 0.488 | 0.551 | 0.463 | 0.508 | 0.479 | 0.355 | 0.454 | 0.523 | 0.55 |
Cascade RPN | 0.538 | 0.537 | 0.528 | 0.389 | 0.407 | 0.482 | 0.508 | 0.472 | 0.483 | 0.461 | 0.335 | 0.445 | 0.508 | 0.532 |
Double Heads | 0.552 | 0.526 | 0.531 | 0.419 | 0.408 | 0.46 | 0.533 | 0.442 | 0.489 | 0.454 | 0.359 | 0.44 | 0.496 | 0.527 |
FPG | 0.462 | 0.473 | 0.451 | 0.413 | 0.395 | 0.415 | 0.466 | 0.408 | 0.424 | 0.405 | 0.317 | 0.409 | 0.444 | 0.45 |
Grid R-CNN | 0.512 | 0.502 | 0.492 | 0.397 | 0.404 | 0.449 | 0.517 | 0.462 | 0.471 | 0.43 | 0.363 | 0.424 | 0.481 | 0.488 |
Guided Anchoring | 0.497 | 0.537 | 0.504 | 0.427 | 0.396 | 0.47 | 0.525 | 0.479 | 0.454 | 0.449 | 0.375 | 0.452 | 0.494 | 0.502 |
HRNet | 0.498 | 0.489 | 0.495 | 0.457 | 0.404 | 0.453 | 0.489 | 0.47 | 0.442 | 0.472 | 0.419 | 0.437 | 0.443 | 0.497 |
Libra R-CNN | 0.535 | 0.517 | 0.479 | 0.452 | 0.404 | 0.446 | 0.468 | 0.433 | 0.447 | 0.431 | 0.374 | 0.421 | 0.442 | 0.494 |
PAFPN | 0.539 | 0.534 | 0.529 | 0.429 | 0.438 | 0.468 | 0.522 | 0.477 | 0.47 | 0.458 | 0.349 | 0.447 | 0.516 | 0.559 |
RepPoints | 0.572 | 0.559 | 0.53 | 0.434 | 0.475 | 0.478 | 0.535 | 0.47 | 0.504 | 0.455 | 0.38 | 0.451 | 0.525 | 0.56 |
Res2Net | 0.449 | 0.437 | 0.435 | 0.403 | 0.301 | 0.402 | 0.458 | 0.406 | 0.407 | 0.386 | 0.324 | 0.387 | 0.428 | 0.451 |
ResNeSt | 0.443 | 0.455 | 0.409 | 0.396 | 0.396 | 0.405 | 0.432 | 0.385 | 0.364 | 0.374 | 0.358 | 0.374 | 0.416 | 0.451 |
SABL | 0.563 | 0.559 | 0.525 | 0.418 | 0.471 | 0.491 | 0.534 | 0.503 | 0.498 | 0.496 | 0.382 | 0.484 | 0.534 | 0.552 |
Sparse R-CNN | 0.492 | 0.481 | 0.477 | 0.38 | 0.347 | 0.434 | 0.484 | 0.386 | 0.396 | 0.407 | 0.352 | 0.389 | 0.455 | 0.493 |
DETR | 0.553 | 0.497 | 0.481 | 0.337 | 0.318 | 0.467 | 0.49 | 0.466 | 0.432 | 0.473 | 0.343 | 0.444 | 0.475 | 0.51 |
Conditional DETR | 0.535 | 0.497 | 0.453 | 0.378 | 0.351 | 0.424 | 0.449 | 0.401 | 0.44 | 0.416 | 0.294 | 0.412 | 0.422 | 0.485 |
DDQ | 0.449 | 0.454 | 0.452 | 0.377 | 0.388 | 0.424 | 0.451 | 0.426 | 0.408 | 0.422 | 0.34 | 0.421 | 0.439 | 0.451 |
DAB-DETR | 0.441 | 0.429 | 0.428 | 0.344 | 0.36 | 0.371 | 0.407 | 0.357 | 0.373 | 0.374 | 0.276 | 0.336 | 0.392 | 0.44 |
Deformable DETR | 0.475 | 0.462 | 0.475 | 0.345 | 0.389 | 0.421 | 0.46 | 0.442 | 0.418 | 0.424 | 0.286 | 0.416 | 0.471 | 0.45 |
DINO | 0.419 | 0.421 | 0.416 | 0.337 | 0.283 | 0.385 | 0.415 | 0.402 | 0.375 | 0.391 | 0.316 | 0.378 | 0.394 | 0.423 |
PVT | 0.474 | 0.465 | 0.418 | 0.378 | 0.368 | 0.383 | 0.405 | 0.359 | 0.369 | 0.393 | 0.381 | 0.391 | 0.41 | 0.426 |
PVTv2 | 0.51 | 0.431 | 0.41 | 0.403 | 0.4 | 0.395 | 0.414 | 0.389 | 0.41 | 0.384 | 0.347 | 0.382 | 0.392 | 0.436 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.157 | 0.148 | 0.147 | 0.121 | 0.113 | 0.127 | 0.155 | 0.138 | 0.143 | 0.128 | 0.098 | 0.132 | 0.138 | 0.151 |
AutoAssign | 0.149 | 0.137 | 0.13 | 0.082 | 0.098 | 0.12 | 0.136 | 0.126 | 0.113 | 0.114 | 0.089 | 0.123 | 0.131 | 0.143 |
CenterNet | 0.161 | 0.145 | 0.147 | 0.12 | 0.111 | 0.141 | 0.151 | 0.132 | 0.132 | 0.135 | 0.102 | 0.124 | 0.142 | 0.166 |
CentripetalNet | 0.143 | 0.145 | 0.144 | 0.142 | 0.083 | 0.145 | 0.139 | 0.142 | 0.131 | 0.128 | 0.109 | 0.127 | 0.138 | 0.143 |
CornerNet | 0.142 | 0.141 | 0.138 | 0.12 | 0.07 | 0.133 | 0.141 | 0.128 | 0.125 | 0.115 | 0.091 | 0.111 | 0.132 | 0.14 |
DDOD | 0.142 | 0.141 | 0.129 | 0.102 | 0.119 | 0.115 | 0.139 | 0.133 | 0.131 | 0.129 | 0.085 | 0.114 | 0.126 | 0.128 |
DyHead | 0.135 | 0.137 | 0.144 | 0.113 | 0.109 | 0.127 | 0.149 | 0.132 | 0.141 | 0.12 | 0.106 | 0.122 | 0.131 | 0.134 |
EfficientNet | 0.123 | 0.115 | 0.119 | 0.11 | 0.113 | 0.106 | 0.116 | 0.105 | 0.11 | 0.106 | 0.104 | 0.104 | 0.109 | 0.115 |
FCOS | 0.126 | 0.117 | 0.117 | 0.1 | 0.102 | 0.108 | 0.118 | 0.109 | 0.106 | 0.107 | 0.094 | 0.106 | 0.109 | 0.123 |
FoveaBox | 0.173 | 0.163 | 0.166 | 0.135 | 0.133 | 0.141 | 0.162 | 0.141 | 0.155 | 0.136 | 0.098 | 0.131 | 0.15 | 0.161 |
FreeAnchor | 0.165 | 0.155 | 0.147 | 0.122 | 0.106 | 0.121 | 0.149 | 0.134 | 0.132 | 0.108 | 0.083 | 0.127 | 0.14 | 0.152 |
FSAF | 0.175 | 0.174 | 0.166 | 0.135 | 0.13 | 0.142 | 0.169 | 0.155 | 0.158 | 0.131 | 0.097 | 0.144 | 0.159 | 0.167 |
GFL | 0.177 | 0.164 | 0.162 | 0.123 | 0.13 | 0.138 | 0.16 | 0.143 | 0.157 | 0.144 | 0.108 | 0.133 | 0.153 | 0.164 |
LD | 0.179 | 0.164 | 0.162 | 0.114 | 0.109 | 0.143 | 0.162 | 0.147 | 0.148 | 0.142 | 0.101 | 0.145 | 0.157 | 0.162 |
NAS-FPN | 0.127 | 0.119 | 0.122 | 0.107 | 0.11 | 0.106 | 0.133 | 0.112 | 0.111 | 0.106 | 0.076 | 0.11 | 0.111 | 0.122 |
PAA | 0.132 | 0.132 | 0.13 | 0.112 | 0.103 | 0.116 | 0.135 | 0.128 | 0.129 | 0.108 | 0.081 | 0.12 | 0.122 | 0.129 |
RetinaNet | 0.155 | 0.167 | 0.151 | 0.135 | 0.123 | 0.13 | 0.158 | 0.152 | 0.142 | 0.134 | 0.106 | 0.144 | 0.144 | 0.154 |
RTMDet | 0.152 | 0.133 | 0.148 | 0.158 | 0.129 | 0.12 | 0.14 | 0.112 | 0.137 | 0.129 | 0.126 | 0.117 | 0.121 | 0.138 |
TOOD | 0.131 | 0.145 | 0.139 | 0.108 | 0.099 | 0.121 | 0.154 | 0.13 | 0.143 | 0.122 | 0.093 | 0.122 | 0.123 | 0.132 |
VarifocalNet | 0.141 | 0.152 | 0.145 | 0.11 | 0.111 | 0.127 | 0.157 | 0.133 | 0.144 | 0.125 | 0.095 | 0.132 | 0.137 | 0.154 |
YOLOv5 | 0.135 | 0.124 | 0.132 | 0.129 | 0.104 | 0.122 | 0.136 | 0.124 | 0.142 | 0.108 | 0.082 | 0.106 | 0.122 | 0.124 |
YOLOv6 | 0.13 | 0.121 | 0.129 | 0.139 | 0.119 | 0.118 | 0.126 | 0.123 | 0.128 | 0.118 | 0.115 | 0.116 | 0.119 | 0.123 |
YOLOv7 | 0.125 | 0.115 | 0.129 | 0.117 | 0.083 | 0.101 | 0.115 | 0.105 | 0.129 | 0.101 | 0.091 | 0.091 | 0.111 | 0.115 |
YOLOv8 | 0.12 | 0.113 | 0.12 | 0.128 | 0.111 | 0.111 | 0.115 | 0.111 | 0.119 | 0.109 | 0.106 | 0.106 | 0.111 | 0.112 |
YOLOX | 0.13 | 0.122 | 0.129 | 0.138 | 0.106 | 0.117 | 0.129 | 0.12 | 0.137 | 0.121 | 0.103 | 0.112 | 0.117 | 0.123 |
Faster R-CNN | 0.159 | 0.166 | 0.146 | 0.114 | 0.11 | 0.119 | 0.164 | 0.136 | 0.133 | 0.122 | 0.083 | 0.121 | 0.152 | 0.158 |
Cascade R-CNN | 0.171 | 0.166 | 0.165 | 0.12 | 0.122 | 0.141 | 0.17 | 0.136 | 0.15 | 0.136 | 0.09 | 0.131 | 0.154 | 0.165 |
Cascade RPN | 0.163 | 0.161 | 0.153 | 0.105 | 0.109 | 0.138 | 0.153 | 0.14 | 0.142 | 0.13 | 0.082 | 0.129 | 0.151 | 0.162 |
Double Heads | 0.169 | 0.161 | 0.164 | 0.12 | 0.11 | 0.134 | 0.165 | 0.128 | 0.144 | 0.128 | 0.085 | 0.126 | 0.151 | 0.16 |
FPG | 0.133 | 0.136 | 0.132 | 0.121 | 0.109 | 0.112 | 0.141 | 0.115 | 0.124 | 0.105 | 0.073 | 0.112 | 0.123 | 0.126 |
Grid R-CNN | 0.151 | 0.145 | 0.145 | 0.109 | 0.11 | 0.123 | 0.15 | 0.13 | 0.139 | 0.116 | 0.091 | 0.114 | 0.137 | 0.144 |
Guided Anchoring | 0.142 | 0.162 | 0.153 | 0.119 | 0.107 | 0.137 | 0.163 | 0.136 | 0.132 | 0.123 | 0.094 | 0.128 | 0.146 | 0.151 |
HRNet | 0.147 | 0.139 | 0.146 | 0.132 | 0.105 | 0.117 | 0.141 | 0.133 | 0.122 | 0.132 | 0.109 | 0.117 | 0.122 | 0.144 |
Libra R-CNN | 0.155 | 0.145 | 0.136 | 0.133 | 0.119 | 0.119 | 0.134 | 0.123 | 0.126 | 0.121 | 0.102 | 0.115 | 0.12 | 0.135 |
PAFPN | 0.163 | 0.157 | 0.158 | 0.118 | 0.12 | 0.124 | 0.156 | 0.137 | 0.137 | 0.125 | 0.086 | 0.121 | 0.149 | 0.165 |
RepPoints | 0.178 | 0.177 | 0.166 | 0.124 | 0.14 | 0.131 | 0.167 | 0.135 | 0.146 | 0.132 | 0.099 | 0.129 | 0.154 | 0.174 |
Res2Net | 0.123 | 0.115 | 0.122 | 0.112 | 0.075 | 0.106 | 0.132 | 0.113 | 0.115 | 0.1 | 0.086 | 0.099 | 0.111 | 0.12 |
ResNeSt | 0.125 | 0.132 | 0.118 | 0.11 | 0.11 | 0.109 | 0.125 | 0.108 | 0.102 | 0.101 | 0.091 | 0.1 | 0.118 | 0.127 |
SABL | 0.177 | 0.169 | 0.166 | 0.118 | 0.133 | 0.145 | 0.166 | 0.151 | 0.147 | 0.143 | 0.096 | 0.141 | 0.161 | 0.172 |
Sparse R-CNN | 0.135 | 0.134 | 0.138 | 0.104 | 0.092 | 0.114 | 0.141 | 0.106 | 0.111 | 0.109 | 0.095 | 0.103 | 0.122 | 0.131 |
DETR | 0.171 | 0.151 | 0.151 | 0.101 | 0.088 | 0.142 | 0.152 | 0.146 | 0.126 | 0.138 | 0.095 | 0.128 | 0.142 | 0.151 |
Conditional DETR | 0.157 | 0.138 | 0.128 | 0.1 | 0.09 | 0.114 | 0.129 | 0.112 | 0.123 | 0.108 | 0.068 | 0.105 | 0.111 | 0.135 |
DDQ | 0.119 | 0.119 | 0.124 | 0.099 | 0.099 | 0.112 | 0.124 | 0.116 | 0.113 | 0.11 | 0.085 | 0.113 | 0.117 | 0.12 |
DAB-DETR | 0.118 | 0.111 | 0.117 | 0.092 | 0.094 | 0.099 | 0.111 | 0.095 | 0.101 | 0.098 | 0.071 | 0.087 | 0.103 | 0.117 |
Deformable DETR | 0.136 | 0.129 | 0.142 | 0.093 | 0.107 | 0.118 | 0.134 | 0.128 | 0.12 | 0.12 | 0.073 | 0.114 | 0.137 | 0.121 |
DINO | 0.112 | 0.11 | 0.112 | 0.086 | 0.068 | 0.097 | 0.113 | 0.106 | 0.098 | 0.101 | 0.083 | 0.097 | 0.1 | 0.112 |
PVT | 0.14 | 0.134 | 0.117 | 0.107 | 0.099 | 0.108 | 0.115 | 0.099 | 0.102 | 0.111 | 0.106 | 0.109 | 0.113 | 0.12 |
PVTv2 | 0.156 | 0.132 | 0.119 | 0.119 | 0.112 | 0.108 | 0.126 | 0.118 | 0.122 | 0.11 | 0.097 | 0.108 | 0.111 | 0.129 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.835 | 0.827 | 0.823 | 0.802 | 0.782 | 0.823 | 0.818 | 0.789 | 0.788 | 0.798 | 0.741 | 0.786 | 0.823 | 0.844 |
AutoAssign | 0.854 | 0.837 | 0.841 | 0.794 | 0.827 | 0.832 | 0.826 | 0.814 | 0.817 | 0.827 | 0.793 | 0.811 | 0.832 | 0.851 |
CenterNet | 0.848 | 0.835 | 0.84 | 0.849 | 0.809 | 0.842 | 0.829 | 0.823 | 0.822 | 0.828 | 0.794 | 0.82 | 0.848 | 0.858 |
CentripetalNet | 0.854 | 0.858 | 0.838 | 0.809 | 0.777 | 0.85 | 0.83 | 0.86 | 0.82 | 0.814 | 0.737 | 0.841 | 0.852 | 0.854 |
CornerNet | 0.871 | 0.87 | 0.853 | 0.814 | 0.778 | 0.853 | 0.841 | 0.848 | 0.826 | 0.805 | 0.74 | 0.817 | 0.857 | 0.858 |
DDOD | 0.737 | 0.729 | 0.739 | 0.723 | 0.746 | 0.716 | 0.732 | 0.713 | 0.725 | 0.72 | 0.678 | 0.694 | 0.72 | 0.742 |
DyHead | 0.725 | 0.731 | 0.747 | 0.702 | 0.7 | 0.723 | 0.745 | 0.711 | 0.711 | 0.725 | 0.724 | 0.716 | 0.727 | 0.744 |
EfficientNet | 0.794 | 0.771 | 0.789 | 0.768 | 0.751 | 0.758 | 0.781 | 0.762 | 0.754 | 0.762 | 0.748 | 0.752 | 0.766 | 0.793 |
FCOS | 0.897 | 0.905 | 0.89 | 0.864 | 0.86 | 0.891 | 0.894 | 0.893 | 0.858 | 0.897 | 0.82 | 0.898 | 0.899 | 0.908 |
FoveaBox | 0.824 | 0.814 | 0.836 | 0.836 | 0.794 | 0.825 | 0.813 | 0.799 | 0.819 | 0.81 | 0.785 | 0.776 | 0.818 | 0.836 |
FreeAnchor | 0.792 | 0.801 | 0.788 | 0.791 | 0.772 | 0.786 | 0.789 | 0.768 | 0.769 | 0.781 | 0.717 | 0.784 | 0.793 | 0.809 |
FSAF | 0.812 | 0.814 | 0.818 | 0.803 | 0.79 | 0.807 | 0.822 | 0.808 | 0.808 | 0.803 | 0.785 | 0.803 | 0.805 | 0.812 |
GFL | 0.827 | 0.803 | 0.805 | 0.796 | 0.782 | 0.806 | 0.784 | 0.781 | 0.763 | 0.785 | 0.743 | 0.777 | 0.804 | 0.823 |
LD | 0.809 | 0.797 | 0.802 | 0.785 | 0.762 | 0.787 | 0.786 | 0.764 | 0.77 | 0.777 | 0.735 | 0.763 | 0.797 | 0.813 |
NAS-FPN | 0.773 | 0.765 | 0.775 | 0.726 | 0.75 | 0.74 | 0.768 | 0.744 | 0.732 | 0.741 | 0.7 | 0.751 | 0.76 | 0.778 |
PAA | 0.816 | 0.819 | 0.821 | 0.808 | 0.791 | 0.805 | 0.805 | 0.792 | 0.784 | 0.78 | 0.73 | 0.783 | 0.807 | 0.837 |
RetinaNet | 0.838 | 0.823 | 0.843 | 0.817 | 0.795 | 0.823 | 0.822 | 0.802 | 0.796 | 0.818 | 0.777 | 0.791 | 0.823 | 0.848 |
RTMDet | 0.976 | 0.978 | 0.974 | 0.973 | 0.972 | 0.975 | 0.974 | 0.966 | 0.971 | 0.97 | 0.966 | 0.976 | 0.977 | 0.976 |
TOOD | 0.73 | 0.732 | 0.731 | 0.708 | 0.691 | 0.734 | 0.737 | 0.716 | 0.723 | 0.724 | 0.69 | 0.714 | 0.723 | 0.733 |
VarifocalNet | 0.794 | 0.772 | 0.786 | 0.767 | 0.738 | 0.765 | 0.77 | 0.744 | 0.755 | 0.758 | 0.729 | 0.745 | 0.771 | 0.791 |
YOLOv5 | 0.814 | 0.832 | 0.836 | 0.773 | 0.782 | 0.825 | 0.831 | 0.841 | 0.814 | 0.829 | 0.817 | 0.836 | 0.831 | 0.843 |
YOLOv6 | 0.96 | 0.963 | 0.957 | 0.951 | 0.939 | 0.959 | 0.943 | 0.964 | 0.943 | 0.95 | 0.921 | 0.956 | 0.963 | 0.966 |
YOLOv7 | 0.729 | 0.73 | 0.744 | 0.723 | 0.7 | 0.716 | 0.731 | 0.712 | 0.721 | 0.731 | 0.721 | 0.723 | 0.722 | 0.743 |
YOLOv8 | 0.677 | 0.667 | 0.686 | 0.67 | 0.669 | 0.676 | 0.676 | 0.679 | 0.687 | 0.674 | 0.675 | 0.674 | 0.668 | 0.679 |
YOLOX | 0.685 | 0.697 | 0.702 | 0.702 | 0.702 | 0.7 | 0.708 | 0.706 | 0.709 | 0.707 | 0.731 | 0.71 | 0.682 | 0.692 |
Faster R-CNN | 0.69 | 0.682 | 0.691 | 0.625 | 0.659 | 0.679 | 0.674 | 0.653 | 0.645 | 0.679 | 0.615 | 0.639 | 0.67 | 0.685 |
Cascade R-CNN | 0.699 | 0.689 | 0.703 | 0.644 | 0.652 | 0.679 | 0.695 | 0.66 | 0.673 | 0.677 | 0.623 | 0.651 | 0.68 | 0.7 |
Cascade RPN | 0.785 | 0.791 | 0.789 | 0.759 | 0.758 | 0.78 | 0.776 | 0.748 | 0.751 | 0.774 | 0.715 | 0.758 | 0.784 | 0.797 |
Double Heads | 0.696 | 0.688 | 0.7 | 0.636 | 0.639 | 0.677 | 0.678 | 0.664 | 0.654 | 0.679 | 0.623 | 0.643 | 0.676 | 0.696 |
FPG | 0.656 | 0.659 | 0.665 | 0.598 | 0.598 | 0.643 | 0.659 | 0.636 | 0.622 | 0.642 | 0.588 | 0.635 | 0.659 | 0.661 |
Grid R-CNN | 0.677 | 0.664 | 0.67 | 0.638 | 0.635 | 0.666 | 0.669 | 0.654 | 0.655 | 0.659 | 0.632 | 0.644 | 0.657 | 0.661 |
Guided Anchoring | 0.74 | 0.732 | 0.748 | 0.729 | 0.709 | 0.738 | 0.745 | 0.733 | 0.711 | 0.736 | 0.72 | 0.718 | 0.739 | 0.739 |
HRNet | 0.665 | 0.669 | 0.679 | 0.645 | 0.644 | 0.671 | 0.673 | 0.661 | 0.646 | 0.676 | 0.65 | 0.645 | 0.649 | 0.678 |
Libra R-CNN | 0.829 | 0.813 | 0.831 | 0.822 | 0.766 | 0.81 | 0.811 | 0.781 | 0.776 | 0.797 | 0.754 | 0.778 | 0.819 | 0.848 |
PAFPN | 0.685 | 0.682 | 0.697 | 0.644 | 0.647 | 0.678 | 0.679 | 0.664 | 0.658 | 0.682 | 0.614 | 0.651 | 0.679 | 0.698 |
RepPoints | 0.826 | 0.813 | 0.822 | 0.816 | 0.813 | 0.804 | 0.807 | 0.799 | 0.808 | 0.801 | 0.783 | 0.779 | 0.813 | 0.828 |
Res2Net | 0.636 | 0.63 | 0.638 | 0.599 | 0.543 | 0.627 | 0.623 | 0.614 | 0.602 | 0.623 | 0.578 | 0.618 | 0.63 | 0.634 |
ResNeSt | 0.648 | 0.651 | 0.625 | 0.61 | 0.61 | 0.636 | 0.619 | 0.614 | 0.586 | 0.607 | 0.593 | 0.609 | 0.64 | 0.654 |
SABL | 0.711 | 0.704 | 0.707 | 0.65 | 0.66 | 0.69 | 0.678 | 0.675 | 0.675 | 0.688 | 0.635 | 0.661 | 0.696 | 0.706 |
Sparse R-CNN | 0.702 | 0.685 | 0.694 | 0.67 | 0.668 | 0.672 | 0.688 | 0.65 | 0.644 | 0.667 | 0.649 | 0.651 | 0.68 | 0.694 |
DETR | 0.893 | 0.9 | 0.877 | 0.761 | 0.798 | 0.904 | 0.887 | 0.891 | 0.827 | 0.882 | 0.789 | 0.889 | 0.895 | 0.892 |
Conditional DETR | 0.73 | 0.729 | 0.717 | 0.706 | 0.708 | 0.697 | 0.703 | 0.67 | 0.691 | 0.683 | 0.643 | 0.692 | 0.69 | 0.712 |
DDQ | 0.774 | 0.774 | 0.768 | 0.768 | 0.758 | 0.746 | 0.757 | 0.728 | 0.744 | 0.748 | 0.725 | 0.745 | 0.746 | 0.773 |
DAB-DETR | 0.743 | 0.739 | 0.745 | 0.737 | 0.76 | 0.729 | 0.742 | 0.705 | 0.73 | 0.717 | 0.703 | 0.716 | 0.722 | 0.743 |
Deformable DETR | 0.703 | 0.699 | 0.707 | 0.68 | 0.693 | 0.671 | 0.692 | 0.666 | 0.675 | 0.673 | 0.591 | 0.67 | 0.679 | 0.693 |
DINO | 0.739 | 0.748 | 0.744 | 0.735 | 0.747 | 0.732 | 0.747 | 0.723 | 0.72 | 0.727 | 0.718 | 0.734 | 0.73 | 0.736 |
PVT | 0.703 | 0.705 | 0.7 | 0.692 | 0.673 | 0.693 | 0.696 | 0.679 | 0.669 | 0.692 | 0.681 | 0.691 | 0.688 | 0.699 |
PVTv2 | 0.738 | 0.733 | 0.714 | 0.736 | 0.731 | 0.712 | 0.733 | 0.715 | 0.705 | 0.725 | 0.7 | 0.716 | 0.709 | 0.731 |
Clean |
Random |
AdvCam |
UPC |
NatPatch |
MTD |
LAP |
InvisCloak |
DAP |
AdvTshirt |
AdvTexture |
AdvPatch |
AdvPattern |
AdvCaT |
ATSS | 0.326 | 0.313 | 0.318 | 0.301 | 0.298 | 0.304 | 0.306 | 0.291 | 0.302 | 0.297 | 0.269 | 0.286 | 0.307 | 0.322 |
AutoAssign | 0.317 | 0.309 | 0.309 | 0.288 | 0.306 | 0.299 | 0.305 | 0.297 | 0.3 | 0.301 | 0.282 | 0.297 | 0.306 | 0.314 |
CenterNet | 0.351 | 0.344 | 0.342 | 0.341 | 0.315 | 0.342 | 0.332 | 0.326 | 0.325 | 0.332 | 0.309 | 0.329 | 0.346 | 0.352 |
CentripetalNet | 0.332 | 0.337 | 0.326 | 0.32 | 0.293 | 0.333 | 0.319 | 0.338 | 0.311 | 0.307 | 0.277 | 0.321 | 0.331 | 0.332 |
CornerNet | 0.331 | 0.333 | 0.325 | 0.306 | 0.281 | 0.322 | 0.314 | 0.314 | 0.306 | 0.297 | 0.27 | 0.298 | 0.321 | 0.326 |
DDOD | 0.273 | 0.267 | 0.275 | 0.262 | 0.28 | 0.256 | 0.267 | 0.256 | 0.264 | 0.262 | 0.245 | 0.251 | 0.262 | 0.273 |
DyHead | 0.283 | 0.284 | 0.294 | 0.267 | 0.273 | 0.275 | 0.293 | 0.273 | 0.279 | 0.279 | 0.27 | 0.274 | 0.279 | 0.291 |
EfficientNet | 0.29 | 0.283 | 0.295 | 0.283 | 0.284 | 0.277 | 0.291 | 0.278 | 0.284 | 0.28 | 0.274 | 0.272 | 0.276 | 0.291 |
FCOS | 0.339 | 0.341 | 0.341 | 0.323 | 0.331 | 0.331 | 0.339 | 0.337 | 0.322 | 0.337 | 0.296 | 0.335 | 0.333 | 0.344 |
FoveaBox | 0.304 | 0.298 | 0.316 | 0.309 | 0.29 | 0.295 | 0.293 | 0.285 | 0.301 | 0.292 | 0.276 | 0.279 | 0.294 | 0.309 |
FreeAnchor | 0.289 | 0.289 | 0.29 | 0.283 | 0.278 | 0.28 | 0.284 | 0.276 | 0.278 | 0.277 | 0.249 | 0.28 | 0.284 | 0.292 |
FSAF | 0.307 | 0.305 | 0.31 | 0.298 | 0.304 | 0.299 | 0.307 | 0.299 | 0.307 | 0.294 | 0.279 | 0.297 | 0.301 | 0.305 |
GFL | 0.311 | 0.298 | 0.304 | 0.29 | 0.29 | 0.292 | 0.29 | 0.283 | 0.285 | 0.284 | 0.267 | 0.281 | 0.295 | 0.308 |
LD | 0.302 | 0.293 | 0.298 | 0.283 | 0.28 | 0.285 | 0.291 | 0.278 | 0.286 | 0.283 | 0.263 | 0.276 | 0.29 | 0.297 |
NAS-FPN | 0.294 | 0.28 | 0.293 | 0.265 | 0.278 | 0.27 | 0.285 | 0.271 | 0.272 | 0.271 | 0.242 | 0.272 | 0.278 | 0.288 |
PAA | 0.321 | 0.322 | 0.323 | 0.308 | 0.303 | 0.312 | 0.32 | 0.312 | 0.313 | 0.303 | 0.269 | 0.304 | 0.318 | 0.33 |
RetinaNet | 0.317 | 0.32 | 0.327 | 0.314 | 0.305 | 0.312 | 0.316 | 0.309 | 0.307 | 0.312 | 0.295 | 0.304 | 0.312 | 0.327 |
RTMDet | 0.246 | 0.242 | 0.248 | 0.235 | 0.233 | 0.236 | 0.244 | 0.229 | 0.237 | 0.238 | 0.233 | 0.233 | 0.234 | 0.238 |
TOOD | 0.277 | 0.276 | 0.278 | 0.262 | 0.258 | 0.271 | 0.277 | 0.266 | 0.274 | 0.269 | 0.249 | 0.265 | 0.268 | 0.278 |
VarifocalNet | 0.295 | 0.287 | 0.296 | 0.281 | 0.277 | 0.28 | 0.286 | 0.271 | 0.28 | 0.277 | 0.262 | 0.274 | 0.284 | 0.294 |
YOLOv5 | 0.241 | 0.239 | 0.243 | 0.235 | 0.222 | 0.237 | 0.242 | 0.238 | 0.241 | 0.229 | 0.179 | 0.223 | 0.238 | 0.235 |
YOLOv6 | 0.241 | 0.237 | 0.24 | 0.232 | 0.233 | 0.231 | 0.236 | 0.232 | 0.238 | 0.231 | 0.228 | 0.228 | 0.232 | 0.233 |
YOLOv7 | 0.242 | 0.236 | 0.242 | 0.229 | 0.207 | 0.221 | 0.237 | 0.218 | 0.235 | 0.226 | 0.215 | 0.207 | 0.232 | 0.233 |
YOLOv8 | 0.244 | 0.241 | 0.244 | 0.235 | 0.235 | 0.237 | 0.241 | 0.238 | 0.242 | 0.237 | 0.232 | 0.233 | 0.239 | 0.238 |
YOLOX | 0.242 | 0.241 | 0.243 | 0.233 | 0.217 | 0.237 | 0.241 | 0.237 | 0.239 | 0.236 | 0.227 | 0.234 | 0.239 | 0.237 |
Faster R-CNN | 0.256 | 0.255 | 0.261 | 0.228 | 0.242 | 0.246 | 0.253 | 0.24 | 0.244 | 0.25 | 0.212 | 0.233 | 0.247 | 0.255 |
Cascade R-CNN | 0.262 | 0.255 | 0.267 | 0.238 | 0.246 | 0.248 | 0.261 | 0.244 | 0.256 | 0.249 | 0.224 | 0.24 | 0.249 | 0.258 |
Cascade RPN | 0.279 | 0.278 | 0.28 | 0.26 | 0.266 | 0.27 | 0.274 | 0.261 | 0.266 | 0.268 | 0.241 | 0.262 | 0.273 | 0.281 |
Double Heads | 0.26 | 0.255 | 0.262 | 0.233 | 0.232 | 0.243 | 0.253 | 0.24 | 0.245 | 0.244 | 0.209 | 0.233 | 0.249 | 0.258 |
FPG | 0.248 | 0.243 | 0.251 | 0.217 | 0.211 | 0.233 | 0.247 | 0.235 | 0.235 | 0.231 | 0.196 | 0.229 | 0.242 | 0.241 |
Grid R-CNN | 0.25 | 0.244 | 0.251 | 0.235 | 0.24 | 0.243 | 0.25 | 0.241 | 0.248 | 0.246 | 0.234 | 0.238 | 0.24 | 0.244 |
Guided Anchoring | 0.271 | 0.266 | 0.271 | 0.254 | 0.252 | 0.262 | 0.272 | 0.259 | 0.257 | 0.265 | 0.244 | 0.253 | 0.264 | 0.268 |
HRNet | 0.253 | 0.248 | 0.256 | 0.243 | 0.233 | 0.246 | 0.25 | 0.246 | 0.243 | 0.252 | 0.239 | 0.234 | 0.238 | 0.254 |
Libra R-CNN | 0.321 | 0.314 | 0.319 | 0.319 | 0.293 | 0.302 | 0.307 | 0.294 | 0.3 | 0.303 | 0.285 | 0.291 | 0.309 | 0.326 |
PAFPN | 0.257 | 0.253 | 0.263 | 0.232 | 0.243 | 0.245 | 0.255 | 0.243 | 0.251 | 0.249 | 0.212 | 0.235 | 0.25 | 0.259 |
RepPoints | 0.314 | 0.31 | 0.319 | 0.304 | 0.306 | 0.295 | 0.301 | 0.291 | 0.305 | 0.298 | 0.28 | 0.287 | 0.306 | 0.316 |
Res2Net | 0.237 | 0.231 | 0.24 | 0.22 | 0.193 | 0.229 | 0.234 | 0.227 | 0.226 | 0.228 | 0.216 | 0.222 | 0.23 | 0.232 |
ResNeSt | 0.244 | 0.248 | 0.236 | 0.224 | 0.226 | 0.236 | 0.236 | 0.231 | 0.217 | 0.226 | 0.214 | 0.221 | 0.241 | 0.245 |
SABL | 0.271 | 0.26 | 0.27 | 0.24 | 0.251 | 0.254 | 0.256 | 0.247 | 0.257 | 0.252 | 0.228 | 0.245 | 0.26 | 0.265 |
Sparse R-CNN | 0.268 | 0.262 | 0.267 | 0.25 | 0.253 | 0.251 | 0.263 | 0.246 | 0.247 | 0.253 | 0.247 | 0.244 | 0.259 | 0.265 |
DETR | 0.42 | 0.42 | 0.391 | 0.334 | 0.352 | 0.399 | 0.39 | 0.399 | 0.351 | 0.383 | 0.328 | 0.392 | 0.404 | 0.397 |
Conditional DETR | 0.293 | 0.302 | 0.29 | 0.278 | 0.281 | 0.285 | 0.28 | 0.268 | 0.269 | 0.268 | 0.236 | 0.275 | 0.278 | 0.286 |
DDQ | 0.322 | 0.32 | 0.32 | 0.303 | 0.304 | 0.301 | 0.313 | 0.29 | 0.29 | 0.305 | 0.284 | 0.299 | 0.3 | 0.322 |
DAB-DETR | 0.301 | 0.301 | 0.303 | 0.3 | 0.31 | 0.295 | 0.296 | 0.277 | 0.294 | 0.285 | 0.27 | 0.28 | 0.292 | 0.303 |
Deformable DETR | 0.281 | 0.271 | 0.278 | 0.258 | 0.282 | 0.256 | 0.27 | 0.252 | 0.261 | 0.26 | 0.221 | 0.257 | 0.261 | 0.267 |
DINO | 0.303 | 0.309 | 0.303 | 0.29 | 0.294 | 0.294 | 0.302 | 0.287 | 0.276 | 0.293 | 0.276 | 0.293 | 0.295 | 0.299 |
PVT | 0.273 | 0.275 | 0.27 | 0.264 | 0.253 | 0.268 | 0.271 | 0.26 | 0.254 | 0.271 | 0.266 | 0.267 | 0.266 | 0.275 |
PVTv2 | 0.284 | 0.283 | 0.277 | 0.278 | 0.277 | 0.271 | 0.281 | 0.276 | 0.266 | 0.279 | 0.264 | 0.273 | 0.273 | 0.28 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.975 | 0.963 | 0.863 | 0.963 | 0.95 | 1.0 | 0.863 | 0.963 | 0.887 | 0.875 |
AutoAssign | 0.975 | 1.0 | 0.975 | 1.0 | 1.0 | 1.0 | 0.975 | 0.988 | 1.0 | 0.988 |
CenterNet | 0.95 | 1.0 | 0.963 | 1.0 | 0.975 | 1.0 | 0.887 | 1.0 | 1.0 | 0.925 |
CentripetalNet | 0.975 | 1.0 | 0.938 | 1.0 | 0.95 | 1.0 | 0.975 | 1.0 | 0.988 | 0.988 |
CornerNet | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.95 |
DDOD | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 |
DyHead | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 0.975 | 1.0 |
EfficientNet | 0.938 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 0.975 |
FCOS | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 0.975 |
FoveaBox | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
FreeAnchor | 1.0 | 0.863 | 0.812 | 0.975 | 0.863 | 1.0 | 0.8 | 0.875 | 0.875 | 0.887 |
FSAF | 0.988 | 0.887 | 0.887 | 0.988 | 0.925 | 1.0 | 0.925 | 0.975 | 0.9 | 0.912 |
GFL | 0.988 | 0.875 | 0.875 | 1.0 | 0.912 | 1.0 | 0.887 | 0.875 | 0.875 | 0.875 |
LD | 1.0 | 0.9 | 0.887 | 1.0 | 0.975 | 1.0 | 0.887 | 1.0 | 0.9 | 0.975 |
NAS-FPN | 1.0 | 0.912 | 1.0 | 1.0 | 0.963 | 1.0 | 0.975 | 1.0 | 0.975 | 0.963 |
PAA | 0.988 | 1.0 | 0.988 | 1.0 | 0.975 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
RetinaNet | 0.988 | 0.938 | 0.925 | 0.963 | 0.925 | 1.0 | 0.863 | 0.988 | 0.863 | 0.925 |
RTMDet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 |
TOOD | 1.0 | 1.0 | 0.912 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 |
VarifocalNet | 0.988 | 0.938 | 0.85 | 0.887 | 0.95 | 1.0 | 0.925 | 0.85 | 0.875 | 0.887 |
YOLOv5 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.963 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.925 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.975 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOX | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 |
Faster R-CNN | 0.988 | 0.85 | 0.562 | 0.775 | 0.85 | 0.938 | 0.762 | 0.887 | 0.725 | 0.875 |
Cascade R-CNN | 1.0 | 0.863 | 0.738 | 0.875 | 0.912 | 0.975 | 0.787 | 0.925 | 0.787 | 0.863 |
Cascade RPN | 1.0 | 1.0 | 0.988 | 1.0 | 0.988 | 0.975 | 0.988 | 1.0 | 0.988 | 0.988 |
Double Heads | 0.975 | 0.9 | 0.812 | 0.875 | 0.825 | 1.0 | 0.838 | 0.975 | 0.838 | 0.875 |
FPG | 0.988 | 0.988 | 0.95 | 1.0 | 0.988 | 1.0 | 0.9 | 0.988 | 0.938 | 0.925 |
Grid R-CNN | 1.0 | 0.863 | 0.738 | 0.963 | 0.85 | 0.988 | 0.838 | 1.0 | 0.875 | 0.875 |
Guided Anchoring | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 1.0 | 0.975 |
HRNet | 0.938 | 0.875 | 0.938 | 0.95 | 0.875 | 0.963 | 0.775 | 0.9 | 0.95 | 0.925 |
Libra R-CNN | 0.988 | 0.938 | 0.938 | 0.963 | 0.963 | 1.0 | 0.975 | 1.0 | 0.9 | 0.975 |
PAFPN | 0.988 | 0.863 | 0.637 | 0.825 | 0.887 | 0.963 | 0.838 | 0.963 | 0.825 | 0.863 |
RepPoints | 0.988 | 0.975 | 0.875 | 1.0 | 0.9 | 1.0 | 0.975 | 0.975 | 0.975 | 0.912 |
Res2Net | 0.975 | 1.0 | 0.925 | 0.988 | 0.988 | 1.0 | 0.912 | 0.938 | 0.988 | 1.0 |
ResNeSt | 0.975 | 0.863 | 0.825 | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 0.863 | 0.963 |
SABL | 1.0 | 0.875 | 0.775 | 0.938 | 0.887 | 0.988 | 0.787 | 0.95 | 0.85 | 0.863 |
Sparse R-CNN | 0.975 | 0.912 | 0.85 | 1.0 | 0.988 | 0.988 | 0.95 | 0.95 | 0.838 | 0.9 |
DETR | 0.95 | 0.537 | 0.388 | 0.237 | 0.8 | 0.713 | 0.812 | 0.5 | 0.487 | 0.8 |
Conditional DETR | 0.988 | 0.887 | 0.975 | 0.975 | 0.975 | 1.0 | 1.0 | 1.0 | 0.875 | 0.912 |
DDQ | 0.988 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.975 | 1.0 |
Deformable DETR | 0.963 | 0.975 | 0.9 | 0.938 | 0.95 | 0.975 | 0.9 | 0.975 | 0.912 | 0.938 |
DINO | 1.0 | 1.0 | 0.988 | 0.925 | 0.975 | 0.988 | 1.0 | 1.0 | 0.9 | 1.0 |
PVT | 0.988 | 1.0 | 0.75 | 0.988 | 1.0 | 1.0 | 0.988 | 1.0 | 1.0 | 0.925 |
PVTv2 | 1.0 | 0.988 | 0.988 | 1.0 | 1.0 | 1.0 | 0.887 | 1.0 | 1.0 | 0.912 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 1.0 | 0.979 | 0.885 | 0.99 | 0.969 | 0.99 | 0.844 | 0.969 | 0.875 | 0.875 |
AutoAssign | 0.958 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 |
CenterNet | 0.99 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 | 0.885 | 0.979 | 0.99 | 0.938 |
CentripetalNet | 1.0 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 |
CornerNet | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.969 | 0.99 |
DDOD | 1.0 | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
DyHead | 1.0 | 1.0 | 0.979 | 1.0 | 0.99 | 1.0 | 1.0 | 0.979 | 0.969 | 1.0 |
EfficientNet | 0.958 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 |
FCOS | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.958 |
FoveaBox | 0.99 | 0.99 | 0.875 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
FreeAnchor | 0.99 | 0.865 | 0.833 | 0.979 | 0.885 | 0.99 | 0.792 | 0.938 | 0.885 | 0.865 |
FSAF | 1.0 | 0.906 | 0.812 | 0.99 | 0.938 | 0.99 | 0.865 | 0.938 | 0.875 | 0.885 |
GFL | 0.99 | 0.927 | 0.969 | 0.99 | 0.938 | 0.99 | 0.885 | 0.927 | 0.917 | 0.896 |
LD | 0.99 | 0.948 | 0.896 | 0.99 | 0.958 | 0.99 | 0.823 | 0.979 | 0.896 | 0.99 |
NAS-FPN | 1.0 | 0.885 | 0.969 | 0.979 | 0.948 | 1.0 | 0.979 | 1.0 | 0.927 | 0.917 |
PAA | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 |
RetinaNet | 0.99 | 0.865 | 0.865 | 0.917 | 0.948 | 0.979 | 0.854 | 0.917 | 0.802 | 0.875 |
RTMDet | 1.0 | 1.0 | 0.99 | 1.0 | 0.99 | 0.979 | 0.99 | 0.99 | 1.0 | 0.99 |
TOOD | 0.99 | 0.99 | 0.927 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
VarifocalNet | 1.0 | 0.958 | 0.802 | 0.865 | 0.938 | 0.99 | 0.885 | 0.823 | 0.833 | 0.875 |
YOLOv5 | 1.0 | 0.99 | 0.99 | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 0.99 | 0.969 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.979 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 0.99 | 0.99 | 0.917 | 1.0 | 1.0 | 0.99 | 0.99 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.958 | 1.0 | 1.0 | 0.99 | 1.0 |
YOLOX | 1.0 | 0.99 | 1.0 | 0.99 | 0.99 | 0.958 | 0.99 | 1.0 | 0.99 | 0.99 |
Faster R-CNN | 0.99 | 0.865 | 0.74 | 0.969 | 0.885 | 0.99 | 0.792 | 0.948 | 0.833 | 0.885 |
Cascade R-CNN | 0.99 | 0.865 | 0.854 | 0.969 | 0.885 | 0.99 | 0.823 | 0.917 | 0.844 | 0.854 |
Cascade RPN | 0.99 | 0.958 | 0.938 | 0.99 | 0.938 | 0.969 | 0.906 | 0.99 | 0.99 | 0.958 |
Double Heads | 0.99 | 0.896 | 0.927 | 0.969 | 0.823 | 0.99 | 0.865 | 0.99 | 0.844 | 0.865 |
FPG | 1.0 | 0.948 | 0.896 | 0.979 | 0.958 | 0.99 | 0.833 | 0.99 | 0.969 | 0.906 |
Grid R-CNN | 0.979 | 0.906 | 0.885 | 0.979 | 0.917 | 0.979 | 0.875 | 0.99 | 0.917 | 0.896 |
Guided Anchoring | 1.0 | 0.979 | 0.99 | 0.99 | 0.99 | 0.917 | 0.979 | 0.99 | 0.99 | 0.979 |
HRNet | 0.969 | 0.906 | 0.958 | 0.958 | 0.854 | 0.969 | 0.812 | 0.958 | 0.969 | 0.927 |
Libra R-CNN | 1.0 | 0.958 | 0.917 | 0.99 | 0.948 | 0.99 | 0.948 | 0.99 | 0.854 | 0.948 |
PAFPN | 0.99 | 0.875 | 0.74 | 0.938 | 0.938 | 0.969 | 0.927 | 0.979 | 0.844 | 0.844 |
RepPoints | 0.99 | 0.979 | 0.823 | 0.99 | 0.885 | 0.99 | 0.948 | 0.979 | 0.927 | 0.896 |
Res2Net | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.927 | 0.99 | 0.979 | 0.979 |
ResNeSt | 0.958 | 0.865 | 0.938 | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 0.917 | 0.969 |
SABL | 0.99 | 0.917 | 0.854 | 0.99 | 0.875 | 0.99 | 0.833 | 0.99 | 0.865 | 0.865 |
Sparse R-CNN | 0.99 | 0.969 | 0.917 | 0.99 | 0.99 | 0.99 | 0.958 | 0.979 | 0.875 | 0.948 |
DETR | 0.99 | 0.677 | 0.448 | 0.302 | 0.885 | 0.875 | 0.854 | 0.74 | 0.594 | 0.844 |
Conditional DETR | 0.99 | 0.958 | 0.938 | 0.979 | 0.979 | 0.948 | 0.969 | 0.99 | 0.906 | 0.896 |
DDQ | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
Deformable DETR | 1.0 | 0.948 | 0.875 | 0.917 | 0.917 | 0.969 | 0.917 | 0.979 | 0.979 | 0.938 |
DINO | 1.0 | 0.99 | 0.948 | 0.938 | 0.979 | 0.979 | 0.979 | 0.99 | 0.927 | 0.99 |
PVT | 0.938 | 0.979 | 0.625 | 0.979 | 0.969 | 0.99 | 0.927 | 0.99 | 0.99 | 0.812 |
PVTv2 | 1.0 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 | 0.802 | 0.99 | 0.99 | 0.875 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.979 | 0.979 | 0.958 | 0.99 | 0.99 | 0.99 | 0.948 | 0.979 | 0.979 | 0.958 |
AutoAssign | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 | 0.958 | 0.99 | 0.979 | 0.99 |
CenterNet | 0.99 | 0.979 | 1.0 | 0.99 | 0.979 | 1.0 | 0.99 | 1.0 | 1.0 | 0.969 |
CentripetalNet | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
CornerNet | 1.0 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 |
DDOD | 0.99 | 0.979 | 0.958 | 0.99 | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.979 |
DyHead | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.979 | 0.99 |
EfficientNet | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.917 | 0.99 | 0.99 | 0.99 | 0.979 |
FCOS | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.969 |
FoveaBox | 0.99 | 0.99 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
FreeAnchor | 0.979 | 0.917 | 0.865 | 0.979 | 0.927 | 0.99 | 0.802 | 0.969 | 0.802 | 0.917 |
FSAF | 1.0 | 0.917 | 0.885 | 0.99 | 0.958 | 0.979 | 0.896 | 0.99 | 0.917 | 0.906 |
GFL | 0.979 | 0.969 | 0.948 | 0.99 | 0.969 | 0.979 | 0.958 | 0.99 | 0.896 | 0.969 |
LD | 0.979 | 0.969 | 0.969 | 0.99 | 0.979 | 0.99 | 0.958 | 0.99 | 0.969 | 0.979 |
NAS-FPN | 0.99 | 0.969 | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.948 | 0.958 |
PAA | 0.99 | 1.0 | 0.979 | 0.99 | 0.99 | 0.979 | 0.969 | 0.99 | 0.99 | 0.99 |
RetinaNet | 0.979 | 0.917 | 0.917 | 0.969 | 0.979 | 0.979 | 0.875 | 0.979 | 0.875 | 0.927 |
RTMDet | 0.99 | 0.99 | 1.0 | 0.938 | 0.99 | 0.958 | 1.0 | 1.0 | 0.99 | 0.99 |
TOOD | 0.99 | 0.99 | 0.948 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 |
VarifocalNet | 0.979 | 0.938 | 0.854 | 0.958 | 0.969 | 0.979 | 0.906 | 0.979 | 0.833 | 0.896 |
YOLOv5 | 0.99 | 1.0 | 1.0 | 0.99 | 1.0 | 0.979 | 0.99 | 1.0 | 0.99 | 1.0 |
YOLOv6 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 | 0.979 | 0.99 |
YOLOv7 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.969 | 0.99 | 0.99 | 0.99 | 0.99 |
YOLOv8 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
YOLOX | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
Faster R-CNN | 0.979 | 0.885 | 0.76 | 0.99 | 0.958 | 0.979 | 0.802 | 0.958 | 0.875 | 0.896 |
Cascade R-CNN | 0.99 | 0.906 | 0.823 | 0.979 | 0.948 | 0.99 | 0.792 | 0.958 | 0.875 | 0.896 |
Cascade RPN | 0.99 | 0.979 | 0.958 | 0.979 | 0.979 | 0.99 | 0.927 | 0.99 | 0.979 | 0.958 |
Double Heads | 0.99 | 0.948 | 0.885 | 0.99 | 0.896 | 0.99 | 0.875 | 0.969 | 0.927 | 0.885 |
FPG | 0.979 | 0.99 | 0.979 | 0.979 | 0.979 | 0.979 | 0.906 | 0.99 | 0.958 | 0.958 |
Grid R-CNN | 0.969 | 0.958 | 0.917 | 0.979 | 0.99 | 0.979 | 0.885 | 0.99 | 0.979 | 0.896 |
Guided Anchoring | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.958 | 0.99 | 0.99 | 0.99 | 0.979 |
HRNet | 0.979 | 0.927 | 0.875 | 0.979 | 0.938 | 0.979 | 0.885 | 0.969 | 0.99 | 0.969 |
Libra R-CNN | 0.99 | 0.969 | 0.969 | 0.979 | 0.979 | 0.979 | 0.969 | 0.99 | 0.958 | 0.938 |
PAFPN | 0.979 | 0.875 | 0.781 | 0.969 | 0.938 | 0.979 | 0.823 | 0.958 | 0.885 | 0.885 |
RepPoints | 0.99 | 0.958 | 0.896 | 0.99 | 0.948 | 0.99 | 0.979 | 1.0 | 1.0 | 0.938 |
Res2Net | 0.979 | 0.99 | 0.979 | 0.99 | 0.979 | 0.979 | 0.979 | 0.99 | 0.99 | 0.979 |
ResNeSt | 0.979 | 0.927 | 0.958 | 0.969 | 0.969 | 0.979 | 0.979 | 0.969 | 0.917 | 0.969 |
SABL | 0.99 | 0.885 | 0.844 | 0.979 | 0.938 | 0.979 | 0.781 | 0.948 | 0.958 | 0.896 |
Sparse R-CNN | 0.979 | 0.979 | 0.865 | 0.99 | 0.99 | 0.979 | 0.938 | 0.99 | 0.906 | 0.979 |
DETR | 0.979 | 0.771 | 0.615 | 0.604 | 0.896 | 0.708 | 0.885 | 0.635 | 0.625 | 0.906 |
Conditional DETR | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 |
DDQ | 0.979 | 0.99 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 | 0.99 | 0.99 | 0.99 |
DAB-DETR | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 0.99 | 1.0 | 0.979 | 0.99 | 0.99 |
Deformable DETR | 0.99 | 0.99 | 0.969 | 0.979 | 0.969 | 0.969 | 0.948 | 1.0 | 0.979 | 0.979 |
DINO | 0.99 | 0.948 | 0.979 | 0.99 | 0.969 | 0.99 | 0.99 | 0.979 | 0.979 | 0.99 |
PVT | 0.958 | 0.979 | 0.792 | 0.958 | 0.969 | 0.979 | 0.875 | 0.979 | 0.969 | 0.958 |
PVTv2 | 0.979 | 0.99 | 0.917 | 1.0 | 1.0 | 0.99 | 0.875 | 0.99 | 0.99 | 0.948 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 1.0 | 0.98 | 0.86 | 0.98 | 0.99 | 1.0 | 0.92 | 1.0 | 0.9 | 0.91 |
AutoAssign | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 |
CenterNet | 0.99 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 0.99 | 0.98 |
CentripetalNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 |
CornerNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DDOD | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DyHead | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
EfficientNet | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FCOS | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FreeAnchor | 1.0 | 0.87 | 0.93 | 0.96 | 0.89 | 1.0 | 0.89 | 1.0 | 0.98 | 0.9 |
FSAF | 1.0 | 0.98 | 0.94 | 0.99 | 0.96 | 1.0 | 0.96 | 1.0 | 0.95 | 0.91 |
GFL | 1.0 | 0.95 | 0.93 | 0.99 | 1.0 | 1.0 | 0.95 | 1.0 | 0.86 | 0.9 |
LD | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 0.93 | 1.0 |
NAS-FPN | 1.0 | 1.0 | 0.98 | 0.99 | 0.98 | 1.0 | 0.98 | 1.0 | 0.99 | 0.97 |
PAA | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
RetinaNet | 1.0 | 0.97 | 0.98 | 0.99 | 1.0 | 1.0 | 0.96 | 1.0 | 0.91 | 0.89 |
RTMDet | 1.0 | 1.0 | 1.0 | 0.97 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 |
TOOD | 1.0 | 1.0 | 0.96 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
VarifocalNet | 0.99 | 0.94 | 0.97 | 0.98 | 0.99 | 1.0 | 0.94 | 1.0 | 0.88 | 0.91 |
YOLOv5 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv8 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOX | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 |
Faster R-CNN | 1.0 | 0.86 | 0.77 | 0.95 | 0.91 | 1.0 | 0.83 | 1.0 | 0.83 | 0.85 |
Cascade R-CNN | 1.0 | 0.89 | 0.91 | 0.99 | 0.97 | 1.0 | 0.93 | 0.99 | 0.88 | 0.9 |
Cascade RPN | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 0.99 | 0.96 | 1.0 | 1.0 | 0.93 |
Double Heads | 1.0 | 1.0 | 0.95 | 1.0 | 0.97 | 1.0 | 0.95 | 1.0 | 0.92 | 0.91 |
FPG | 1.0 | 1.0 | 0.96 | 0.99 | 1.0 | 0.99 | 0.97 | 1.0 | 1.0 | 0.99 |
Grid R-CNN | 0.98 | 0.97 | 0.89 | 1.0 | 0.95 | 0.99 | 0.95 | 1.0 | 0.94 | 0.9 |
Guided Anchoring | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 |
HRNet | 1.0 | 0.98 | 1.0 | 0.99 | 0.94 | 0.99 | 0.84 | 1.0 | 0.98 | 0.95 |
Libra R-CNN | 1.0 | 1.0 | 0.91 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 0.96 | 0.96 |
PAFPN | 1.0 | 0.93 | 0.83 | 0.97 | 0.99 | 1.0 | 0.95 | 1.0 | 0.9 | 0.9 |
RepPoints | 1.0 | 1.0 | 0.89 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 0.96 | 0.95 |
Res2Net | 0.98 | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 0.93 | 1.0 |
ResNeSt | 0.96 | 0.85 | 0.94 | 0.99 | 1.0 | 0.96 | 0.99 | 0.99 | 0.9 | 0.99 |
SABL | 1.0 | 1.0 | 0.89 | 0.99 | 0.96 | 1.0 | 0.92 | 1.0 | 0.89 | 0.89 |
Sparse R-CNN | 1.0 | 1.0 | 0.88 | 0.99 | 1.0 | 0.98 | 1.0 | 1.0 | 0.92 | 0.95 |
DETR | 1.0 | 0.84 | 0.72 | 0.49 | 0.93 | 0.93 | 0.96 | 0.87 | 0.81 | 0.94 |
Conditional DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
Deformable DETR | 0.99 | 1.0 | 0.99 | 1.0 | 0.99 | 1.0 | 0.99 | 1.0 | 0.98 | 1.0 |
DINO | 1.0 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.98 | 1.0 |
PVT | 1.0 | 0.98 | 0.7 | 0.98 | 0.96 | 1.0 | 0.92 | 1.0 | 0.96 | 0.85 |
PVTv2 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.82 | 1.0 | 1.0 | 0.97 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.98 | 0.96 | 0.47 | 0.73 | 0.94 | 0.85 | 0.77 | 1.0 | 0.76 | 0.7 |
AutoAssign | 1.0 | 0.86 | 0.74 | 0.92 | 0.91 | 0.97 | 0.71 | 0.93 | 0.87 | 0.98 |
CenterNet | 1.0 | 1.0 | 0.62 | 0.98 | 0.98 | 0.99 | 0.71 | 0.91 | 0.92 | 1.0 |
CentripetalNet | 1.0 | 1.0 | 0.78 | 1.0 | 0.96 | 1.0 | 0.74 | 1.0 | 0.88 | 0.98 |
CornerNet | 0.99 | 0.79 | 0.49 | 0.75 | 0.93 | 0.78 | 0.51 | 0.95 | 0.66 | 0.6 |
DDOD | 0.99 | 0.97 | 0.61 | 0.99 | 0.85 | 0.98 | 1.0 | 1.0 | 0.96 | 0.97 |
DyHead | 1.0 | 0.81 | 0.52 | 0.53 | 0.98 | 0.81 | 0.61 | 0.55 | 0.62 | 0.93 |
EfficientNet | 1.0 | 0.79 | 0.81 | 1.0 | 0.84 | 1.0 | 0.78 | 1.0 | 0.97 | 1.0 |
FCOS | 1.0 | 0.94 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 1.0 | 1.0 | 0.49 | 0.96 | 0.87 | 0.99 | 0.67 | 0.87 | 0.96 | 0.92 |
FreeAnchor | 1.0 | 0.82 | 1.0 | 0.98 | 0.9 | 1.0 | 0.99 | 0.98 | 0.88 | 1.0 |
FSAF | 1.0 | 1.0 | 0.64 | 1.0 | 0.82 | 1.0 | 0.99 | 1.0 | 1.0 | 1.0 |
GFL | 1.0 | 0.97 | 0.45 | 0.95 | 0.93 | 0.96 | 0.94 | 0.98 | 0.73 | 1.0 |
LD | 1.0 | 0.99 | 0.56 | 1.0 | 1.0 | 0.93 | 0.98 | 0.88 | 0.96 | 1.0 |
NAS-FPN | 1.0 | 0.92 | 0.52 | 0.86 | 1.0 | 0.79 | 0.72 | 0.94 | 0.65 | 0.92 |
PAA | 1.0 | 1.0 | 0.97 | 0.99 | 0.94 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 |
RetinaNet | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
RTMDet | 1.0 | 1.0 | 0.99 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.88 |
TOOD | 0.83 | 0.85 | 0.52 | 0.87 | 0.76 | 0.84 | 0.8 | 0.97 | 0.96 | 0.8 |
VarifocalNet | 1.0 | 0.77 | 0.48 | 0.55 | 0.86 | 0.83 | 0.64 | 0.62 | 0.61 | 0.65 |
YOLOv5 | 1.0 | 1.0 | 0.74 | 1.0 | 1.0 | 1.0 | 0.92 | 1.0 | 1.0 | 1.0 |
YOLOv6 | 1.0 | 1.0 | 0.73 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
YOLOv7 | 0.98 | 0.76 | 0.78 | 0.72 | 0.91 | 1.0 | 0.78 | 1.0 | 0.75 | 0.79 |
YOLOv8 | 1.0 | 0.92 | 0.64 | 0.82 | 1.0 | 0.82 | 0.64 | 1.0 | 0.83 | 0.97 |
YOLOX | 1.0 | 0.9 | 0.62 | 0.92 | 0.95 | 0.99 | 0.88 | 1.0 | 0.97 | 0.98 |
Faster R-CNN | 0.85 | 0.49 | 0.46 | 0.52 | 0.66 | 0.73 | 0.53 | 0.63 | 0.55 | 0.49 |
Cascade R-CNN | 0.75 | 0.64 | 0.47 | 0.54 | 0.67 | 0.67 | 0.62 | 0.65 | 0.55 | 0.6 |
Cascade RPN | 1.0 | 0.92 | 0.51 | 0.79 | 0.88 | 1.0 | 0.96 | 1.0 | 0.85 | 0.82 |
Double Heads | 0.76 | 0.72 | 0.46 | 0.57 | 0.68 | 0.71 | 0.63 | 0.87 | 0.6 | 0.62 |
FPG | 1.0 | 0.89 | 0.5 | 0.93 | 0.97 | 0.9 | 0.72 | 0.91 | 0.96 | 0.53 |
Grid R-CNN | 0.91 | 0.68 | 0.48 | 0.68 | 0.71 | 0.77 | 0.63 | 0.85 | 0.6 | 0.56 |
Guided Anchoring | 1.0 | 1.0 | 0.87 | 0.99 | 1.0 | 1.0 | 0.8 | 1.0 | 1.0 | 0.9 |
HRNet | 0.96 | 0.78 | 0.55 | 0.59 | 0.73 | 0.83 | 0.62 | 0.81 | 0.62 | 0.62 |
Libra R-CNN | 1.0 | 1.0 | 0.67 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
PAFPN | 0.74 | 0.63 | 0.4 | 0.51 | 0.63 | 0.71 | 0.62 | 0.66 | 0.57 | 0.44 |
RepPoints | 1.0 | 1.0 | 0.63 | 1.0 | 0.94 | 0.77 | 1.0 | 1.0 | 1.0 | 0.9 |
Res2Net | 0.9 | 0.64 | 0.57 | 0.55 | 0.73 | 0.65 | 0.66 | 0.77 | 0.66 | 0.72 |
ResNeSt | 0.97 | 0.61 | 0.5 | 0.73 | 0.72 | 0.96 | 0.61 | 0.93 | 0.61 | 0.58 |
SABL | 0.77 | 0.53 | 0.45 | 0.54 | 0.66 | 0.66 | 0.61 | 0.66 | 0.56 | 0.47 |
Sparse R-CNN | 1.0 | 0.96 | 0.63 | 0.91 | 1.0 | 0.72 | 0.8 | 0.78 | 0.87 | 0.85 |
DETR | 0.77 | 0.56 | 0.39 | 0.45 | 0.71 | 0.64 | 0.51 | 0.55 | 0.55 | 0.57 |
Conditional DETR | 1.0 | 0.88 | 0.67 | 0.84 | 1.0 | 0.8 | 0.86 | 1.0 | 0.84 | 0.95 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 | 0.98 |
Deformable DETR | 1.0 | 1.0 | 0.94 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DINO | 1.0 | 1.0 | 0.93 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
PVT | 1.0 | 1.0 | 0.95 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.96 |
PVTv2 | 1.0 | 0.93 | 0.64 | 0.94 | 0.87 | 1.0 | 0.76 | 1.0 | 0.76 | 0.6 |
Clean |
Random |
ACTIVE |
DTA |
FCA |
APPA |
POOPatch |
3D2Fool |
CAMOU |
RPAU |
ATSS | 0.97 | 0.71 | 0.45 | 0.79 | 0.84 | 0.85 | 0.68 | 0.98 | 0.71 | 0.73 |
AutoAssign | 0.98 | 0.9 | 0.74 | 0.92 | 0.94 | 1.0 | 0.83 | 0.98 | 0.91 | 0.99 |
CenterNet | 0.98 | 0.91 | 0.58 | 0.91 | 0.92 | 0.89 | 0.93 | 0.9 | 0.88 | 0.94 |
CentripetalNet | 0.98 | 0.78 | 0.66 | 0.79 | 0.9 | 0.86 | 0.7 | 0.85 | 0.71 | 0.86 |
CornerNet | 0.96 | 0.75 | 0.61 | 0.87 | 0.93 | 0.86 | 0.7 | 0.94 | 0.76 | 0.82 |
DDOD | 0.99 | 0.96 | 0.74 | 0.99 | 0.86 | 1.0 | 0.82 | 1.0 | 0.97 | 0.94 |
DyHead | 1.0 | 0.79 | 0.53 | 0.79 | 0.92 | 0.8 | 0.79 | 0.73 | 0.77 | 0.99 |
EfficientNet | 1.0 | 0.94 | 0.86 | 0.99 | 0.98 | 1.0 | 0.98 | 0.98 | 0.96 | 0.93 |
FCOS | 1.0 | 0.99 | 0.99 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
FoveaBox | 0.94 | 0.84 | 0.5 | 0.76 | 0.87 | 0.81 | 0.72 | 0.8 | 0.76 | 0.91 |
FreeAnchor | 0.97 | 0.73 | 0.79 | 0.88 | 0.81 | 0.98 | 0.74 | 0.85 | 0.84 | 0.88 |
FSAF | 0.93 | 0.75 | 0.49 | 0.8 | 0.72 | 0.81 | 0.67 | 0.89 | 0.74 | 0.84 |
GFL | 1.0 | 0.8 | 0.48 | 0.9 | 0.95 | 0.95 | 0.89 | 0.92 | 0.81 | 0.9 |
LD | 0.98 | 0.86 | 0.52 | 0.9 | 0.92 | 0.87 | 0.75 | 0.94 | 0.85 | 0.93 |
NAS-FPN | 1.0 | 0.92 | 0.66 | 0.96 | 0.98 | 0.97 | 0.73 | 0.82 | 0.75 | 0.88 |
PAA | 0.99 | 0.96 | 0.9 | 0.97 | 0.93 | 0.99 | 0.86 | 0.99 | 1.0 | 0.95 |
RetinaNet | 1.0 | 0.81 | 0.67 | 0.85 | 0.89 | 0.86 | 0.81 | 0.84 | 0.83 | 0.8 |
RTMDet | 1.0 | 1.0 | 0.93 | 0.99 | 1.0 | 0.98 | 0.99 | 1.0 | 1.0 | 0.98 |
TOOD | 0.9 | 0.72 | 0.48 | 0.9 | 0.73 | 0.79 | 0.72 | 0.84 | 0.74 | 0.86 |
VarifocalNet | 0.99 | 0.75 | 0.45 | 0.66 | 0.87 | 0.83 | 0.67 | 0.8 | 0.67 | 0.79 |
YOLOv5 | 1.0 | 1.0 | 0.98 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.97 | 0.99 |
YOLOv6 | 1.0 | 1.0 | 0.97 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 0.99 | 1.0 |
YOLOv7 | 1.0 | 0.94 | 0.97 | 0.94 | 0.98 | 1.0 | 0.96 | 1.0 | 0.98 | 0.99 |
YOLOv8 | 1.0 | 0.95 | 0.87 | 0.93 | 1.0 | 0.98 | 0.89 | 0.99 | 0.84 | 0.99 |
YOLOX | 1.0 | 0.97 | 0.77 | 0.97 | 0.99 | 1.0 | 0.98 | 0.92 | 0.94 | 0.98 |
Faster R-CNN | 0.85 | 0.45 | 0.35 | 0.47 | 0.54 | 0.59 | 0.49 | 0.61 | 0.44 | 0.49 |
Cascade R-CNN | 0.82 | 0.51 | 0.4 | 0.52 | 0.58 | 0.64 | 0.53 | 0.62 | 0.48 | 0.51 |
Cascade RPN | 1.0 | 0.91 | 0.62 | 0.93 | 0.96 | 0.99 | 0.96 | 1.0 | 0.86 | 0.96 |
Double Heads | 0.83 | 0.61 | 0.44 | 0.58 | 0.58 | 0.72 | 0.54 | 0.8 | 0.57 | 0.57 |
FPG | 1.0 | 0.78 | 0.5 | 0.93 | 0.82 | 0.99 | 0.7 | 0.94 | 0.9 | 0.71 |
Grid R-CNN | 0.89 | 0.51 | 0.37 | 0.6 | 0.58 | 0.73 | 0.55 | 0.76 | 0.47 | 0.55 |
Guided Anchoring | 1.0 | 0.97 | 0.98 | 0.99 | 0.99 | 1.0 | 0.89 | 1.0 | 0.97 | 0.97 |
HRNet | 0.89 | 0.6 | 0.52 | 0.59 | 0.54 | 0.73 | 0.51 | 0.84 | 0.56 | 0.6 |
Libra R-CNN | 0.94 | 0.76 | 0.49 | 0.8 | 0.9 | 0.74 | 0.85 | 0.81 | 0.72 | 0.93 |
PAFPN | 0.84 | 0.55 | 0.37 | 0.52 | 0.59 | 0.65 | 0.58 | 0.71 | 0.5 | 0.54 |
RepPoints | 0.97 | 0.89 | 0.52 | 0.84 | 0.83 | 0.84 | 0.8 | 0.95 | 0.94 | 0.84 |
Res2Net | 0.88 | 0.59 | 0.52 | 0.66 | 0.77 | 0.83 | 0.54 | 0.79 | 0.55 | 0.73 |
ResNeSt | 0.91 | 0.63 | 0.43 | 0.8 | 0.78 | 0.91 | 0.68 | 0.92 | 0.64 | 0.62 |
SABL | 0.82 | 0.49 | 0.4 | 0.52 | 0.56 | 0.57 | 0.53 | 0.71 | 0.46 | 0.49 |
Sparse R-CNN | 0.99 | 0.89 | 0.55 | 0.79 | 0.95 | 0.81 | 0.83 | 0.9 | 0.82 | 0.86 |
DETR | 0.71 | 0.44 | 0.3 | 0.32 | 0.57 | 0.52 | 0.63 | 0.43 | 0.38 | 0.5 |
Conditional DETR | 1.0 | 0.9 | 0.71 | 0.94 | 1.0 | 0.92 | 0.8 | 0.98 | 0.9 | 0.94 |
DDQ | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 | 1.0 |
DAB-DETR | 1.0 | 0.94 | 1.0 | 0.99 | 0.99 | 1.0 | 0.98 | 0.96 | 1.0 | 0.99 |
Deformable DETR | 1.0 | 0.9 | 0.74 | 0.86 | 0.87 | 0.93 | 0.95 | 0.96 | 0.86 | 0.91 |
DINO | 1.0 | 0.92 | 0.95 | 0.99 | 0.96 | 1.0 | 1.0 | 1.0 | 0.92 | 1.0 |
PVT | 0.98 | 0.95 | 0.78 | 0.87 | 0.96 | 0.97 | 0.91 | 0.94 | 0.87 | 0.8 |
PVTv2 | 1.0 | 0.99 | 0.77 | 0.97 | 1.0 | 1.0 | 0.78 | 0.99 | 0.98 | 0.8 |
Clean | AdvCam | RP2 | ShapeShifter | |
ATSS | 0.929 | 0.93 | 0.892 | 0.919 |
AutoAssign | 0.921 | 0.943 | 0.896 | 0.915 |
CenterNet | 0.903 | 0.91 | 0.875 | 0.914 |
CentripetalNet | 0.951 | 0.946 | 0.924 | 0.951 |
CornerNet | 0.942 | 0.951 | 0.929 | 0.947 |
DDOD | 0.916 | 0.926 | 0.9 | 0.915 |
DyHead | 0.927 | 0.933 | 0.866 | 0.921 |
EfficientNet | 0.921 | 0.913 | 0.887 | 0.919 |
FCOS | 0.939 | 0.932 | 0.913 | 0.929 |
FoveaBox | 0.915 | 0.917 | 0.871 | 0.913 |
FreeAnchor | 0.921 | 0.919 | 0.877 | 0.912 |
FSAF | 0.901 | 0.907 | 0.864 | 0.899 |
GFL | 0.927 | 0.942 | 0.887 | 0.908 |
LD | 0.933 | 0.931 | 0.88 | 0.92 |
NAS-FPN | 0.93 | 0.942 | 0.883 | 0.925 |
PAA | 0.928 | 0.921 | 0.886 | 0.91 |
RetinaNet | 0.921 | 0.912 | 0.863 | 0.899 |
RTMDet | 0.929 | 0.942 | 0.862 | 0.927 |
TOOD | 0.922 | 0.93 | 0.895 | 0.921 |
VarifocalNet | 0.929 | 0.932 | 0.902 | 0.921 |
YOLOv5 | 0.941 | 0.943 | 0.882 | 0.945 |
YOLOv6 | 0.94 | 0.954 | 0.901 | 0.936 |
YOLOv7 | 0.945 | 0.942 | 0.89 | 0.942 |
YOLOv8 | 0.942 | 0.943 | 0.868 | 0.944 |
YOLOX | 0.923 | 0.922 | 0.866 | 0.906 |
Faster R-CNN | 0.891 | 0.897 | 0.863 | 0.861 |
Cascade R-CNN | 0.929 | 0.924 | 0.891 | 0.895 |
Cascade RPN | 0.927 | 0.93 | 0.887 | 0.901 |
Double Heads | 0.887 | 0.895 | 0.847 | 0.88 |
FPG | 0.921 | 0.935 | 0.859 | 0.897 |
Grid R-CNN | 0.913 | 0.911 | 0.865 | 0.91 |
Guided Anchoring | 0.928 | 0.921 | 0.899 | 0.925 |
HRNet | 0.923 | 0.915 | 0.91 | 0.904 |
Libra R-CNN | 0.921 | 0.922 | 0.885 | 0.905 |
PAFPN | 0.901 | 0.89 | 0.85 | 0.882 |
RepPoints | 0.915 | 0.908 | 0.865 | 0.887 |
Res2Net | 0.91 | 0.897 | 0.861 | 0.899 |
ResNeSt | 0.929 | 0.901 | 0.872 | 0.889 |
SABL | 0.92 | 0.913 | 0.88 | 0.898 |
Sparse R-CNN | 0.931 | 0.925 | 0.9 | 0.927 |
DETR | 0.908 | 0.904 | 0.878 | 0.933 |
Conditional DETR | 0.93 | 0.931 | 0.907 | 0.924 |
DDQ | 0.949 | 0.95 | 0.901 | 0.932 |
DAB-DETR | 0.931 | 0.94 | 0.902 | 0.919 |
Deformable DETR | 0.943 | 0.955 | 0.893 | 0.933 |
DINO | 0.94 | 0.944 | 0.885 | 0.936 |
PVT | 0.906 | 0.886 | 0.868 | 0.861 |
PVTv2 | 0.915 | 0.899 | 0.877 | 0.905 |
B.5 Additional ablation experiments
Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 12, 13, 14, 15, 16, and 17. In addition, the visualized evaluation results are shown in Fig. 16, 17, and 18.
Appendix C Reproducibility
The reproducibility of all evaluation results in our benchmark is ensured through the provision of the necessary code in the associated Github Repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack). Furthermore, all datasets and model architectures evaluated are accessible to the public.
Appendix D License
This work is under Apache-2.0 license. More detailed information can be found in the LICENSE file in the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack

### v3

@@
 Abstract
-Physical attacks against object detection have gained increasing attention due to their significant practical implications.
-However, conducting physical experiments is extremely time-consuming and labor-intensive.
-Moreover, physical dynamics and cross-domain transformation are challenging to strictly regulate in the real world, leading to unaligned evaluation and comparison, severely hindering the development of physically robust models.
-To accommodate these challenges, we explore utilizing realistic simulation to thoroughly and rigorously benchmark physical attacks with fairness under controlled physical dynamics and cross-domain transformation.
-This resolves the problem of capturing identical adversarial images that cannot be achieved in the real world.
-Our benchmark includes 20 physical attack methods, 48 object detectors, comprehensive physical dynamics, and evaluation metrics.
-We also provide end-to-end pipelines for dataset generation, detection, evaluation, and further analysis.
-In addition, we perform 8064 groups of evaluation based on our benchmark, which includes both overall evaluation and further detailed ablation studies for controlled physical dynamics.
-Through these experiments, we provide in-depth analyses of physical attack performance and physical adversarial robustness, draw valuable observations, and discuss potential directions for future research.
-Codebase: GitHub Repository.
+Physical attacks against object detection have gained significant attention due to their practical implications.
+However, conducting physical experiments is time-consuming and labor-intensive, and controlling physical dynamics and cross-domain transformations in the real world is challenging, leading to inconsistent evaluations and hindering the development of robust models.
+To address these issues, we rigorously explore realistic simulations to benchmark physical attacks under controlled conditions.
+This approach ensures fairness and resolves the problem of capturing strictly aligned adversarial images, which is challenging in the real world.
+Our benchmark includes 23 physical attacks, 48 object detectors, comprehensive physical dynamics, and evaluation metrics.
+We provide end-to-end pipelines for dataset generation, detection, evaluation, and analysis.
+The benchmark is flexible and scalable, allowing easy integration of new objects, attacks, models, and vision tasks.
+Based on this benchmark, we generate comprehensive datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies.
+These experiments provide detailed analyses from detection and attack perspectives, highlight limitations of existing algorithms, and offer revealing insights.
+The code and datasets are publicly available at https://github.com/JiaweiLian/PADetBench.
+keywords:
+[…]
 1 Introduction
-Deep neural networks (DNNs) have achieved remarkable success in various research fields and industrial applications, such as computer vision [65], natural language processing [64], and speech recognition [61].
-However, previous studies [78, 27, 2, 41, 3] have shown that DNNs are susceptible and vulnerable to adversarial attacks, which can be divided into digital attacks and physical attacks according to different attack domains in computer vision tasks.
-Technically, digital attacks are usually performed by adding imperceptible perturbations to the input image of victim models after imaging, while physical attacks can cheat DNNs by directly modifying the physical properties of the targets before imaging, such as changing the texture of the object [77, 114], and adding stickers [96, 44].
-Intuitively, physical attacks are more dangerous and practical than digital attacks because they can be easily implemented in the real world with fewer prerequisites, which arouses significant concerns, especially in safety-critical applications like autonomous driving [92, 6], security surveillance [63, 95], remote sensing [94, 48], and so on.
-Object detection is a fundamental and pragmatic task in computer vision, which has been widely deployed in various intelligent systems [126, 113].
-Consequently, most of the physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has attracted increasing attention in recent years.
+Deep neural networks (DNNs) have achieved remarkable success in various fields such as computer vision, natural language processing, and speech recognition.
+However, studies [1, 2, 3, 4] show that DNNs are vulnerable to adversarial attacks, which can be categorized into digital and physical attacks.
+Digital attacks add imperceptible perturbations to input images post-imaging, while physical attacks modify the physical properties of targets pre-imaging, such as changing textures [5, 6] or adding stickers [7, 8].
+Physical attacks are more practically dangerous as they can be easily implemented in real-world scenarios, raising significant concerns in safety-critical applications like autonomous driving [9, 10, 11], security surveillance [12, 13], and remote sensing [14, 15].
+Object detection is a fundamental and pragmatic task in computer vision, widely deployed in various intelligent systems [16, 17, 18, 19].
+Consequently, many physical attacks aim to fool object detectors in real-world scenarios, and the physical adversarial robustness of object detection models has garnered increasing attention in recent years.
 However, the absence of regulated and easy-to-follow benchmarks hinders the development of physical attack and physically robust detection methods.
-The main reasons for discouraging the development of physical attack benchmarks are as follows.
-1) Time-consuming and expensive: To evaluate the physical attack performance and physical adversarial robustness of object detection models, researchers need to conduct a large number of experiments in the real world, which is time-consuming and expensive.
-2) Physical dynamics align: To ensure comparison fairness, physical dynamics should be strictly controlled and kept the same, while this can’t be achieved in real-world scenarios since it is impossible to take two identical pictures.
-3) Cross-domain loss: Physical attacks are commonly conducted by elaborating conspicuous adversarial patterns that should be able to survive after transforming from the physical to the digital domain, while this cross-domain loss is uncontrollable.
-4) Hard to compare: With the progressive evolution of physical attacks, physical adversarial perturbations have evolved from 2D space to 3D space, which builds a barrier to fairly compare different kinds of physical attack methods.
-As the above challenges exist, it is tough to effectively verify the physical attack efficacy and physical adversarial robustness of object detection models without thorough evaluation and impartial comparisons.
-As a result, the researchers can not identify the actual progress of physical adversarial attacks and physical adversarial robustness development, which may gear down the progress of the research field.
-In this paper, we propose a fair and rigorous benchmarking of physical attacks on object detection models.
-We utilize CARLA [17], a highly realistic autonomous driving simulator, to benchmark physical attacks under consistent physical dynamics such as weather, viewing angle, and distance.
-These conditions are challenging to align for impartial comparisons in the real world due to the impossibility of capturing two identical photographs.
-Our benchmark includes 20 physical attack methods, clean and random noised objects for comparison, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, and a comprehensive pipeline for evaluating physical attacks and subsequent analysis.
-We conduct exhaustive evaluation of each physical attack against a wide range of object detectors under various physical dynamic conditions, using multiple evaluation metrics.
-Following our experimental procedures, we delve into detailed analyses concerning the performance of physical attacks and the robustness against physical adversarial actions.
-We highlight intriguing findings and discuss potential avenues for future research.
+The main reasons for the lack of physical attack benchmarks are concluded as follows:
+Time-consuming and expensive: Evaluating the performance of physical attacks and the adversarial robustness of object detection models requires numerous real-world experiments, which are both time-consuming and costly.
+Physical dynamics alignment: Ensuring fair comparisons necessitates strictly controlled and consistent physical dynamics, which is difficult to achieve in real-world scenarios where it is nearly impossible to capture identical physical dynamics.
+Cross-domain loss: Physical attacks often involve creating conspicuous adversarial perturbations that must survive the transformation from the physical to the digital domain and vice versa.
+This cross-domain loss is often uncontrollable and unpredictable.
+Difficulty in comparison: With the evolution of physical attacks from 2D to 3D space, it becomes increasingly challenging to compare different types of physical attack methods fairly.
+Due to these challenges, it is difficult to effectively verify the efficacy of physical attacks and the adversarial robustness of object detection models without thorough evaluation and impartial comparisons.
+As a result, researchers cannot accurately gauge the progress of physical adversarial attacks and robustness development, which slows down advancements in the field.
+This paper proposes utilizing realistic simulations to benchmark physical attacks under controlled conditions such as weather, viewing angle, and location.
+These conditions are challenging to align for impartial comparisons in the real world.
+Our benchmark includes 23 physical attack methods, 48 object detectors, diverse physical dynamics, evaluation metrics from different perspectives, comprehensive pipelines for data generation, attack and detection evaluation, and subsequent analysis.
+Moreover, the benchmark is highly flexible and scalable, allowing for easy integration of new physical attacks, models, and other vision tasks.
+Based on the benchmark, we generate comprehensive and strictly aligned datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies for controlled physical dynamics.
+Through these experiments, we provide detailed analyses from detection and attack perspectives, highlight algorithm limitations, and convey valuable insights.
 In summary, our contributions are as follows:
 We propose a robust and equitable benchmark for physical attacks against object detection models.
-This benchmark deeply explores the potential of real-world simulators to consistently evaluate physical attacks under a variety of continuous physical dynamics.
-We offer 8064 groups of evaluation on 20 physical attacks against 48 object detectors, which are conducted under continuous physical dynamics with diverse evaluation metrics, providing a holistic perspective on the performance of attack and detection.
-We delve into detailed analyses of the fooling efficacy of physical attacks and the robustness of object detection models against these attacks.
-Through these analyses, we highlight valuable findings and discuss potential avenues for future research in this area.
+This benchmark deeply explores the potential of real-world simulators to evaluate physical attacks under various continuous and strictly aligned physical dynamics.
+The benchmark includes 23 physical attacks, 48 object detectors, comprehensive physical dynamics, and rigorous evaluation metrics.
+We provide end-to-end pipelines for dataset generation, detection, evaluation, and analysis, ensuring a thorough evaluation process.
+The benchmark is designed to be highly flexible and scalable, facilitating the easy integration of new objects, attacks, detectors, and even other vision tasks.
+This adaptability enhances the utility of our framework for ongoing research and development in the field.
+Based on our benchmark, we generate comprehensive datasets and perform over 8,000 evaluations, including overall assessments and detailed ablation studies.
+These experiments highlight the limitations of existing algorithms and illuminate informative insights.
 2 Related work
 2.1 Object detection
-Object detection serves as a cornerstone in computer vision, with the primary objective of identifying and localizing objects within images or video sequences.
-With the advent of deep learning, DNNs have been widely adopted for object detection, significantly improving performance.
-R-CNN [26] and its successors [25, 70, 56, 66, 99, 106, 75], have made substantial strides in improving detection speed and accuracy by introducing region proposal networks and shared convolution computations.
-Single Shot MultiBox Detector (SSD) [55] and You Only Look Once (YOLO) [67] series [67, 68, 69, 1, 38, 43, 86, 37, 87, 85] further accelerate the detection process by eliminating the need for region proposals and performing detection in a single pass, making them suitable for real-time applications.
-Recently, more advanced architectures based on transformers have been proposed, such as DETR [7], DAB-DETR [52], ViTDet [47], DINO [104], and Co-DETR [125], which further push the boundaries of object detection performance by leveraging attention mechanisms.
-Despite these advancements, object detection in adversarial environments remains a challenging task, necessitating continuous research and development in this area.
+Object detection is a fundamental task in computer vision, aiming to identify and localize objects within images or videos.
+It can be formulated as a mapping function , where is the input space and is the output space (e.g., bounding boxes and class labels).
+Deep learning has significantly advanced object detection.
+R-CNN [20] and its successors [21, 22, 23, 24, 25, 26, 27] improved detection speed and accuracy with region proposal networks and shared convolution computations.
+SSD [28] and YOLO series [29, 30, 31, 32, 33, 34, 35, 36, 37, 38] further accelerated detection by eliminating region proposals, enabling real-time applications.
+Recently, transformer-based architectures like DETR [39], DAB-DETR [40], ViTDet [41], DINO [42], and Co-DETR [43] have pushed performance boundaries using attention mechanisms.
+Despite increasing advancements [44, 45, 46], object detection in adversarial environments remains challenging, requiring ongoing research.
 2.2 Physical attack
-Kurakin et al.
-[41] first illustrate that machine learning systems remain susceptible to adversarial examples, even within these physical world contexts.
-They substantiate this by introducing adversarial images, captured via a cell phone camera, which still can dramatically degrade the performance of intelligent vision systems.
-Brown et al.
-[2] first design adversarial perturbations in patch form, i.e., adversarial patch, which localizes adversarial perturbations to a specific region of the image without imperceptibility constraints.
-Adversarial patches are more practical and effective in the real world, as they can be easily printed and attached to objects to fool detectors [73, 81, 100, 124, 122, 91, 121, 34, 112, 72, 35, 29].
-Nonetheless, the adversarial patches mentioned above are overly noticeable, potentially arousing the suspicion of the victims.
-To mitigate this concern, scholars have suggested the generation of adversarial patches in natural styles to circumvent drawing attention [36, 32, 28].
-Beyond adversarial patches, the academic community has also investigated physical perturbations in a multitude of forms, including light [31, 98], viewpoint [15], and 3D objects [51], among others.
-To reconcile the disparity between 2D perturbations and 3D physical space, the concept of adversarial perturbations has been expanded into the 3D realm [111, 88, 76, 77, 117].
-This extension has proven to be more efficacious and applicable in real-world scenarios.
-The considerable variation in physical perturbations and experimental settings presents a hurdle to the equitable comparison of disparate physical attack methodologies.
+Adversarial attacks typically add imperceptible perturbations to the clean input in the digital domain, fooling DNNs into incorrect predictions.
+This is formulated as: , where is the attack loss and is the ground-truth.
+In contrast, physical attacks often manipulate the physical properties of objects to deceive detection models, formulated as: , where and are transformations between digital and physical domains.
+[47] first showed that machine learning systems are vulnerable to adversarial examples in physical contexts.
+They demonstrated this with adversarial images captured via a cell phone camera, significantly degrading vision system performance.
+[48] introduced adversarial patches, which localize perturbations to specific image regions without imperceptibility constraints.
+These patches are practical and effective in the real world, easily printed and attached to objects to fool detectors [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60].
+To avoid suspicion, natural-style adversarial patches have been proposed [61, 62, 63].
+Beyond patches, physical perturbations include light [64, 65], viewpoint [66], and 3D objects [67].
+Extending adversarial perturbations to 3D space [68, 69, 70, 5, 71] has proven more effective and applicable in real-world scenarios.
+The variety in perturbations and settings complicates fair comparisons of physical attack methods.
 2.3 Robustness benchmark
-The benchmarking of adversarial attacks is crucial for evaluating the robustness of DNNs-based models and further developing more robust ones.
-To address the issue of overestimated robustness, [12] first establish a standardized benchmark of adversarial robustness, which accurately reflects the robustness of the considered models within a reasonable computational budget.
-Wu et al.
-[97] build a comprehensive benchmark for backdoor attacks in image classification models.
-[60] provides an easy-to-use benchmark to assess the performance of object detection models under conditions of deteriorating image quality, such as image distortions or diverse weather conditions instead of adversarial perturbations.
-In contrast, [115] benchmark the adversarial robustness of image classifiers in black-box settings.
-Except for RGB images, Dong et al.
-[14] benchmark the robustness of 3D object detection to common corruptions from weather, sensor, motion, etc., in both LiDAR and camera data.
-Li et al.
-[45] towards benchmarking and assessing the visual naturalness of physical adversarial perturbations.
-Hingun et al.
-[30] construct a large-scale benchmark for evaluating adversarial patches with traffic sign dataset.
-CARLA [17], as a realistic autonomous driving simulator, has been progressively involved in the research of physical adversarial robustness.
-[62] presents CARLA-GEAR: a dataset generator for evaluating the adversarial robustness of vision models.
-The work in [109] put forth a pipeline for instance-level data generation, leveraging the CARLA simulator.
-Utilizing this pipeline, they created the DCI dataset and carried out experiments involving three detectors and three physical attacks.
-In summary, there is still a lack of a comprehensive and rigorous benchmark for physical attacks against object detection models.
-We aim to fill the gap in this work and significantly lower the threshold of this path with easy-to-follow instructions and codebase.
+Benchmarking adversarial attacks is crucial for evaluating and improving the robustness of DNN-based models.
+[72] established a standardized benchmark for adversarial robustness, accurately reflecting model robustness within a reasonable computational budget.
+[73] created a comprehensive benchmark for backdoor attacks in image classification models.
+[74] provided a benchmark to assess object detection models under deteriorating image quality, such as distortions or adverse weather conditions.
+[75] benchmarked the adversarial robustness of image classifiers in black-box settings.
+[76] evaluated the robustness of 3D object detection to common corruptions in LiDAR and camera data.
+[77] focused on benchmarking the visual naturalness of physical adversarial perturbations.
+[78] constructed a large-scale benchmark for evaluating adversarial patches using a traffic sign dataset.
+CARLA [79], a realistic autonomous driving simulator, has been widely used in physical adversarial robustness research.
+[80] presented CARLA-GEAR, a dataset generator for evaluating the adversarial robustness of vision models.
+[81] proposed a pipeline for instance-level data generation using CARLA, creating the DCI dataset and conducting experiments with three detectors and three physical attacks.
+Despite these efforts, a comprehensive and rigorous benchmark for physical attacks against object detection models is still lacking.
+This work aims to fill that gap by providing easy-to-follow instructions and a robust codebase.
 3 PADetBench
-The benchmark encompasses four integral facets: datasets, physical attack implementations, object detection models, and comprehensive evaluation & analysis procedures.
-From a technical standpoint, we have engineered each constituent of the benchmark as modular, end-to-end pipelines within our publicly accessible codebase, ensuring straightforward adoption and replication.
-3.1 Dataset Generation
+The benchmark encompasses four integral facets: dataset generation, physical attacks, object detection, and comprehensive evaluation & analysis procedures, as shown in Fig.
+1.
+From a technical standpoint, we have engineered each benchmark constituent as modular, end-to-end pipelines within the codebase, ensuring straightforward adoption and replication.
+3.1 Datasets Generation
 It is common to use COCO [82], PASCAL VOC [83], KITTI [84], etc., as benchmark datasets for object detection.
-However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility required to create manipulated, real-world adversarial scenarios.
+However, these datasets are ill-suited for assessing physical attacks since they are static and lack the flexibility to create manipulated, real-world adversarial scenarios.
 Physical attacks typically entail altering the physical attributes of objects before capturing their images.
 To fairly and accurately evaluate and compare such attacks, experiments necessitate applying perturbations in real-world conditions with controlled physical dynamics, which are excessively time-consuming, labor-intensive, and theoretically infeasible.
-Simulated environments, like CARLA [17], present a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
+Simulated environments, like CARLA [79], offer a viable solution to these obstacles by enabling the straightforward manipulation of physical dynamics through configurable parameters.
 This work contributes an end-to-end pipeline for dataset generation within our codebase, significantly streamlining the dataset generation process and enhancing research productivity.
-Our pipeline prioritizes user-friendliness, enabling researchers to swiftly generate datasets embodying diverse physical conditions through a concise series of steps.
-These conditions encompass variations in weather, viewing angles, and distances, along with the capacity to impose physical perturbations on objects.
-Comprehensively, our pipeline supports over 10 distinct environments ranging from downtowns to small towns and rural landscapes (detailed listing available in the Appendix), coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures.
-It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning both location and orientation.
+Our pipeline prioritizes user-friendliness, enabling researchers to generate datasets embodying diverse physical conditions through concise steps swiftly.
+These conditions encompass variations in weather, viewing angles, distances, and the capacity to impose physical perturbations on objects.
+Comprehensively, our pipeline supports over ten distinct environments ranging from downtowns to small towns and rural landscapes.
+It is coupled with a library of more than 40 vehicles and 40 pedestrian models, all customizable concerning their hues and surface textures.
+It further integrates continuous manipulation of physical dynamics such as fluctuating weather patterns, precise sun positioning, and flexible camera placements concerning location and orientation.
+We provide a detailed illustration of the physical dynamics alignment in Fig.
+2 and Fig.
+3.
+Specifically, it is observed from Fig.
+2 that the imaging settings and lighting conditions are not strictly aligned in the comparison experiments, such as the different view angles and shadows, which have been demonstrated to have a significant impact on fooling deep neural networks [85, 66].
+To address this issue, we align the physical dynamics in the benchmark, as shown in Fig.
+3, where the physical dynamics are strictly controlled and aligned, ensuring a fair and impartial comparison.
+Moreover, we also provide a detailed illustration of the physical dynamics in Fig.
+4, which includes the weather conditions, camera settings, and lighting conditions.
+The lighting conditions vary aligned with real-world laws as shown in Fig.
+5, such as the sun positions of 24 hours, the intensity of the light, and the shadow, which are strictly controlled and aligned in the benchmark.
+In addition, we detail the diverse scenes and objects in A.2.
+Our benchmark comprises three datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies.
+To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline.
 To ensure accessibility, we accompany the pipeline with step-by-step guidelines for personalizing object perturbations and seamlessly integrating these modifications within CARLA’s [79] simulation framework.
-Our benchmark comprises three categories of datasets: a clean dataset serving as a control group, a dataset with random noise perturbations, and several datasets featuring adversarial perturbations generated through various attack methodologies.
-To ensure fair comparisons, scene compositions and camera perspectives are meticulously synchronized and regulated across all datasets, achievable effortlessly through our provided pipeline.
 Moreover, our pipeline facilitates the automatic generation of supplementary annotations, including 2D and 3D bounding boxes, depth maps, and instance segmentation maps.
-Consequently, our benchmark extends its utility beyond 2D object detection, also catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
+Consequently, our benchmark extends its utility beyond 2D object detection, catering to tasks like 3D object detection, instance segmentation, depth estimation, and more, thereby enhancing the scope of research and application in computer vision.
 3.2 Physical attacks
-We select 20 representative physical attack methods, in which FCA [88], DTA [76], ACTIVE [77], 3D2Fool [114], POOPatch [11], RPAU [54], APPA [48], and CAMOU [111], are for vehicle detection, while DAP [29], AdvPattern [95], UPC [36], NatPatch [32], MTD [13], AdvCaT [33], AdvTexture [34], AdvTshirt[101], AdvPatch [81], LAP [79], InvisCloak [100], and AdvCam [18], are for person detection.
+Physical attacks are usually tailored for specific objects, and the commonly targeted objects are vehicles, persons, and traffic signs, as evidenced by [86].
+Consequently, we adopt typical objects from these categories as examples to illustrate the proposed benchmark.
+Specifically, we select 23 representative physical attack methods, categorized into three types according to their target objects: vehicle, person, and traffic sign, as shown in Table 1.
 The corresponding physical perturbations of these methods are imported into Unreal Engine 4 for CARLA (CarlaUE4) [79], as shown in the physical attacks part of Fig.
 1, to generate the physical adversarial datasets.
-We adhere to two principles similar to [97] when selecting physical attacks.
-First, the methods are representative or advanced in the research field, which can serve as baseline and state-of-the-art (SOTA) methods for comparison, respectively.
-Second, physical attacks are easily conducted and with reproducible performance, which can be conveniently followed and reproduced by other researchers.
+When selecting physical attacks, we adhere to two principles similar to [73].
+First, the methods are representative or advanced in the research field, which can serve as a baseline and state-of-the-art (SOTA) methods for comparison, respectively.
+Second, physical attacks are efficiently conducted with reproducible performance, which other researchers can conveniently follow and reproduce.
 Since our benchmark evaluates physical attacks based on their crafted perturbations, novel physical attack methods can be easily integrated into the benchmark by following the provided pipeline.
 We will continue to update the physical attacks in the benchmark to keep pace with the latest research progress.
+Target objects | Physical attacks |
+Vehilcle | FCA [69], DTA [70], ACTIVE [5], |
+3D2Fool [6], POOPatch [87], | |
+RPAU [88], CAMOU [68] | |
+Person | DAP [60], AdvPattern [13], UPC [61], |
+NatPatch [62], MTD [89], AdvCaT [90], | |
+AdvTexture [56], AdvTshirt[91], AdvPatch [50], | |
+LAP [92], InvisCloak [51], AdvCam [93] | |
+Traffic sign | AdvCam [93], RP2 [94], ShapeShifter[95] |
 Backbone | Category | Detectors |
-CNN | One-stage | ATSS[107], AutoAssign[119], CenterNet[118], CentripetalNet[16] |
-CornerNet[42], DDOD[10], DyHead[99], EfficientNet[80], FCOS[82], | ||
-FoveaBox[40], FreeAnchor[110], FSAF[120], GFL[46],LD[116], NAS-FPN[24], | ||
-PAA[39], RetinaNet[49], RTMDet[58], TOOD[20], VarifocalNet[105], | ||
-YOLOX[22], YOLOv5[38], YOLOv6[43], YOLOv7[86], YOLOv8[37] | ||
-Two-stage | Faster R-CNN[70], Cascade R-CNN[5], Cascade RPN[84], | |
-Double Heads[99], FPG[8],Grid R-CNN[56], Guided Anchoring[89], | ||
-HRNet[74], Libra R-CNN[66], PAFPN[53],RepPoints[102], | ||
-Res2Net[21], ResNeSt[103], SABL[90], Sparse R-CNN[75] | ||
-Transformer | - | DETR[7], Conditional DETR[59], DDQ[108], DAB-DETR[52], |
-Deformable DETR[123], DINO[104],PVT[93], PVTv2[93] |
+CNN | One-stage | ATSS[96], AutoAssign[97], NAS-FPN [98] |
+LD[99], GFL[100], CornerNet[101], PAA[102], | ||
+DDOD[103], DyHead[25], EfficientNet[104], | ||
+FCOS[105], FoveaBox[106], FreeAnchor[107], | ||
+CenterNet[108], CentripetalNet[109], FSAF[110], | ||
+RTMDet[111], TOOD[112], VarifocalNet[113], | ||
+YOLOX[114], YOLOv5[33], YOLOv6[34], | ||
+YOLOv7[35], RetinaNet[115], YOLOv8[36] | ||
+Two-stage | Faster R-CNN[22], Cascade R-CNN[116], | |
+Cascade RPN[117], Double Heads[25], FPG[118], | ||
+Libra R-CNN[24], PAFPN[119], HRNet[120], | ||
+ResNeSt[121], Res2Net[122], SABL[123], | ||
+Guided Anchoring[124], Sparse R-CNN[27], | ||
+RepPoints[125], Grid R-CNN[23] | ||
+Transformer | - | DETR[39], PVT[126], PVTv2[126], |
+DDQ[127], DAB-DETR[40], DINO[42], | ||
+Deformable DETR[128], Conditional DETR[129] |
+Clean |
+Random |
+ACTIVE |
+DTA |
+FCA |
+APPA |
+POOPatch |
+3D2Fool |
+CAMOU |
+RPAU |
+ATSS | 0.973 | 0.808 | 0.518 | 0.84 | 0.883 | 0.811 | 0.732 | 0.924 | 0.549 | 0.826 |
+AutoAssign | 0.946 | 0.846 | 0.703 | 0.876 | 0.849 | 0.888 | 0.763 | 0.938 | 0.856 | 0.901 |
+CenterNet | 0.976 | 0.926 | 0.674 | 0.893 | 0.901 | 0.897 | 0.806 | 0.97 | 0.873 | 0.895 |
+CentripetalNet | 0.943 | 0.867 | 0.753 | 0.908 | 0.901 | 0.918 | 0.749 | 0.958 | 0.81 | 0.853 |
+CornerNet | 0.948 | 0.8 | 0.692 | 0.905 | 0.847 | 0.902 | 0.697 | 0.959 | 0.722 | 0.829 |
+DDOD | 0.95 | 0.936 | 0.756 | 0.923 | 0.9 | 0.934 | 0.863 | 0.948 | 0.949 | 0.909 |
+DyHead | 0.977 | 0.836 | 0.606 | 0.792 | 0.903 | 0.845 | 0.838 | 0.843 | 0.685 | 0.927 |
+EfficientNet | 0.977 | 0.974 | 0.912 | 0.975 | 0.974 | 0.97 | 0.951 | 0.973 | 0.966 | 0.948 |
+FCOS | 0.981 | 0.974 | 0.93 | 0.968 | 0.951 | 0.975 | 0.931 | 0.972 | 0.939 | 0.959 |
+FoveaBox | 0.955 | 0.873 | 0.623 | 0.806 | 0.896 | 0.832 | 0.748 | 0.881 | 0.768 | 0.878 |
+FreeAnchor | 0.973 | 0.841 | 0.855 | 0.928 | 0.864 | 0.939 | 0.781 | 0.89 | 0.68 | 0.887 |
+FSAF | 0.95 | 0.836 | 0.567 | 0.886 | 0.823 | 0.84 | 0.632 | 0.918 | 0.659 | 0.8 |
+GFL | 0.978 | 0.837 | 0.575 | 0.81 | 0.908 | 0.839 | 0.836 | 0.913 | 0.549 | 0.903 |
+LD | 0.98 | 0.893 | 0.579 | 0.862 | 0.927 | 0.871 | 0.743 | 0.917 | 0.676 | 0.91 |
+NAS-FPN | 0.975 | 0.916 | 0.768 | 0.932 | 0.944 | 0.946 | 0.817 | 0.937 | 0.766 | 0.925 |
+PAA | 0.966 | 0.923 | 0.861 | 0.952 | 0.905 | 0.937 | 0.895 | 0.938 | 0.951 | 0.92 |
+RetinaNet | 0.98 | 0.859 | 0.764 | 0.915 | 0.934 | 0.89 | 0.775 | 0.921 | 0.729 | 0.9 |
+RTMDet | 0.982 | 0.954 | 0.943 | 0.971 | 0.958 | 0.971 | 0.975 | 0.98 | 0.99 | 0.937 |
+TOOD | 0.908 | 0.763 | 0.666 | 0.83 | 0.834 | 0.826 | 0.717 | 0.847 | 0.622 | 0.853 |
+VarifocalNet | 0.977 | 0.802 | 0.486 | 0.77 | 0.9 | 0.832 | 0.671 | 0.866 | 0.466 | 0.829 |
+YOLOv5 | 0.975 | 0.979 | 0.967 | 0.974 | 0.977 | 0.974 | 0.974 | 0.972 | 0.985 | 0.966 |
+YOLOv6 | 0.985 | 0.982 | 0.968 | 0.974 | 0.983 | 0.979 | 0.974 | 0.984 | 0.978 | 0.973 |
+YOLOv7 | 0.962 | 0.924 | 0.931 | 0.948 | 0.939 | 0.958 | 0.932 | 0.96 | 0.932 | 0.931 |
+YOLOv8 | 0.975 | 0.919 | 0.903 | 0.93 | 0.942 | 0.96 | 0.916 | 0.965 | 0.885 | 0.917 |
+YOLOX | 0.955 | 0.877 | 0.853 | 0.902 | 0.91 | 0.945 | 0.926 | 0.918 | 0.868 | 0.891 |
+Faster R-CNN | 0.846 | 0.479 | 0.268 | 0.493 | 0.595 | 0.593 | 0.417 | 0.752 | 0.337 | 0.532 |
+Cascade R-CNN | 0.854 | 0.539 | 0.382 | 0.591 | 0.65 | 0.689 | 0.448 | 0.78 | 0.368 | 0.602 |
+Cascade RPN | 0.973 | 0.884 | 0.741 | 0.933 | 0.898 | 0.957 | 0.885 | 0.967 | 0.868 | 0.891 |
+Double Heads | 0.849 | 0.597 | 0.416 | 0.654 | 0.621 | 0.725 | 0.459 | 0.819 | 0.488 | 0.594 |
+FPG | 0.912 | 0.763 | 0.624 | 0.848 | 0.761 | 0.866 | 0.587 | 0.907 | 0.61 | 0.748 |
+Grid R-CNN | 0.873 | 0.585 | 0.39 | 0.672 | 0.653 | 0.794 | 0.488 | 0.836 | 0.495 | 0.614 |
+Guided Anchoring | 0.975 | 0.962 | 0.914 | 0.966 | 0.962 | 0.966 | 0.839 | 0.961 | 0.929 | 0.956 |
+HRNet | 0.844 | 0.655 | 0.429 | 0.687 | 0.627 | 0.732 | 0.423 | 0.875 | 0.532 | 0.609 |
+Libra R-CNN | 0.959 | 0.828 | 0.599 | 0.824 | 0.892 | 0.775 | 0.805 | 0.92 | 0.563 | 0.858 |
+PAFPN | 0.856 | 0.61 | 0.337 | 0.591 | 0.661 | 0.659 | 0.49 | 0.805 | 0.434 | 0.569 |
+RepPoints | 0.978 | 0.903 | 0.595 | 0.874 | 0.885 | 0.884 | 0.833 | 0.945 | 0.883 | 0.848 |
+Res2Net | 0.911 | 0.692 | 0.541 | 0.757 | 0.763 | 0.793 | 0.511 | 0.852 | 0.646 | 0.761 |
+ResNeSt | 0.929 | 0.646 | 0.497 | 0.727 | 0.701 | 0.691 | 0.685 | 0.802 | 0.515 | 0.716 |
+SABL | 0.856 | 0.545 | 0.387 | 0.646 | 0.656 | 0.636 | 0.488 | 0.774 | 0.407 | 0.571 |
+Sparse R-CNN | 0.959 | 0.877 | 0.513 | 0.759 | 0.913 | 0.746 | 0.733 | 0.882 | 0.605 | 0.871 |
+DETR | 0.746 | 0.328 | 0.232 | 0.256 | 0.468 | 0.383 | 0.457 | 0.369 | 0.234 | 0.468 |
+Conditional DETR | 0.962 | 0.831 | 0.73 | 0.931 | 0.865 | 0.934 | 0.881 | 0.964 | 0.839 | 0.917 |
+DDQ | 0.983 | 0.976 | 0.972 | 0.979 | 0.975 | 0.975 | 0.977 | 0.974 | 0.983 | 0.969 |
+DAB-DETR | 0.98 | 0.909 | 0.928 | 0.97 | 0.948 | 0.968 | 0.946 | 0.924 | 0.91 | 0.934 |
+Deformable DETR | 0.954 | 0.907 | 0.704 | 0.902 | 0.879 | 0.924 | 0.766 | 0.905 | 0.91 | 0.88 |
+DINO | 0.975 | 0.895 | 0.883 | 0.953 | 0.923 | 0.958 | 0.922 | 0.953 | 0.912 | 0.924 |
+PVT | 0.948 | 0.953 | 0.827 | 0.936 | 0.957 | 0.956 | 0.901 | 0.963 | 0.954 | 0.886 |
+PVTv2 | 0.973 | 0.942 | 0.884 | 0.952 | 0.934 | 0.973 | 0.835 | 0.967 | 0.941 | 0.867 |
+Please note that the categorization is based on the selected version of the object detection methods, which means the category may vary with their different versions, such as the backbone of a detector can be CNN or Transformer.
+Please refer to A.3 for the corresponding config files of the detectors.
 3.3 Object detectors
@@
 3.4 Evaluation and Analysis
+Evaluation metrics.
 To rigorously assess the efficacy of physical attacks on object detection systems, we furnish baseline datasets: clean datasets (without perturbations) and those infused with randomized noise (incorporating arbitrary disturbances in -bounded space).
-This dual-baseline approach sets the stage for a thorough examination.
-Quantifying performance entails employing evaluation metrics that consider both detection accuracy and the performance of the attacks.
-These metrics comprise by several widely adopted indicators, including average recall (AR) and Attack Success Rate (ASR).
-AR is calculated as the average of the recall rates:
-where TP and FN denote the true positive and false negative counts of the detector, respectively.
-AR signifies the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given target.
-On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as:
-where and denote the average recall rates of the attack and clean datasets, respectively.
-ASR provides a direct measure of the extent to which the attacks undermine the detector’s confidence in its predictions.
-Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating deeper insights into model behavior.
+This dual-baseline approach sets the stage for a thorough and fair examination.
+Quantifying performance entails employing evaluation metrics that consider object detection and adversarial attack performance.
+These metrics comprise several widely adopted indicators, including mean average precision (mAP), mean average recall (mAR), and attack success rate (ASR).
+mAP and mAR are calculated as the mean value of average precisions and recalls at recall and precision levels over classes, respectively, i.e., and .
+Precision rate and recall rate are calculated as and , respectively, where TP, FP, and FN denote the true positive, false positive, and false negative counts of the detector, respectively.
+On the other hand, ASR quantifies the effectiveness of the adversarial perturbations, calculated as , where and denote the value of adopted metric on the attack and clean datasets, respectively.
+ASR directly measures the extent to which the attacks undermine the detector’s performance.
+Advocacy of mAR for physical attacks. Adversarial attacks aim to induce mispredictions, i.e., to maximize error rate, which is the mathematical expectation of incorrect predictions written as:
+where is 1 for a correct prediction and 0 otherwise, and and represent the ground truths and predicted results of all objects, respectively.
+According to the calculation of performance metrics for detection, we can rewrite the error rate as follows:
+Therefore, mAR is a more direct and intuitive metric for evaluating the effectiveness of physical attacks on object detection models.
+We use mAR as the primary metric in the main manuscript, while mAP is also provided for reference.
+Evaluation perspectives.
+Specifically, we use mAP50, i.e., the confidence threshold of 0.5, to evaluate the overall performance of object detection, which is widely adopted in the object detection community.
+mAR50 is adopted to signify the proportion of correctly identified instances relative to the actual total in the dataset, offering an intuitive gauge of how physical attacks degrade the detection capability of a given adversarial target.
+However, mAR50 and mAP50 cannot fully reflect the performance of object detection models, especially when the confidence score of an adversarial object is significantly dropped but still higher than the threshold.
+To address this issue, we also use mAR50:95 and mAP50:95, calculated as the mean value over the range of 0.5 to 0.95 of the confidence threshold, to provide a more comprehensive evaluation of the object detection models.
+From the perspective of physical attacks, we use ASR over the detection metrics mAP50, mAR50, mAP50:95, and mAR50:95 to evaluate the effectiveness of physical attacks on object detection models, ensuring a comprehensive and impartial assessment.
+Moreover, we also visualize the distribution of evaluation performance using violin plots, which can provide a more intuitive understanding of the performance of object detection models and physical attacks, respectively.
+Analysis tools.
+Furthermore, we enhance our codebase by incorporating several ready-to-use explainability visualization tools, facilitating more profound insights into model behavior.
 These include Grad-CAM [132] for visualizing the regions of input data that contribute most to the model’s prediction, Shapley value [133] to quantify the individual feature contributions, and t-SNE [134] for reducing dimensionality and visualizing high-dimensional data in a more interpretable manner.
@@
 1) Overall experiments.
-We generate overall datasets with 3 objects, 10 weather conditions, 2 altitude angles, 8 azimuth angles, 5 radius values, 3 spawn points, and 20 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison.
+We generate overall datasets with three objects, ten weather conditions, two altitude angles, eight azimuth angles, five radius values, three spawn points, and 23 physical perturbations, i.e., 7200 samples () for each attack method, in which the physical dynamics are strictly aligned and controlled for impartial comparison as shown in Fig.
+3.
+Please note that these parameters are adjustable in the pipeline, and the datasets can be easily generated with different settings as needed.
 2) Ablation Studies.
-We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, altitude angle within a hemispherical space.
+We conduct in-depth examinations to explore the individual impact of core physical dynamics: weather conditions, venue, camera distance, azimuth angle, and altitude angle within a hemispherical space.
 Accomplishing this involves generating focused sub-benchmarks, each consisting of 100 samples.
 Physical attacks.
-we generate 24 datasets, including 20 physically noised datasets that correspond to 20 physical attacks, an extra 2 clean datasets and 2 randomly noised datasets for comparison of vehicle detection and person detection, respectively.
+We generate 24 datasets for comprehensive evaluation, including 20 physically noised datasets corresponding to 20 physical attacks, an extra two clean datasets, and two randomly noised datasets for comparison of vehicle and person detection, respectively.
+To evaluate the attack transferability, we also adopt perturbations optimized for aerial detection [135] and depth estimation [6, 87] in the experiments.
+Furthermore, we generate four extra datasets concerning traffic sign detection to show the easy extension of the benchmark to other objects (refer to A.2 for more details).
+The involved physical attacks are detailed in Table 1.
 Object detectors.
-We evaluate 48 object detectors, as shown in Table 1, by integrating MMDetection [9] into our evaluation pipeline, in which the selected detectors are trained on the COCO [50] dataset.
-Therefore, we conduct a total of 8064 () groups of the experiment, which are conducted with NVIDIA Geforce 4090.
+We evaluate 48 object detectors covering mainstream types, such as one- and two-stage and transformer-based detectors, as shown in Table 2, by integrating MMDetection [131] into our evaluation pipeline.
+Therefore, we conduct 8256 () groups of the experiment, which are conducted with NVIDIA GeForce 4090.
 4.2 Overall experiments and analysis
-We present the comprehensive results of vehicle detection against physical attacks in Fig.
-2.
-Additionally, Fig.
-3 and Fig.
-4 show visualized analyses of the experimental results from detection and attack perspectives, respectively.
-Experimental results of person detection and corresponding detailed numerical results are listed in the Appendix.
-From these analyses, several key observations emerge:
-Detection perspective: 1) Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing up to 50%, as shown in Fig.
-4.
-However, pedestrian detection performance is less affected regarding various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Fig.
-3.
-The potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as shown in the ablation experiments, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks.
-2) For different detectors, the performance varies significantly, with some detectors exhibiting superior robustness against physical attacks like EfficientNet, YOLO series and RTMDet of one-stage detectors.
-Besides, DDQ also shows significant adversarial robustness among transformer-based detectors.
-While the rest of the detectors show varying lower levels of robustness, the SOTA detection performance is not necessarily indicative of adversarial robustness.
-Consequently, the benchmark also serves as a robustness indicator.
-Attack perspective: 1) For vehicle detection, different physical attacks exhibit varying levels of effectiveness, with some attacks achieving ASR values exceeding 70% like ACTIVE, and others failing to surpass 20%.
-Most of the physical attacks hard to fool the latest SOTA detectors, such as EfficientNet, YOLO series, and RTMDet.
-This phenomenon is caused by the victim models of the attack method lagging behind the development of the detection method, which also motivates us to fill this gap.
-2) For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with the majority of attacks achieving ASR values below 20%, even lower than 0% (see more explanation in the Appendix).
-The relatively strongest attack method is AdvTexture, which elaborates on a 2D patch but with tricks for 3D space.
-This also demonstrates the gap between 2D perturbations and 3D physical space.
+We present the comprehensive results of vehicle detection against physical attacks in Table 3 and Figure 6.
+Additionally, Figures 7 and 8 provide visualized analyses of the experimental results from a detection perspective, while Figures 9 and 10 present the results from an attack perspective.
+More experimental results and corresponding detailed numerical data are listed in B.
+From these evaluations, several vital observations emerge.
+1) Detection Perspective.
+Vehicle detection performance is significantly impacted by physical attacks, with the average recall rates of detectors decreasing by up to 50%, as shown in Figure 8.
+In contrast, pedestrian detection performance is less affected by various attacks, with the average recall rates of detectors decreasing by less than 20%, as shown in Figure 7.
+The potential reason for this difference is that stronger physical perturbations targeting vehicles are optimized considering 3D space and accommodate more complex physical dynamics.
+On the other hand, physical attacks aimed at fooling person detectors are commonly performed using optimized 2D patches, which are effective under specific physical dynamics.
+This observation is further detailed in the ablation experiments (B.2.2), which empirically demonstrate the pressing need for a comprehensive and rigorous benchmark for physical attacks.
+The performance of different detectors varies significantly.
+Some detectors exhibit superior robustness against physical attacks, such as EfficientNet, the YOLO series, and RTMDet among one-stage detectors.
+Additionally, DDQ demonstrates notable adversarial robustness among transformer-based detectors.
+Other detectors show varying degrees of lower robustness, and interestingly, state-of-the-art detection performance does not necessarily correlate with adversarial robustness.
+Consequently, the benchmark also serves as an essential indicator of robustness.
+2) Attack perspective.
+For vehicle detection, different physical attacks exhibit varying levels of effectiveness.
+Some attacks achieve ASR values exceeding 70%, such as ACTIVE, while others fail to surpass 20%.
+Most physical attacks struggle to fool the latest state-of-the-art (SOTA) detectors, including EfficientNet, the YOLO series, and RTMDet.
+This phenomenon is attributed to the attack methods often lagging behind the rapid development of detection techniques, which motivates us to bridge this gap.
+For person detection, the ASR values of physical attacks are generally lower than those for vehicle detection, with most attacks achieving ASR values below 20%.
+The most effective attack method is AdvTexture, which uses a 2D patch with additional techniques to account for 3D space.
+This phenomenon highlights the significant gap between 2D perturbations and 3D physical environments, emphasizing the challenges in effectively transferring adversarial attacks from controlled 2D settings to more complex 3D scenarios.
+Furthermore, it underscores the need to develop more sophisticated attack strategies that can better account for the intricacies of 3D physical dynamics.
 4.3 Ablation experiments and analysis
-Except for the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors.
+In addition to the overall experiments, we also conduct ablation experiments to investigate the impact of physical world factors.
 We show the results of 3 physical dynamics, including weather, distance, and camera viewing angle, in Fig.
 9 and Fig.
-6, respectively.
-More experiments on other dynamics are provided in the Appendix.
-From these analyses, several key observations emerge: 1) Physical attack performance can be easily swayed by physical dynamics.
-This phenomenon is consistent with our intuition and emphasizes the importance of strictly aligning physical dynamics when evaluating physical attacks, which are often underestimated by previous works.
-2) We also observe a gap between the attack performance of our benchmark and the reported performance in the original papers (supplemented in Appendix).
+10, respectively.
+More experiments on other dynamics are provided in B.3 and B.5.
+From these evaluations, several key observations emerge:
+We also observe a gap between the ablation attack performance of our benchmark and the reported performance in the original papers (refer to B.2.1 for more details).
 Two reasons may contribute to this gap: the first is the adopted SOTA detectors in our benchmark, which are more robust than the victim models in the original papers, and the second is that our benchmark provides more comprehensive and strict evaluation datasets and physical dynamics, which are more challenging for the attack methods.
-Please refer to the Appendix for more experiments, detailed analysis and discussion.
+These observations empirically demonstrate the pressing need for a comprehensive and rigorous benchmark for physical attacks. Please refer to B for more experiments, detailed analysis, and discussion.
 5 Discussion
-Limitations.
-For now, PADetBench mainly focuses on the physical attacks on object detection models.
-In the future, we will extend the benchmark to evaluate other vision tasks, such as instance segmentation, 3D object detection, and depth estimation.
-Potential impacts.
-1) Positive impacts: In-depth understanding of the robustness of object detection models against physical attacks, which can guide the development of more robust models.
-2) Negative impacts: The benchmark may be misused to conduct physical attacks in real life, which may pose a threat to the security-critical applications involving intelligent visual perception.
+5.1 Where are we?
+1) Lack of alignment and comprehensiveness in physical dynamics.
+Existing works often suffer from limitations in comprehensiveness and fail to strictly control and align physical dynamics, as illustrated in Fig.
+2.
+Previous studies, such as those by Zhong et al.
+[85] and Dong et al.
+[66], have demonstrated that physical dynamics can be exploited to deceive DNNs.
+This observation underscores the critical necessity of aligning these dynamics to ensure fair and accurate evaluations.
+Without a comprehensive and rigorously aligned study, researchers cannot accurately gauge the progress in this research domain, which impedes advancements in the field.
+A robust benchmark that meticulously controls and aligns physical dynamics is essential for fostering the development of more resilient and reliable object detection models in the face of physical adversarial attacks.
+2) Discrete and naive physical adaptation.
+While theoretically, well-studied digital attacks should enhance the effectiveness of physical attacks, the practical outcomes often fall short of expectations.
+This discrepancy primarily stems from the fact that theoretical advancements do not always survive the cross-domain transformations ( as discussed in Section 2.2.
+These transformations are essential for translating digital models into real-world applications, but introduce significant challenges that are not adequately addressed by existing methods.
+Current research typically employs discrete and simplistic augmentation techniques to simulate physical dynamics.
+However, these methods fail to capture the real-world physical scenarios’ continuous and complex nature.
+As a result, the performance gaps observed in our ablation experiments (see B.2.1) are not surprising.
+These gaps highlight the limitations of current approaches and emphasize the need for a more comprehensive and rigorous benchmark.
+5.2 Where to go?
+1) Comprehensive and Physically Aligned Benchmark
+A comprehensive and physically aligned benchmark is essential for evaluating physical attacks on object detection models.
+Such a benchmark ensures rigorous and unbiased assessments, enabling researchers to identify the strengths and weaknesses of various attack methods and defense mechanisms.
+Providing detailed insights facilitates the development of more robust and resilient object detection models.
+This move is crucial for enhancing the security and reliability of AI systems in real-world applications, where the ability to withstand physical attacks is paramount.
+A well-designed benchmark should encompass various physical scenarios, including lighting conditions, angles, and environmental factors.
+It should also include diverse datasets and evaluation metrics to ensure the models are tested under realistic conditions.
+By doing so, the benchmark can help identify specific areas where current models and defenses fall short, guiding future research and development efforts.
+Ultimately, a comprehensive and physically aligned benchmark will drive innovation and improve AI systems’ overall security and performance in practical settings.
+2) Rigorous and Differentiable Modeling of Cross-Domain Transformations
+Accurate modeling of cross-domain transformations is crucial for both physical attacks and defenses.
+These transformations involve converting digital perturbations into physical perturbations and vice versa, which introduces significant challenges due to the differences in the domains.
+While existing works have attempted to use differentiable neural renderers to generate adversarial examples automatically, these methods often have limited modeling capabilities and struggle to align physical factors between perturbed and clean images accurately.
+The advent of large foundation models offers a promising direction for addressing these limitations.
+Large-scale data and advanced foundation models can provide the necessary complexity and detail to more rigorously model physical dynamics.
+By leveraging these resources, researchers can develop more sophisticated methods for simulating real-world physical perturbations, leading to more effective attack simulations and defensive strategies.
+For example, large foundation models can be trained on extensive datasets that capture various physical environments and conditions.
+This training can help the models learn to account for various physical factors, such as lighting variations, texture properties, and camera distortions.
+Differentiable rendering techniques can also be integrated with these models to ensure the generated perturbations are realistic and physically plausible.
+In summary, exploring how to model physical dynamics more rigorously using large-scale data and foundation models is a promising direction.
+This approach has the potential to significantly improve the realism and effectiveness of both attack simulations and defensive strategies, thereby advancing the field of physical security in AI systems.
 6 Conclusion
-In conclusion, this work aims to meet the critical need for a standardized and rigorous benchmark for physical attacks on object detection models.
-Our benchmark leverages the CARLA simulator for consistent physical dynamics and includes 20 attack methods, 48 detectors, comprehensive dynamics, and metrics from different perspectives.
-It also provides end-to-end pipelines for dataset generation, attack and detection evaluation, and experimental results analysis.
-Our extensive evaluation and analyses provide valuable insights for future research directions.
-The supporting codebase is publicly available and we will continuously supplement it, aiding the development of physically robust object detection methods.
-7 Acknowledgment
-The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision funded by The Hong Kong Jockey Club Charities Trust.
-Checklist
-1.
-For all authors…
-Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
-Did you describe the limitations of your work? [Yes]
-Did you discuss any potential negative societal impacts of your work? [Yes]
-Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
-2.
-If you are including theoretical results…
-Did you state the full set of assumptions of all theoretical results? [N/A]
-Did you include complete proofs of all theoretical results? [N/A]
-3.
-If you ran experiments (e.g. for benchmarks)…
-Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)?
-[Yes] Please refer to the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack for the code and instructions.
-Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A] This benchmark does not involve a training process.
-Did you report error bars (e.g., to the random seed after running experiments multiple times)? [N/A]
-Did you include the total amount of computing and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes]
-4.
-If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
-- (a)
-Did you mention the license of the assets? [Yes] Please refer to the Appendix for the license.
-Did you include any new assets either in the supplemental material or as a URL? [Yes] Please refer to the GitHub repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack) for details.
-Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [N/A]
-Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
-5.
-If you used crowdsourcing or conducted research with human subjects…
-Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
-Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
-Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]
+In conclusion, we develop a comprehensive simulation-based benchmark to evaluate physical attacks under controlled conditions rigorously.
+This benchmark encompasses 23 physical attacks, 48 object detectors, and detailed physical dynamics, all supported by end-to-end pipelines.
+The benchmark is designed to be flexible and scalable, facilitating the easy integration of new attacks, models, and vision tasks.
+Through extensive evaluations involving over 8,000 tests, we have identified vital algorithm limitations and provided valuable insights.
+We hope this benchmark will advance research in physical adversarial attacks, fostering the development of more robust and reliable models.
+CRediT authorship contribution statement
+Jiawei Lian: Conceptualization, Data curation, Formal analysis, Writing - original draft, Writing - review and editing.
+Jianhong Pan: Data curation, Formal analysis, Writing - original draft.
+Lefan Wang: Formal analysis, Writing - original draft.
+Yi Wang, Shaohui Mei, and Lap-Pui Chau: Supervision, Writing - review and editing.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgements
+The research work was conducted in the JC STEM Lab of Machine Learning and Computer Vision, funded by The Hong Kong Jockey Club Charities Trust.
+This work was supported partly by the National Natural Science Foundation of China (No.
+62171381).
+Data availability
+The datasets used in this study are publicly available and can be accessed through https://github.com/JiaweiLian/PADetBench.
+Appendices
+A Additional content of the Benchmark
+A.1 Mini-test
+A.2 The adaptability of the benchmark
+A.3 Corresponding config files of the selected detectors
+A.4 Explanation about the selected objects
+A.5 The necessity of the benchmark
+A.5.1 Utilities of the benchmark
+A.5.2 Potential applications of the benchmark
+A.6 Limitations and potential impacts
+B Additional content of the Experiments
+B.1 Generated data for ablation studies
+B.2 A detailed illustration of the performance gap
+B.2.1 Performance gap between the benchmark and the original papers
+B.2.2 Performance gap between attacks against vehicle and person detection
+B.3 Supplemented experiments analysis and discussion
+B.3.1 Detection perspective
+B.3.2 Attack perspective
+B.4 Additional oversall experimental experiments
+B.5 Additional ablation experiments
+B.5.1 Ablation study on physical dynamics
+B.5.2 Ablation study on training dataset
+B.5.3 Ablation study on 2D and 3D perturbations
+C User feedback
 Appendix A Additional content of the Benchmark
@@
 Town3 | A larger, urban map with a roundabout and large junctions. |
-Town4 | A small town embedded in the mountains with a special "figure of 8" infinite highway. |
-Town5 | Squared-grid town with cross junctions and a bridge. |
-It has multiple lanes per direction. Useful to perform lane changes. | |
+Town4 | A small town embedded in the mountains with a special ”figure of 8” infinite highway. |
+Town5 | Squared-grid town with cross junctions and a bridge. It has multiple lanes per direction. Useful to perform lane changes. |
 Town6 | Long many lane highways with many highway entrances and exits. It also has a Michigan left. |
 Town7 | A rural environment with narrow roads, corn, barns and hardly any traffic lights. |
-Town8 | Secret "unseen" town used for the Leaderboard challenge. |
-Town9 | Secret "unseen" town used for the Leaderboard challenge. |
+Town8 | Secret ”unseen” town used for the Leaderboard challenge. |
+Town9 | Secret ”unseen” town used for the Leaderboard challenge. |
 Town10 | A downtown urban environment with skyscrapers, residential buildings and an ocean promenade. |
-Town11 | A Large Map that is undecorated. |
-Serves as a proof of concept for the Large Maps feature. | |
-Town12 | A Large Map with numerous different regions, |
-including high-rise, residential and rural environments. |
-Full list of the optional maps, where Town8 and Town9 are unseen for competition. Please refer to CARLA [17] documentary for more details.
-A.2 Corresponding config files of the selected detectors
+Town11 | A Large Map that is undecorated. Serves as a proof of concept for the Large Maps feature. |
+Town12 | A Large Map with numerous different regions, including high-rise, residential and rural environments. |
+A.2 The adaptability of the benchmark
+We provide a detailed illustration of the scene diversity of the benchmark in Table 4 and Fig.
+12, where the optional maps are listed with their descriptions.
+In addition, we display the extendable vehicles, pedestrians, and traffic signs in Fig.
+13, Fig.
+14, and Fig.
+15, respectively, which can be easily extended to evaluate other objects in the benchmark.
+The users are also allowed to export any customized scenes and objects to the benchmark as needed, which can be easily integrated into the benchmark.
+A.3 Corresponding config files of the selected detectors
 The corresponding config files of the selected detectors are listed in Table 5.
+Specifically, 1-25 and 26-40 are CNN-based One-stage and Two-stage object detectors, respectively.
+41-48 are Transformer-based object detectors.
+The corresponding config files of the detectors are available in our codebase or MMDetection [131] toolbox.
 Number | Config Files | Detectors |
@@
 48 | retinanet_pvtv2-b0_fpn_1x_coco | PVTv2[126] |
+A.4 Explanation about the selected objects
+According to a survey [86] published in TPAMI 2024, most physical attacks against object detection are optimized for specific target categories, such as vehicles, persons, and a few for traffic signs.
+In line with this, we have chosen vehicles and pedestrians as the representative target categories, to evaluate the robustness of object detectors against physical attacks.
+.
+In order to ensure the validity of our benchmark for different types of objects, we have demonstrated that our benchmark can be easily extended to other target categories, as shown by the experiments conducted on traffic sign in Table 20.
+The benchmark is designed to evaluate the robustness of object detectors against physical attacks in various aligned scenarios for ensuring fairness.
+It can be extended to other target categories with minimal modifications.
+We have thoroughly reviewed over forty physical attack methods, and we found that most of these methods conducted experiments under unaligned conditions and without fair comparisons.
+This lack of clarity hinders the accurate assessment of the progress of physical adversarial attacks and the development of physical adversarial robustness.
+Therefore, we are motivated to establish a comprehensive and rigorous benchmark for physical attacks to address these limitations and provide a solid foundation for future research.
+A.5 The utility of the benchmark
+In this section, we sumarize our motivation and provide the potential applications of the benchmark.
+A.5.1 Utilities of the benchmark
+Standardization and Fair Evaluation: The primary utility of PADetBench lies in its ability to standardize the evaluation of physical attacks against object detection models.
+By ensuring that all evaluations are conducted under the same physical dynamics, PADetBench eliminates inconsistencies found in real-world experiments, making it a fair and rigorous benchmark.
+Comprehensive Coverage: PADetBench includes 23 physical attack methods and evaluates 48 state-of-the-art object detectors, providing a comprehensive coverage that enables researchers to compare and contrast various models and attack strategies.
+A.5.2 Potential applications of the benchmark
+Research and Development: Researchers developing robust object detection models or physical attack strategies need a benchmark to evaluate and compare their approaches.
+Security Assessments: Security teams need to assess the robustness of deployed object detection systems in critical infrastructure.
+Regulatory Compliance: Regulatory bodies require evidence of robustness and security for autonomous systems.
+Product Testing: Companies developing autonomous vehicles or security systems need to test their products under various physical attack scenarios.
+Educational Purposes: Educators and students need resources to understand the vulnerabilities of object detection models.
+A.6 Limitations and potential impacts
+Limitations
+For now, PADetBench primarily focuses on evaluating the robustness of object detection models against physical attacks.
+In the future, we plan to extend the benchmark to include other vision tasks, such as instance segmentation, 3D object detection, and depth estimation.
+This expansion will provide a more comprehensive evaluation framework that covers a broader range of computer vision applications.
+Potential Impacts
+1) Positive Impacts: The in-depth understanding gained through PADetBench will contribute significantly to the development of more robust object detection models.
+By identifying vulnerabilities and limitations, researchers and practitioners can design improved algorithms that are better equipped to handle physical adversarial attacks.
+This enhanced robustness is crucial for real-world applications where reliability and accuracy are paramount.
+2) Negative Impacts: While the benchmark provides valuable insights, there is a risk that it could be misused to conduct physical attacks in real-life scenarios.
+Such misuse could threaten the security of critical applications involving intelligent visual perception systems.
+Therefore, it is essential to promote responsible use of the benchmark and to emphasize the importance of ethical considerations in research and development.
 Appendix B Additional content of the Experiments
-B.1 Genetated data for the ablation studies
+B.1 Genetated data for ablation studies
 We provide the generated data samples for the ablation studies in Fig. 16.
-B.2 Detailed illustration of the gap between the attack performance of our benchmark and the reported performance in the original papers
-Our benchmark addresses two important aspects that were not considered in the original papers:
-1.
-Strict physical dynamics: Our benchmark leverages the CARLA simulator, which provides a realistic and consistent physical environment for evaluating physical attacks on object detection models.
-This ensures that the attacks are evaluated under conditions that closely resemble the physical world, allowing for more accurate assessment of the robustness of the models.
-2.
-Evaluation in 3D space perspective: Unlike the original papers, which mainly focus on 2D evaluation, our benchmark evaluates the attack performance in a 3D space perspective.
-This is crucial because physical attacks can exploit vulnerabilities in the 3D structure of the objects, such as comprehensive and continuous view, scale, distance, etc.
-By considering the 3D space perspective, our benchmark provides a more comprehensive evaluation of the models’ robustness against physical attacks.
-By addressing these two aspects, our benchmark provides a more realistic and comprehensive evaluation of the attack performance of object detection models, bridging the gap between the reported performance in the original papers and the real-world scenarios.
+B.2 A detailed illustration of the performance gap
+B.2.1 Performance gap between the benchmark and the original papers
+In this section, we provide an explanation for the performance gap between the reported attack performance in the original papers and the results in our benchmark.
+Our benchmark encompasses a wide range of physical dynamics, whereas previous validation settings are often limited to a few specific scenarios.
+The comprehensive physical dynamics in our benchmark reveal the shortcomings of existing object detectors and physical attacks and this is the main motivation of our work.
+Therefore, our benchmark might not be captured by previous validation settings, leading to the discrepancy between our results and the reviewer’s individual experiences.
+In comparison, we removed various physical dynamics including weather (rain, snow, fog), lighting (nighttime), and distance (far positions), and reproduced the results of several attack methods on YOLOv7 as reported in ACTIVE [5], which are listed in Table 23.
+It is worth noting that these reported results are also included in our benchmark with particular evaluation settings.
+Contrary to the simplified settings of these reproduced experiments, more comprehensive physical dynamics incorporated into our benchmark significantly highlight the ineffectiveness of existing physical attacks.
+These aspects may not have been adequately captured by previous validation settings.
+As illustrated in Tabel 23, when we exclude various dynamics, the effectiveness of physical attacks notably increases, thereby reducing the performance of object detectors.
+Therefore, our benchmark strive to encompass and align these physical dynamics for comprehensive and equitable comparisons.
+B.2.2 Performance gap between attacks against vehicle and person detection
+For the gap between attacks against vehicle and person detection, one reason is that these attacks are optimized to fool object detectors in particular target detection during training process.
+Consequently, we follow the attack purpose of the original works in this benchmark to attack specified target category accordingly for fairness, which partially accounts for the phenomenon that pedestrian detection performance is less affected regarding various attacks in comparison with car detection.
+Another potential reason is that the stronger physical perturbations are optimized with consideration of 3D space and accommodate more complex physical dynamics, while physical attacks aiming to fool person detectors are commonly performed with optimized 2D patches, which work well in particular physical dynamics, as evidenced by the ablation experiments in B.2.1, which empirically demonstrate the pressing need and necessity of a comprehensive and rigorous benchmark for physical attacks.
 B.3 Supplemented experiments analysis and discussion
@@
 B.4 Additional overall experimental results
-Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 4, 5, 6, 7, 8, 9, 10, and 11.
+Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 6, 7, 8, 9, 10, 11, 12, and 13.
 In addition, the visualized evaluation results are shown in Fig.
-9, 10, 11, 12, 13, 14, and 15.
+17, 18, 19, 20, 21, 22, and 23.
 Clean |
@@
 B.5 Additional ablation experiments
-Due to space constraints, we provide additional overall experimental results in this part, as shown in Table 12, 13, 14, 15, 16, and 17.
+B.5.1 Ablation study on physical dynamics
+Due to space constraints, we provide additional ablation experimental results in this part, as shown in Table 14, 15, 16, 17, 18, and 19.
 In addition, the visualized evaluation results are shown in Fig.
-16, 17, and 18.
-Appendix C Reproducibility
-The reproducibility of all evaluation results in our benchmark is ensured through the provision of the necessary code in the associated Github Repository (https://github.com/JiaweiLian/Benchmarking_Physical_Attack).
-Furthermore, all datasets and model architectures evaluated are accessible to the public.
-Appendix D License
-This work is under Apache-2.0 license. More detailed information can be found in the LICENSE file in the GitHub repository: https://github.com/JiaweiLian/Benchmarking_Physical_Attack
+24, 25, and 26.
+B.5.2 Ablation study on training dataset
+To further investigate the impact of the training dataset on the physical attacks, we collected ten physical attacks for fooling person detection, and the results are shown in Table 21, where the Median ASR represents the median attack success rate across the 48 detectors.
+It can be observed that physical attacks trained on the INRIA and COCO datasets achieve comparable performance in general.
+Physical attacks | Training datasets | Median ASR |
+AdvCam | ImageNet | 0 |
+AdvCaT | 376 self-collected images | 0 |
+MTD | - | 2 |
+LAP | INRIA | 2 |
+AdvPattern | Market1501 | 2 |
+AdvTshirt | 40 self-collected videos | 3 |
+DAP | INRIA | 5 |
+NaTPatch | INRIA | 5 |
+InvisCloak | COCO | 5 |
+AdvTexture | INRIA | 7 |
+B.5.3 Ablation study on 2D and 3D perturbations
+Physical attacks that evaluate 2D adversarial patches from a frontal perspective have a significant limitation, as they do not account for the effects of multiple viewing angles in a 3D environment.
+Our study aims to bridge this gap by developing a comprehensive benchmark for assessing physical attacks from various angles and incorporating a broader range of physical dynamics.
+During our investigation, we noted a substantial drop in performance (detection rate: ) when adversarial patches were only applied to the frontal view of objects.
+To ensure a fair comparison and enhance the efficacy of the attacks, we expanded the application of these patches to cover the entirety of the object’s surface.
+Additional experiments were conducted to assess the impact of adversarial patches on frontal views using several object detection algorithms.
+The results are summarized in Table 22.
+The ’Entire Surface’ column highlights cases where the adversarial patch was applied across the entire surface of an object.
+The values in parentheses indicate the relative decrease in performance compared to full-surface patching.
+Perturbations | Entire surface | CornerNet | VarifocalNet |
+Clean | - | 87 | 80 |
+Random | - | 87 | 77 |
+AdvTexture | ✓ | 74 | 73 |
+AdvTexture | 81(7) | 77(4) | |
+AdvPatch | ✓ | 82 | 75 |
+AdvPatch | 85(3) | 79(4) | |
+NatPatch | ✓ | 78 | 74 |
+NatPatch | 83(5) | 77(3) |
+Appendix C User feedback
+To ensure ease of use, we have addressed potential barriers by user feedback, such as CARLA deployment and customizing adversarial objects, by providing a comprehensive Docker installation guide for CARLA and a tutorial on customizing adversarial objects in our documentation.
+These resources enable users to install CARLA and customize objects in just a few minutes.
+We also conducted usability testing with five researchers from a well-known University and got feedback from them in the form of a survey questionnaire as shown in Table 25.
+The users consistently found the benchmark easy to use and provided positive feedback on its usability.
+Clean | Random | CAMOU | DTA | ACTIVE | ||
+YOLOv3 | Reported | 86 | 67 | 60 | 32 | 23 |
+Reproduced | 86 | 66 | 62 | 33 | 23 | |
+YOLOv7 | Reported | 93 | 86 | 83 | 59 | 42 |
+Reproduced | 93 | 85 | 83 | 60 | 41 | |
+PVT | Reported | 89 | 78 | 69 | 56 | 52 |
+Reproduced | 89 | 78 | 69 | 56 | 51 |
+Number | Questions |
+Q1 | How easy was it to follow the Docker installation guide for CARLA? (Rating 1-5) |
+Q2 | How helpful was the tutorial on customizing adversarial objects in the documentation? (Rating 1-5) |
+Q3 | Were you able to successfully deploy CARLA using the provided resources? (Yes or No) |
+Q4 | Were you able to successfully customize adversarial objects using the provided resources? (Yes or No) |
+Q5 | Overall, how satisfied are you with the ease of CARLA deployment and customizing adversarial objects? (Rating 1-5) |
+Questions | User1 | User2 | User3 | User4 | User5 |
+Q1 | 4 | 5 | 5 | 4 | 5 |
+Q2 | 5 | 5 | 5 | 5 | 5 |
+Q3 | Yes | Yes | Yes | Yes | Yes |
+Q4 | Yes | Yes | Yes | Yes | Yes |
+Q5 | 4 | 5 | 5 | 4.5 | 5 |